### v1

# Depth-Based Local Center Clustering: A Framework for Handling Different Clustering Scenarios

Abstract
Cluster analysis, or clustering, plays a crucial role across numerous scientific and engineering domains. Despite the wealth of clustering methods proposed over the past decades, each method is typically designed for specific scenarios and presents certain limitations in practical applications. In this paper, we propose depth-based local center clustering (DLCC). This novel method makes use of data depth, which is known to produce a center-outward ordering of sample points in a multivariate space. However, data depth typically fails to capture the multimodal characteristics of data, something of the utmost importance in the context of clustering. To overcome this, DLCC makes use of a local version of data depth that is based on subsets of data. From this, local centers can be identified as well as clusters of varying shapes. Furthermore, we propose a new internal metric based on density-based clustering to evaluate clustering performance on non-convex clusters. Overall, DLCC is a flexible clustering approach that seems to overcome some limitations of traditional clustering methods, thereby enhancing data analysis capabilities across a wide range of application scenarios.
Index Terms:
Clustering, statistical depth, similarity matrix, Mahalanobis depth, spatial depth, internal metric.
I Introduction
Clustering, or cluster analysis, is one of the most important topics in data science. A cluster is usually seen as a set of similar objects that have been gathered or grouped together [1]. Generally speaking, clustering is the task of partitioning a collection of objects, into separate clusters, in such a way that objects in the same cluster are as similar as possible and objects in separate clusters as distinct as possible [2]. Research on clustering methods has a long history in several disciplines, including biology, psychology, geography, and marketing [3]. In recent years, the rapid development of sensing and storage technology has resulted in the collection of vast quantities of high-dimensional, large-scale, data. In turn, this has increased the need for flexible and computationally efficient clustering methods [4] in many fields, such as pattern recognition [5], computer vision [6, 7] and image processing [8].
The complexity of data distributions has led to the development of numerous clustering algorithms, each having a unique set of strengths and constraints. We next briefly discuss different traditional clustering methods to highlight their distinct characteristics and limitations. In this, and in what follows, we consider the objects to be clustered as points in the $d$ dimensional real space $\mathbb{R}^{d}$.
I-A Clustering Methods
I-A1 Center-Based Clustering
Center-based clustering methods, with k-means clustering [9] being a prime example, are often utilized due to their simplicity and effectiveness. However, k-means clustering has several limitations. Firstly, it is sensitive to outliers as it uses the means of the points in each cluster as cluster centers. Secondly, even in outlier-free contexts, this approach is most adequate when clusters are spherically symmetric, which may not always be representative of the actual data distribution. Lastly, the algorithm is sensitive to initialization and can easily get trapped in local minima. To address some of these issues, several variants of the k-means algorithm have been developed:
  • Partitioning around medoids (PAM; [10]), was introduced to combat the sensitivity to outliers. Instead of cluster means, PAM uses medoids (a form of multivariate median) as cluster centres.
  • Weighted k-means (W-k-means; [11]) and lasso weighted k-means (LW-k-means; [12]) were designed to address the underlying spherical symmetry assumption for clusters. These both provide a more flexible k-means approach by incorporating a weight vector for each variable. LW-k-means improves W-k-means by adding a lasso penalty to the objective function, enabling feature selection in high-dimensional data.
  • Power k-means [13] and kernel power k-means [14] were developed to address the sensitivity of k-means to initialization and tendency to get trapped in local minima by redefining the objective function using power means. Kernel power k-means extends this by incorporating multiple kernel configurations to handle non-linear separable data via higher-dimensional projections.
I-A2 Density-Based and Hierachical Clustering
Density-based clustering assumes that clusters are regions of high density separated by low-density areas, making it well-suited for detecting arbitrarily shaped clusters. For instance, density-based spatial clustering of applications with noise (DBSCAN; [15]) sets a threshold and minimum point count for a region to be considered dense. Then, dense regions that can be directly connected with each other are merged into a cluster. However, DBSCAN’s reliance on a fixed threshold limits its ability to handle clusters with varying densities.
Hierarchical clustering, on the other hand, adopts a tree-like approach to iteratively merge or split clusters. It offers two main strategies: agglomerative (bottom-up) and divisive (top-down). The resulting dendrogram represents cluster relationships, from an all-inclusive cluster at the top to individual objects at the bottom [16]. While hierarchical clustering provides an intuitive structure, its high computational cost and irreversible decisions make it less suitable for large datasets or complex cluster shapes.
Density peak clustering (DPC) [17], introduced in $2014$, takes a different approach by combining density and distance to identify cluster centers. It locates density peaks based on a decision graph, where density is plotted on the y-axis and the distance to the nearest higher-density point on the x-axis. This method allows for manual selection of cluster centers and performs well with varying-density clusters. However, it can struggle with robustness, as density peaks may correspond to satellite points rather than true cluster centers.
An improved version, fast main density peak clustering within relevant regions (R-MDPC) [18], proposed in $2024$, incorporates hierarchical concepts by applying single-linkage clustering to density peaks, defining relevant regions as density-connected areas. This enhancement better handles complex cluster shapes by merging single-density peak clusters and reducing satellite peak interference. However, for highly overlapping or extremely high-dimensional datasets, R-MDPC may collapse all points into one region, behaving similarly to the original DPC.
I-A3 Model-Based Clustering
Model-based clustering typically assumes the points to be clustered stem from an underlying mixture distribution [19], whose components are naturally associated with the clusters to be constructed. For example, working from a Gaussian mixture model (GMM), one can identify elliptical clusters with varying orientations, shapes, and sizes. Other mixture models handling asymmetric distributions have also been developed, see, e.g., [20], [21], [22], and [23]. These different families of mixture models lead to clustering methods typically relying on the expectation-maximization (EM) algorithm [24] to find optimal parameter values, which can be computationally intensive. Moreover, selection of the most appropriate family of mixture models can be challenging unless knowledge about the data-producing mechanism and/or subject area expertise is available.
I-A4 Spectral and Subspace Clustering
Spectral clustering [25], a notable graph-based method, treats clustering as a graph partitioning problem, converting data points into nodes and their similarities into connecting edge weights. It is well-suited for high-dimensional datasets, often utilizing a sparse adjacency matrix derived from the k-nearest neighbors methods. After representing the data as a similarity graph, the clustering problem can essentially be seen as partitioning the graph in such a way that edges between groups are given low weights and edges within groups high weights [26].
Spectral clustering has been extended and has led to other clustering techniques. It is, for instance, an essential component of modern subspace clustering algorithms. These algorithms have gained attention due to their capacity for identifying low-dimensional representations of high-dimensional data [27]. In particular, sparse subspace clustering (SSC) [28] is a key evolution in this domain. Making use of the self-expressiveness property, points are expressed as linear combinations of other points within the same subspace. Therefore, SSC merely requires the coefficient matrix of these linear combinations. As a result, SSC overcomes two major challenges of spectral clustering: selecting an appropriate neighborood size and handling points near the intersection of two subspaces. However, for low-dimensional data, the premise that data points can be effectively represented as linear combinations of the lower-dimensional subspace may falter, potentially leading to a diminished efficacy of SSC. This makes SSC a tool that is primarily designed for high-dimensional datasets.
I-B Motivation
Our objective is to develop a versatile clustering framework capable of addressing diverse challenges, from convex to non-convex cluster shapes, and from low- to high-dimensional data. A common theme across clustering methods is their reliance on representative points, or exemplars. Center-based methods define exemplars by minimizing within-cluster distances, often relying on distance metrics. However, a specific distance measure often performs well only for certain types of data, and adapting to more diverse structures may require mapping data into higher-dimensional spaces via kernel methods. The effectiveness of such approaches, however, hinges on selecting an appropriate kernel, which can be challenging and highly sensitive to the data. Density-based methods, on the other hand, define exemplars through local density variations. These methods identify clusters by connecting dense regions, but the appropriate definition of density often depends on the dimensionality of the data. For example, density measures that work well in low-dimensional spaces may lose effectiveness in high-dimensional settings due to the sparsity of data. Moreover, in overlapping regions, density connections often struggle to distinguish between clusters. Both center-based and density-based approaches share a reliance on exemplars but are inherently limited by their respective definitions of exemplars.
A versatile clustering framework must overcome these limitations by introducing an adaptive similarity measure that avoids strong parametric assumptions and generalizes across diverse data geometries. Equally important is the need for a robust exemplar definition that is consistent across dimensions and cluster shapes, along with effective grouping strategies for these exemplars to form meaningful clusters.
In response, our method introduces a more flexible exemplar definition by leveraging statistical data depth, a nonparametric tool for multivariate analysis. Tukey [29] pioneered the idea of center-outward ordering in multivariate data, with the introduction of the first depth definition, half-space depth. Data depth enables the center-outward ordering of data, making it a potent tool for locating a center for any dataset. When considering subsets of the data, data depth with respect to each subset aids in defining multiple exemplars. Naturally, as the subset size grows, the number of exemplars diminishes, and vice versa.
Building on this concept, we propose the depth-based local center clustering (DLCC) algorithm. Equipped with two grouping strategies for depth-defined representative points, DLCC inherits strengths from both center-based and density-based methods, while offering greater flexibility to tackle a wide range of clustering problems. Here, the term “depth-based” encompasses two aspects: (1) data depth is used to construct a similarity matrix instead of relying on distance, adaptively capturing the geometric characteristics of each point’s location; (2) data depth measures the centrality of points within each point’s neighborhood, helping to identify exemplars.
Furthermore, ground truth is typically not available in real-world clustering applications, necessitating the use of internal metrics to evaluate performance. Established metrics like the average silhouette width (ASW; [49]), which measures the average distance within clusters relative to the nearest cluster, and the Calinski-Harabasz index (CHI; [50]), which relies on distances from points to their cluster centroids, inherently favor spherical clusters. This preference limits their effectiveness in capturing non-convex shapes, as illustrated in Figure 1, where they will incorrectly favor clustering result (c) over (a). To address these shortcomings, we propose a new internal metric, the depth-based clustering (DC) metric, designed to more accurately handle complex cluster shapes and assist in parameter selection for our algorithm. Further details will be discussed in Section IV-B.
Figure 1: Clustering results for the synthetic dataset SmileyF. Panel (a) shows the ground truth, (b) misclusters several points in the eyes, and (c) fails to capture the face pattern.
I-C Main Contributions and Organization of the Paper
This paper makes the following main contributions:
  • Depth-Based Similarity Matrix: We improve the efficiency when constructing a depth-based similarity matrix using Mahalanobis depth (MD) and Spatial depth (SD). For the MD-based similarity matrix, model-based clustering methods are employed to predefine covariance matrices that approximate the data’s distribution shapes. For the SD-based similarity matrix, we develop an efficient matrix-based computation method.
  • DLCC Clustering Framework: We develop a novel clustering framework, DLCC, which includes the definition of a similarity measure and local centers, and the introduction of methods for filtering and grouping local centers. Experiments shows it achieves competitive performance across different types of clustering challenges among the representative algorithms from each type of clustering method referenced in Section I-A.
  • Internal Evaluation Metric: We propose a depth-based clustering metric as an internal evaluation criterion. This metric combines the density-based clustering concept with the depth-based similarity matrix to evaluate clustering performance, particularly for non-convex cluster structures.
The remainder of the paper is organized as follows: Section II reviews key concepts of data depth and introduces the depth-based similarity matrix. Section III provides a detailed description of the DLCC framework. In Section IV, we discuss the hierarchical structure of DLCC and present the proposed internal metric for parameter selection. Section V showcases experimental results of DLCC using both synthetic and real-world datasets. Lastly, Section VI concludes the paper.
Figure 2 provides the flowchart of the DLCC method, and Table I summarizes the key notations used in this paper to help readers navigate its contents more easily. Additional graphs providing further insights into our methods are included in the supplementary file.
DatasetDepth-basedsimilarity matrixSection IINeighborhoods &Local centersSection III Filtered centersSection III-AGroups of filtered centersSection III-A &\&& Section IVTemporaryclustersSection III-BUnlabelled pointsSection III-BFinal clustersclassification
Figure 2: Flowchart for the DLCC algorithm, detailing the sections that include the key contents represented in the nodes.
TABLE I: Summary of the main notations
Notation Description
$n$ and $d$ Number of observations and dimensions, respectively;
$D_{\mathrm{MD/SD}}(\mathbf{z}|\mathbf{X})$ Mahalanobis or spatial depth of $\mathbf{z}$ with respect to $\mathbf{X}$;
$\mathbf{X}$ A given dataset;
$\mathbf{X}_{Ri}$ Dataset reflected around $\mathbf{x}_{i}$, i.e., $\mathbf{X}_{Ri}=\mathbf{X}\cup\{2\mathbf{x}_{i}-\mathbf{x}_{j}\,|\,j\neq i\}$;
$\mathbf{E}$, $\mathbf{\tilde{E}}$ Matrices of unit vector sums from $\mathbf{X}$ and reflected points to $\mathbf{X}$;
$\mathbb{S}$ Depth-based similarity matrix, with $\mathbb{S}_{ij}$ its $ij$-th entry;
$s$ Neighborhood size;
$\mathbf{N}_{i}$ Depth-based neighborhood of $\mathbf{x}_{i}$;
$r_{i}(\mathbf{x}_{j})$ Depth rank of $\mathbf{x}_{j}$ within $\mathbf{N}_{i}$;
$c$, $\mathbf{c}$, $\mathbf{C}$ A local center, a certain set of local centers, and the set of all filtered centers, respectively;
$\mathbb{M}$ Local center neighborhoods’ similarity matrix, with $\mathbb{M}_{ij}$ its $ij$-th entry;
$\delta$ Similarity threshold for local center grouping ($\mathbb{M}$-based);
$\mathbf{g}$ A group of local centers;
$\mathbf{U}_{\mathbf{g}}$ Unique neighbors of group $\mathbf{g}$;
$G$ and $K$ Number of groups and clusters, typically $G=K$;
$k$ Cluster index;
$\mathcal{X}_{k}$ Observations in cluster $k$;
$f$ Frequency of a local center;
$\mathcal{P}$ Cumulative proportion of neighbors of local centers to all points;
$\mathcal{S}_{k}$, $\mathcal{\hat{S}}_{k}$ Score pools for accepting and leaving points unlabelled in cluster $k$;
$\eta_{k}$, $\eta_{X}$ Depth similarity thresholds for cluster $k$ and the dataset ($\mathbb{S}$-based);
$\mathcal{I}_{k}$ Indices of points in $\mathbf{g}_{k}$;
$\mathcal{J}_{k}^{i}$ Indices of core neighbors of $\mathbf{x}_{i}$ in $\mathcal{X}_{k}$ based on $\eta_{k}$;
$\mathcal{H}_{k}^{i}$ Indices of core neighbors of $\mathbf{x}_{i}$ outside $\mathcal{X}_{k}$ based on $\eta_{X}$;
$\mathbf{I}_{k}$ Inner observations of cluster $k$;
$\mathbf{O}_{k}$ DBCA-outliers of cluster $k$;
II Depth-based Similarity Matrix
II-A Preliminaries
A depth function quantifies how “central” a point is within a multidimensional data distribution, providing a meaningful center-outward ordering of points. The term “depth” reflects the analogy to physical depth: points closer to the “central area” of a distribution are considered “deeper”, while points farther away are “shallower”. It is worth to note that a point with high depth is central but does not necessarily coincide with a region of high density. Consequently, clustering approaches based on depth are conceptually distinct from those based on density estimates. This centrality-based ordering is particularly useful in multivariate data analysis, where traditional univariate notions of order are not well-defined.
Serfling and Zuo [30] formulated several desirable properties of depth functions: affine invariance, maximality at center, monotonicity on rays, and vanishing at infinity. A deepest point $\mathbf{m}=\text{argmax}_{\mathbf{z}\in\mathbb{R}^{d}}D(\mathbf{x}\mid F)$ is usually referred to as a depth median. In applications, where the inherent distribution $F$ is usually unknown, depth is calculated with respect to the empirical distribution $F_{n}$ of $\mathbf{X}$, that is $D(\mathbf{z}\mid F_{n})$ is the depth of $\mathbf{z}$ with respect to the sample $\mathbf{X}$. Hereinafter, we will denote the latter as $D(\mathbf{z}\mid\mathbf{X})$ to emphasize the role of the sample on calculating the depth of the point $\mathbf{z}$. For a broader overview of various depth notions, see Mosler and Mozharovskyi [31].
Previous studies have explored the application of data depth in clustering. Jornsten [32] proposed DDclust, an algorithm that refines the output of other clustering methods like PAM by recalculating observation depths within each cluster and reallocating points for better results. Jeong et al. [33] introduced data depth-based clustering analysis (DBCA), an alternative to DBSCAN. DBCA replaces the sample Mahalanobis depth (MD) function
$D_{\mathrm{MD}}(\mathbf{z}\mid\mathbf{X})=\left[1+(\mathbf{z}-\bar{\mathbf{X}})^{\prime}\hat{\mathbf{\Sigma}}^{-1}(\mathbf{z}-\bar{\mathbf{X}})\right]^{-1}$ (1)
with fixed depth median versions
$D_{\mathrm{MD}}^{i}(\mathbf{z})=\left[1+(\mathbf{z}-\mathbf{x}_{i})^{\prime}\hat{\mathbf{\Sigma}}^{-1}(\mathbf{z}-\mathbf{x}_{i})\right]^{-1},$ (2)
where $\mathbf{z}$ denotes a point in $\mathbb{R}^{d}$, $\bar{\mathbf{X}}$ is the mean vector, $\mathbf{x}_{i}$ is an observation in $\mathbf{X}$, while $\hat{\mathbf{\Sigma}}$ is the estimated covariance matrix of dataset $\mathbf{X}$. Unlike DBSCAN, which uses fixed Euclidean distance for neighborhoods, DBCA employs (2) to define elliptical, affine-invariant neighborhoods. Recent work has also explored data depth to identify centroids for k-means initialization in diverse datasets [34, 35].
So far, the notion of depth serves an auxiliary role to the more traditional clustering methods and is not employed directly in examining the multimodality of data. Herein, two depth functions are considered for our DLCC algorithm: MD and spatial depth (SD). For SD [36], the sample version is
$D_{\mathrm{SD}}(\mathbf{z}\mid\mathbf{X})=1-\left\|\sum_{i=1}^{n}\frac{\mathbf{z}-\mathbf{x}_{i}}{n\|\mathbf{z}-\mathbf{x}_{i}\|}\right\|,$ (3)
where $n$ is the total number of observations. Although SD is only invariant under similarity transformations, it is nevertheless a sensible choice in the current context due to its straightforward computation in high-dimensional spaces.
DLCC requires a similarity matrix to determine the neighbors of each observation. Traditional distance measures, such as Euclidean distance, are typically based on assumptions about the underlying cluster shapes (e.g., spherical or linear structures), which may not hold in more complex scenarios. In contrast, data depth offers a nonparametric, robust, and adaptive alternative. As the depths (1) and (3) provide a center-outward ordering and their value ranges from 00 to $1$, the depth values of other points can be viewed as a measure of similarity to the depth median. Thus, similarity matrices can be constructed based on data depth. For this, however, one requires an approach capable of defining neighbors of any point $\mathbf{x}_{i}$ in $\mathbf{X}$, i.e., the observations in $\mathbf{X}$ that are most similar to $\mathbf{x}_{i}$. It should be clear that the usual notion of depth cannot be applied directly here, as that can only provide similarity measurements related to the global median.
Now, consider a point $\mathbf{x}_{i}$ in $\mathbf{X}$ whose neighbors need to be identified and introduce the set
$\mathbf{X}_{Ri}=\mathbf{X}\cup\{2\mathbf{x}_{i}-\mathbf{x}_{j}\,|\,j\neq i,j=1,2,\ldots,n\}.$
We note that this new set is comprised of all the points belonging to the original sample $\mathbf{X}$ as well as of the reflection of all the points of $\mathbf{X}$ through $\mathbf{x}_{i}$. Crucially, $\mathbf{x}_{i}$ is a depth median of $\mathbf{X}_{Ri}$ as it is a center of symmetry for this set. It is also the case that $D(\mathbf{x}_{i}|\mathbf{X}_{Ri})=1$ and that $D(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ can be used to quantify the similarity of any point $\mathbf{x}_{j}$ ($j\neq i$) to the point of interest $\mathbf{x}_{i}$. Subsequently, by repeatedly computing the depth values for points in $\mathbf{X}$ with respect to $\mathbf{X}_{R1},\mathbf{X}_{R2},\ldots,\mathbf{X}_{Rn}$, a $n\times n$ non-mutual depth-based similarity matrix can be defined. We denote it as $\mathbb{S}$, where $\mathbb{S}_{ij}=D(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ represents the entry in the $i$th row and $j$th column. As will be seen shortly, this can cause some computational challenges.
This approach to localize the concept of depth was introduced by Paindaveine and Van Bever [37] when they introduced local depth, which was shown to yield non-spherical depth neighborhoods in some contexts. The same authors [38] used this concept and introduced an affine-invariant k-nearest neighbor classification algorithm that adapts to the local geometry of the sample and was shown to be consistent.
II-B Mahalanobis Depth based Similarity Matrix
Previous studies [39] have employed the approach of using individual observations $\mathbf{x}_{i}$, $i=1,2,\ldots,n$, in place of the mean vector, while incorporating the globally estimated covariance matrix, for depth value computations as given in (2). In other words, the center used for calculating the Mahalanobis depth is shifted to each observation, and the depth values of the remaining observations are subsequently recorded. These studies commonly employ the minimum covariance determinant (MCD) method [40] to achieve robust results. In light of the previous discussion, the mean of $\mathbf{X}_{Ri}$ can easily be seen to be $\mathbf{x}_{i}$, which leads to
$D_{\mathrm{MD}}(\boldsymbol{z}\mid\mathbf{X}_{Ri})=D_{\mathrm{MD}}^{i}(\boldsymbol{z}),$
assuming the use of a consistent covariance matrix. This implies that making use of (2) to define a neighborhood of any point is in line with the approach of Paindaveine and Van Bever.
However, the method of Paindaveine and Van Bever necessitates estimating $n$ distinct covariance matrices, substantially increasing computational load. In contrast, the DBCA approach estimates only a single global covariance matrix, which may not adequately reflect the diversity within different clusters. To mediate between these extremes, we propose a covariance matrix selection strategy that balances computational efficiency with the ability to capture the unique characteristics of each cluster more effectively than a global approach.
In the framework of GMM, it is feasible to estimate multiple covariance matrices, ideally keeping the number of such matrices less than or equal to the number of clusters — substantially fewer than $n$ — yet more than one. The covariance matrix for cluster $k$, i.e., component $k$, can be decomposed as
${\mathbf{\Sigma}}_{k}=\lambda_{k}\mathbf{\Gamma}_{k}\mathbf{\Delta}_{k}\mathbf{\Gamma}_{k}^{\prime},$ (4)
where $\lambda_{k}=|{\mathbf{\Sigma}}_{k}|^{1/d}$, $\mathbf{\Gamma}_{k}$ denotes the matrix of eigenvectors, and $\mathbf{\Delta}_{k}$ is a diagonal matrix of eigenvalues with the condition $|\mathbf{\Delta}_{k}|=1$ [41], [42]. These elements, respectively, represent size, orientation, and shape of the cluster; see [43] for details.
In our study, we focus on the EEV model from the Gaussian parsimonious clustering models (GPCM) family, wherein different clusters estimated by the EEV model share common $\lambda_{k}$ and $\mathbf{\Delta}_{k}$ — i.e., $\lambda_{k}=\lambda$ and $\mathbf{\Delta}_{k}=\mathbf{\Delta}$ — but possess distinct $\mathbf{\Gamma}_{k}$. When model performance metrics, such as the Bayesian information criterion (BIC; [44]), are similar for results with varying numbers of components, those with fewer components are preferred. We do not consider more flexible models within the GPCM family as these could result in GMMs that overfit the data, and the size of the estimated covariance matrices may be too small for constructing the similarity matrix.
If we assume clusters are all convex, we use a test based on principal component analysis (PCA; [45]) to decide on the method for constructing the similarity matrix. With the close correlation between the Mahalanobis distance and PCA, the squared Mahalanobis distance is equal to the sum of squares of standardized principal component scores [46]. If a significant proportion of variance can be explained by the leading principal components, then the Mahalanobis distance in the space spanned by these components is a suitable metric, especially in the presence of strong inter-variable correlations. In contrast, if the variables show no clear correlations, the standard Euclidean distance is deemed a more appropriate measure.
The test procedure can be summarized as follows: Let $\mathbf{vp}=\{vp_{1},vp_{2},\ldots,vp_{H},\ldots,vp_{d}\}$, which contains the proportion of variance explained by each PC from high to low, where $H$ indicates the number of PCs that explain over $5\%$ variance, but if $H=d$, then $H$ redefined to be $d-1$. If either (5) or (6) is satisfied, the dataset passes the test:
$vp_{1}$ $>$ $0.6,$ (5)
$\sum_{h=1}^{H}vp_{h}$ $>$ $0.95$ (6)
When the dataset passes the test, we use the estimated global covariance matrix. If it fails, the EEV model is brought into play, offering the benefit of estimating clusters with identical eigenvalues. The test can be repeated using the eigenvalues derived from the EEV model. If it continues to fail, indicating weak inter-variable correlations even within clusters, and the Euclidean distance becomes sufficient for constructing the similarity matrix, setting the covariance matrix in (2) as the identity matrix. However, if it succeeds, suggesting the presence of within-cluster correlations, we employ the covariance matrices estimated by the EEV model. Specifically, consider two points, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, where $\mathbf{x}_{i}$ is assigned to cluster $1$ and $\mathbf{x}_{j}$ to cluster $2$ by the EEV model. The estimated covariance matrices, $\hat{\mathbf{\Sigma}}_{1}$ for cluster $1$ and $\hat{\mathbf{\Sigma}}_{2}$ for cluster $2$, are used to compute the $i$th and $j$th rows of the similarity matrix, respectively. While both covariance matrices are of the same size and shape, their orientations differ to capture the inter-variable correlation within each cluster.
II-C Spatial Depth based Similarity Matrix
A spatial-depth-based similarity matrix can be constructed from calculating $D_{\mathrm{SD}}(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ for all $i,j=1,2,\ldots,n$. One advantage of using this depth measure is that the shape of the resulting neighborhoods is determined by the local geometry of the sample and does not rely on any parameters. This is a potentially very interesting advantage, but it does entail the computationally intensive process of repeatedly calculating depth values for the $n$ sample points with respect to $n$ distinct datasets $\mathbf{X}_{R1},\mathbf{X}_{R2},\ldots,\mathbf{X}_{Rn}$. Hence, we delve into certain properties of spatial depth in the current context (working with reflected samples) and introduce an efficient approach to calculate the similarity matrix.
In order to compute the spatial depth of a point $\mathbf{x}$ from (3), we note that it is necessary to find the vectors from all sample points to $\mathbf{x}$ and their corresponding norms. In any of the combined dataset $\mathbf{X}_{Ri}$, the $n$ points from the original dataset $\mathbf{X}$ remain unchanged. Consequently, we can retain a $n\times d$ matrix, denoted as $\mathbf{E}$, to prevent redundant computations. Here, the $q^{\text{th}}$ row is the sum of the unit vectors from other observations in $\mathbf{X}$ to $\mathbf{x}_{q}$, expressed as:
$\mathbf{E}_{q}=\sum_{\mathbf{x}\in\mathbf{X}\setminus\{\mathbf{x}_{q}\}}\frac{\mathbf{x}_{q}-\mathbf{x}}{\|\mathbf{x}_{q}-\mathbf{x}\|}.$
Taking the combined dataset $\mathbf{X}_{Ri}$ as an example, the subsequent step is to derive a similar $n\times d$ matrix $\mathbf{\tilde{E}}_{i}$. In this matrix, the $q$th row is represented by
$\mathbf{\tilde{E}}_{i,q}=\sum_{\mathbf{x}\in\mathbf{X}_{Ri}\setminus\mathbf{X}}\frac{\mathbf{x}_{q}-\mathbf{x}}{\|\mathbf{x}_{q}-\mathbf{x}\|},$
and corresponds to the the sum of the unit vectors from all the reflected points in $\mathbf{X}_{Ri}$ to $\mathbf{x}_{q}$.
Proposition 1 (Symmetry).
Consider two distinct points $\mathbf{u}$ and $\mathbf{v}$, and a point $\mathbf{x}_{i}\in\mathbf{X}$. Also, let $\mathbf{u}^{i}$ and $\mathbf{v}^{i}$ denote the points obtained by reflecting $\mathbf{u}$ and $\mathbf{v}$, respectively, about $\mathbf{x}_{i}$, that is $\mathbf{u}^{i}=2\mathbf{x}_{i}-\mathbf{u}$ and $\mathbf{v}^{i}=2\mathbf{x}_{i}-\mathbf{v}$. Then, we have that
$\mathbf{u}-\mathbf{v}^{i}=\mathbf{v}-\mathbf{u}^{i},$
and that
$\lVert\mathbf{v}^{i}-\mathbf{u}\rVert^{2}=2\lVert\mathbf{v}-\mathbf{x}_{i}\rVert^{2}+2\lVert\mathbf{u}-\mathbf{x}_{i}\rVert^{2}-\lVert\mathbf{u}-\mathbf{v}\rVert^{2}.$
The above result implies that we can express the squared norm of a vector from any point after reflection to an arbitrary point in $\mathbf{X}$, as a linear combination of the squared norms of vectors linking points in $\mathbf{X}$. Therefore, if the norm information is concurrently stored while generating the matrix $\mathbf{E}$, the norm of any vector from the reflected point to a point in $\mathbf{X}$ can be directly computed. Let us denote the $n\times n$ square norm matrix as $\mathbf{L}$, where the entry in position $(i,j)$ stores the norm of the vector between point $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$. This computation no longer involves calculating the root of the sum of squares of the vector coordinates, thus making the norm computation independent of the dimension $d$.
From Proposition 1, it can also be observed that the vector from any reflected point $\mathbf{v}^{i}$ to a vector $\mathbf{u}$ can be expressed as $\mathbf{u}+\mathbf{v}-2\mathbf{x}_{i}$. Then, the sum of unit vectors from all reflection points in $\mathbf{X}_{Ri}$ to $\mathbf{x}_{q}$ can be reformulated as follows:
$\begin{split}\sum_{\mathbf{x}\in\mathbf{X}_{Ri}\setminus\mathbf{X}}&\frac{\mathbf{x}_{q}-\mathbf{x}}{C_{\mathbf{x}}}=\sum_{\mathbf{x}\in\mathbf{X}\setminus\{\mathbf{x}_{i}\}}\frac{\mathbf{x}_{q}+\mathbf{x}-2\mathbf{x}_{i}}{C_{2\mathbf{x}_{i}-\mathbf{x}}}\\ &\quad=\mathbf{x}_{q}\sum_{\mathbf{x}\,\in\,\mathbf{X}\setminus\{\mathbf{x}_{i}\}}\frac{1}{C_{2\mathbf{x}_{i}-x}}\ \ +\sum_{\mathbf{x}\,\in\,\mathbf{X}\setminus\{x_{i}\}}\frac{\mathbf{x}-2\mathbf{x}_{i}}{C_{2\mathbf{x}_{i}-\mathbf{x}}},\end{split}$
where $C_{\mathbf{x}}=\|\mathbf{x}_{q}-\mathbf{x}\|$ and, by separating $\mathbf{x}_{q}$ from the summation, we save some vector addition computations.
The matrix $\mathbf{\tilde{E}}_{i}$ can now be constructed using matrix computations. The necessary notations to understand this process are summarized below:
  • The $(n-1)\times n$ matrix $\mathbf{\tilde{L}}_{i}$ stores the norms of vectors from $\mathbf{x}^{i}_{j}$ to $\boldsymbol{x_{q}}$ at each entry $\tilde{L}_{i,jq}$.
  • The $(n-1)\times n$ matrix $\mathbf{R}_{i}$ is obtained by taking the element-wise reciprocal of $\mathbf{\tilde{L}}_{i}$, i.e., $R_{i,jq}=1/\tilde{L}_{i,jq}$.
  • The $(n-1)\times d$ matrix $\mathbf{\tilde{X}}_{i}$ contains each reflected point (excluding $\mathbf{x}_{i}$ itself) as a row.
  • The $n\times d$ matrix $\mathbf{B}_{i}$ consists of points in $\mathbf{X}$, each multiplied by the corresponding column sum of $\mathbf{R}_{i}$.
With these, $\mathbf{\tilde{E}}_{i}$ is expressed as: $\mathbf{\tilde{E}}_{i}=\mathbf{B}_{i}-\mathbf{R}^{\prime}_{i}\mathbf{\tilde{X}}_{i}.$ Through the above procedure, $\mathbf{\tilde{E}}_{i}$ can be derived efficiently. For instance, in a $2000\times 600$ dataset, while $\mathbf{E}$ requires approximately two seconds to compute, $\mathbf{\tilde{E}}_{i}$ only takes around $0.1$ second per loop — run on a PC with Intel(R) Core(TM) i7-11700K processor (3.6 GHz) with sufficiently large physical memory, Matlab R2023a. Having obtained $\mathbf{E}$ and $\mathbf{\tilde{E}}$, the reflection spatial depth values of any $\mathbf{x}_{q}$ in $\mathbf{X}$ based on $\mathbf{X}_{Ri}$ can be computed via
$D_{\mathrm{SD}}(\mathbf{x}_{q}|\mathbf{X}_{Ri})=1-\left\|\frac{\mathbf{E}_{q}+\mathbf{\tilde{E}}_{i,q}}{2n-1}\right\|.$ (7)
III Depth-based Local Center Clustering
In this section, we first introduce some concepts that are key to DLCC and then provide a detailed explanation of the clustering procedure.
As explained in Section II, it is necessary to use an appropriately localized version of data depth for clustering applications. In the framework of DLCC, we establish two foundational concepts: neighborhoods and local ranks. These concepts are formalized in the following definition:
Definition 1 (DLCC Neighborhoods and Local Ranks).
Neighbors and Neighborhoods: Given an integer parameter $s$, the neighbors of $\mathbf{x}_{i}\in\mathbf{X}$ are defined as the set of $s-1$ points $\mathbf{x}_{q}\in\mathbf{X}\setminus\{\mathbf{x}_{i}\}$ with the highest depth values with respect to the dataset $\mathbf{X}_{Ri}$, i.e., the points with the largest $D\left(\mathbf{x}_{q}\mid\mathbf{X}_{Ri}\right)$. The neighborhood of $\mathbf{x}_{i}$, denoted as $\mathbf{N}_{i}$, is thus formed by including $\mathbf{x}_{i}$ and its $s-1$ neighbors, constituting a subset of $\mathbf{X}$ with size $s$.
Local Depth [37] and Local Ranks: The local depth of an observation $\mathbf{x}_{i}$ is its depth value within its neighborhood $\mathbf{N}_{i}$. The local rank of an observation $\mathbf{x}_{j}$ within the neighborhood $\mathbf{N}_{i}$ is defined as:
$r_{i}(\mathbf{x}_{j})=1+\sum_{\mathbf{x}_{q}\in\mathbf{N}_{i}}\mathbb{I}\left[D\left(\mathbf{x}_{j}\mid\mathbf{N}_{i}\right)<D\left(\mathbf{x}_{q}\mid\mathbf{N}_{i}\right)\right],$ (8)
where $\mathbb{I}$ denotes the indicator function.
If $r_{i}(\mathbf{x}_{j})=1$, then $\mathbf{x}_{j}$ is characterized as the center of the subset $\mathbf{N}_{i}$. We define central points of $\mathbf{N}_{i}$ as a local center. These local centers can be understood as representative points for the entire dataset $\mathbf{X}$.
Figure 3: Local center example ($s=50$) for the first two PCs of the Iris dataset, where a focus point (green triangle) with its neighbors (pink points) and the corresponding local center (blue diamond) are highlighted.
Notes:
  1. 1.
    While $\mathbf{x}_{i}$ is deepest in $\mathbf{X}_{Ri}$ by construction, it is not generally deepest in $\mathbf{N}_{i}$, as illustrated in Figure 3.
  2. 2.
    There should be less than $n$ unique local centers due to the occurrence of shared local centers among the sets $\mathbf{N}_{i}$. The frequency of a local center, denoted by $f$, is defined as the number of neighborhoods that share this local center.
The DLCC method uncovers clusters through analyzing and grouping local centers. However, as $n$ subsets might yield a large number of local centers, many of them might not be useful for clustering. For example, an outlier residing between two clusters can be the deepest point in a subset containing observations from both clusters. Hence, it is crucial to filter out unnecessary local centers before grouping them. The remaining centers after filtering are denoted as filtered centers.
DLCC introduces two strategies designed to address diverse types of clustering challenges. The conditions for grouping the filtered centers under these two strategies are summarized as follows:
Definition 2.
Let ${c_{1},\ldots,c_{A}}$ denote a finite set of filtered centers and ${\mathbf{N}_{c_{1}},\ldots,\mathbf{N}_{c_{A}}}$ their corresponding neighborhoods. Define the similarity between any two neighborhoods as the proportion of overlapping observations among them. Two grouping strategies are as follows:
  • Min strategy: for any $c\in\mathbf{g}_{i}$:
[…]
  • Max strategy: Given a parameter $\delta$, termed as the similarity threshold, $\forall$ $c$ in group $\mathbf{g}$, $\exists c^{*}$ in the same group s.t. $\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{*}})\,>\delta$, and for $\forall c^{\dagger}\notin\mathbf{g}$, $\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}})\,\leq\delta$.
Note that, “group” refers specifically to a set of filtered centers and should not be confused with “cluster”. Some insights into the strategy chosen for grouping local centers are important in applications. Under the min strategy, DLCC attempts to find centroids of clusters from filtered centers. The variance in the number of filtered centers across different groups allows for better flexibility than traditional center-based clustering methods, especially when dealing with clustering problems that have overlapping regions. The min strategy tends to be more suitable when the sizes of all clusters are comparable.
The max strategy assumes each cluster can be built by connecting multiple finite convex shapes created by filtered centers. Filtered centers in the same group are connected with a chain of similarities. For example, if two filtered centers $c_{y}$ and $c_{z}$ are in the same group, then there must exist a chain of filtered centers, $c_{p_{1}},c_{p_{2}},\ldots,c_{p_{h}}$, where $c_{p_{1}}=c_{y}$, $c_{p_{h}}=c_{z}$, and the similarity between any two adjacent centers’ neighborhoods in that chain is larger than $\delta$. Generally, the parameter $s$ is small under the max strategy. When using Mahalanobis depth to construct the similarity matrix with a small $s$, the max strategy is less sensitive to variations in the matrix. In such cases, a globally estimated similarity matrix often suffices, similar to DBCA. However, unlike DBCA, which builds clusters with chains of fixed-shape ellipses centered at each observation, DLCC with the max strategy considers only chains of filtered centers, with a fixed number of neighbors rather than a fixed region. Like density-based methods, the max strategy works well for disjoint clusters and can handle non-convex clusters of varying sizes.
Definition 3 (Unique neighbor).
Define $\bigcup_{c_{i}\in\mathbf{g}}\mathbf{N}_{c_{i}}$ and $\bigcup_{c_{j}\notin\mathbf{g}}\mathbf{N}_{c_{j}}$ as the unions of neighborhoods for filtered centers within and outside of group $\mathbf{g}$, respectively. The set of unique neighbors for $\mathbf{g}$, $\mathbf{U}_{\mathbf{g}}$, is the largest subset of $\bigcup_{c_{i}\in\mathbf{g}}\mathbf{N}_{c_{i}}$ satisfying:
$\mathbf{U}_{\mathbf{g}}\cap\bigcup_{c_{j}\notin\mathbf{g}}\mathbf{N}_{c_{j}}=\emptyset.$
The number of clusters, denoted by $K$, aligns with the grouping of the filtered centers. Temporary clusters are derived from the unique neighbors of each group. The details of constructing these temporary clusters based on the unique neighbors and filtered centers will be further elaborated later in this section.
With these temporary clusters, it is expected that most observations have been allocated to a particular cluster. Nevertheless, there may still be some points that remain unclustered. In such cases, the clustering problem can be reframed as a classification problem. Intuitively, classification methods can be used to assign the remaining points to a cluster. It should be noted that the choice of classification method is very flexible here. Indeed, we could potentially consider any classification method at this stage, depending on the specific requirements of the task at hand. DLCC incorporates three built-in classification methods: maximum-depth classifier (MDC), K-nearest neighbors (KNN) based on the similarity matrix, and random forests (RF).
Lastly, self-improvement procedures of DLCC are discussed. Although the clustering problem is converted to a classification problem in the final stage, it has some special features, because the observations that are already labelled may still be updated. Two logical parameters are introduced here:
  • maxdepth: This parameter dictates whether DLCC will continue iterations until all observations in the assigned cluster attain the highest depth values.
  • ifloop: This parameter, only applicable under the min strategy, determines whether only a single filtered center should remain in each cluster.
Let the $K$ clusters constructed so far be denoted by $\mathcal{X}_{k}$, where $k=1,2,\ldots,K$. For any observation $\mathbf{x}_{j}$, $D(\mathbf{x}_{j}\mid\mathcal{X}_{k})$ corresponds to the depth of this observation within cluster $\mathcal{X}_{k}$. If the depth value of $\mathbf{x}_{j}$ in the cluster it is currently assigned to does not match $\max_{k}D(\mathbf{x}_{j}\mid\mathcal{X}_{k})$, and if the “maxdepth” parameter is set to TRUE, $\mathbf{x}_{j}$ will be relocated. This procedure loops until all observations are assigned to the cluster where they obtain the largest depth value. Regarding the min strategy, it can be adapted to resemble a centroid-based clustering method. If the “ifloop” parameter is TRUE, the clustering algorithm updates clusters using only the deepest filtered center within each cluster. However, other filtered centers are not disregarded. They may become the deepest center for a cluster in future iterations, as clustering results are reassessed after each update. This iterative process continues, regenerating clusters and potentially re-evaluating the roles of filtered centers, until the clustering outcome stabilizes.
III-A Local Center Selection and Grouping
Locating and filtering local centers is at the core of DLCC. For neighborhoods $\mathbf{N}_{1},\mathbf{N}_{2},\ldots,\mathbf{N}_{n}$ of size $s$, local centers are identified. To filter local centers, two benchmarks are suggested:
  • a local center should be central to its neighborhood; and
  • a local center with higher frequency $f$ is more reliable.
The first step of filtering local centers is to pick up local centers satisfying $r_{i}(\mathbf{x}_{i})\leq 2$ (Definition 1), which aims to meet the first benchmark. The requirement of a central position is relaxed from $r_{i}(\mathbf{x}_{i})=1$ to $r_{i}(\mathbf{x}_{i})\leq 2$ for adding flexibility. Local centers with $f=1$ are discarded next, as they are less reliable. The remaining steps in filtering local centers differ between the min and max strategies.
III-A1 Min strategy
After the first filtering step, assume there are $T$ local centers remaining. Index these local centers in the decreasing order of their frequencies, i.e., $\{c_{1},\ldots,c_{T}\},$ which corresponds to $f_{1}\geq f_{2}\geq\cdots\geq f_{T}$. With this order, the cumulative proportion of neighbors for the local centers down the list can be calculated as
$\mathcal{P}_{t}=\frac{1}{n}\left|\bigcup_{p=1}^{t}\mathbf{N}_{c_{p}}\right|,$ (11)
where ||$\cdot$|| ⋅ | represents the cardinality of a set.
The filtering procedure consists of two steps: (1) selecting a suitable frequency cutoff, and (2) adjusting the grouping results to satisfy the min strategy defined in Definition 2. According to Definition 2, for fixed local centers, the grouping result satisfying the max strategy is unique, whereas multiple solutions may satisfy the min strategy. Thus, a reasonable initial grouping method that provides high-quality results and estimates the number of groups is essential. We summarize the designed grouping method in Algorithm 1. This method leverages local depths and the similarities between neighborhoods of local centers. Here, local depth values are computed during defining local centers, and the similarity matrix for neighborhoods of local centers is denoted as $\mathbb{M}$, with $\mathrm{sim}(\mathbf{N}_{c_{i}},\mathbf{N}_{c_{j}})$ as its $ij$-th entry.
Algorithm 1 Local depth and similarity-based grouping
1:Local centers $c_{1},\ldots,c_{T}$, corresponding top $s/2$ neighbors, local depth values (LD), similarity matrix $\mathbb{M}$, and a similarity threshold range $[0.4,0.75]$.
2:Grouping results $\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{G}\}$.
3:For each local center $c_{i}$, set valid neighbors $\mathrm{VN}_{i}$ as local centers in its top $s/2$ neighbors (excluding itself).
4:Initialize mean similarities and closest neighbors: MS=zero array of size $\mathrm{MS}=\text{zero array of size }T$, $\mathrm{CN}=1{:}T$
5:for $t=1$ to $T$ do
6:     Compute mean similarity: $\mathrm{MS}_{t}=\frac{\sum_{c_{j}\in\mathrm{VN}_{t}}\mathbb{M}_{tj}}{|\mathrm{VN}_{t}|}$.
7:     Find the closest neighbor: $\mathrm{CN}_{t}=\arg\max_{j\neq t}\mathbb{M}_{tj}$.
8:end for
9:Set similarity threshold $\delta_{i}$ for each local center $c_{i}$:
$\delta_{i}=\min(\mathrm{MS}_{i},\mathrm{MS}_{\mathrm{CN}_{i}})$
and if $\delta_{i}\notin[0.4,0.75]$ adjust it to the nearest bound.
10:Initialize stable centers $\mathrm{SC}=\emptyset$.
11:for $t=1$ to $T$ do
12:     if All $c_{j}\in\mathrm{VN}_{t}$ with $\mathrm{sim}(\mathbf{N}_{c_{t}},\mathbf{N}_{c})\geq\delta_{t}$ have $\text{LD}_{j}<\text{LD}_{t}$ then
13:         Add $c_{t}$ to $\mathrm{SC}$.
14:     end if
15:end for
16:Set $G=|\mathrm{SC}|$.
17:Assign each remaining local center to the group of its most similar stable center using $\mathbb{M}$.
18:Return groups $\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{G}\}$.
Algorithm 1 uses local depth to assess the centrality of each center and employs neighborhood similarity to evaluate the relationship between centers. The similarity threshold is determined automatically, eliminating the need for manual parameter selection. Specifically, the threshold for each center $c$ is set as the minimum of the mean similarity between $c$ and its closest neighbor. This approach provides a center-specific threshold while ensuring that centers with similar neighborhoods share comparable thresholds. The empirical threshold range $[0.4,0.75]$ prevents the values from being excessively small or large. Moreover, the algorithm naturally estimates the number of groups, which corresponds to the number of estimated clusters.
A limitation of Algorithm 1 is if multiple clusters overlap with each other, the local center defined in the overlapped region may have high local depth values. These local centers can be removed by analyzing the number of unique neighbors (Definition 3) within each group, as groups containing local centers from overlapped regions tend to have very few unique neighbors. Specifically, we consider a group of centers should be dropped if the number of unique neighbors of this group is less than $s/10$. Additionally, local centers highly similar to those in the identified group should also be dropped to avoid ambiguity.
With the grouping of local centers, we determine candidates for the cutoff in the order of frequency. The first candidate index is defined as
$\max\left(\arg\min_{t}\{\mathcal{P}_{t}\geq 0.75\},\max_{\mathbf{g}}\{\arg\min_{t:c_{t}\in\mathbf{g}}t\}\right),$ (12)
where the first term ensures the final cumulative proportion $\mathcal{P}_{t}$ exceeds $0.75$ to prevent excessive elimination of local centers, and the second term guarantees at least one local center from each group is retained.
Additional candidates include local centers after the begin index that provide complementary information beyond those in the same group with higher frequency. Specifically, a candidate index $i$ must satisfy two conditions: First for all $c_{j}$ where $j\leq i$ and $c_{j}$ belongs to the same group $\mathbf{g}$ as $c_{i}$,
$\sum_{j}\mathbb{M}_{ij}=\min_{l\leq i,c_{l}\in\mathbf{g}}\sum_{j}\mathbb{M}_{lj},$ (13)
This ensures $c_{i}$ is the least similar to other local centers with higher frequencies in its group. Second, for all $c_{v}$ where $v<i$ and $c_{v}\notin\mathbf{g}$,
$\max_{v}\mathbb{M}_{iv}\leq\min_{j<i,c_{j}\in\mathbf{g}}\max_{v}\mathbb{M}_{jv},$ (14)
ensuring $c_{i}$ is less similar to points in other groups compared to any existing local center in its own group. In essence, these candidates are distant from local centers in other groups while also comparatively distinct within their own group, making them potential sources of complementary information.
Lastly, we evaluate each candidate in ascending order and count the number of “bad points” between candidates, defined as local centers that fail to satisfy Equations (9) or (10) after removing local centers in groups with indices larger than the candidate. If the number of bad points between two consecutive candidates $c_{i}$ and $c_{i^{\prime}}$ exceeds $1$, we stop at $c_{i}$.
In the above setting, the initial candidate should have a cumulative proportion of at least $0.75$. However, if $\mathcal{P}_{T}<0.75$, it may indicate that overlapping clusters cause local centers to concentrate in shared regions rather than capturing true subcluster centers. The robustness of the depth median (e.g. SD’s breakdown point is $0.5$) biases local centers toward overlapping regions when most points in a cluster have neighborhoods with similarities exceeding $0.5$ to points in other clusters. To address this, if $\mathcal{P}_{T}<0.75$, we retain only centers dissimilar to those in other groups after grouping $c_{1},\ldots,c_{T}$ using Algorithm 1. For $c\in\mathbf{g}$, we check:
$\frac{\sum_{j:c_{j}\in\mathbf{g}\setminus c_{i}}\mathbb{M}_{ij}}{|\mathbf{g}|-1}-\max_{v:c_{v}\notin\mathbf{g}}\mathbb{M}_{iv}>0.$ (15)
Then, we iteratively add back local centers from the dropped ones, including those with $f=1$ or $r_{i}(\mathbf{x}_{i})>2$. In each step, the local center that maximizes the increase in cumulative proportion is added back. The process stops when the cumulative proportion exceeds $90\%$ of the cumulative proportion of all local centers.
Algorithm 2 Local center filtering under the min strategy
1:All local centers, their ranks, frequencies, local depth values, neighborhoods, size $s$ and similarity matrix $\mathbb{M}$.
2:Filtered local centers and their local depth values.
3:Perform initial filtering to obtain $\{c_{1},\ldots,c_{T}\}$.
4:if $\mathcal{P}_{T}\geq 0.75$ then
5:     Apply Algorithm 1 to obtain $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$.
6:     Set Continue as True.
7:     while Continue do
8:         Obtain unique neighbors $\{\mathbf{U}_{\mathbf{g}_{1}},\ldots,\mathbf{U}_{\mathbf{g}_{G}}\}$.
9:         Identify $\mathrm{di}=\arg\min_{i:|\mathbf{U}_{\mathbf{g}_{i}}|<s/10}|\mathbf{U}_{\mathbf{g}_{i}}|$.
10:         if $\mathrm{di}\neq\emptyset$ then
11:              Set $\delta=\max_{i,j\in\mathbf{g}_{\mathrm{di}},i\neq j}\mathbb{M}_{ij}$.
12:              Mark centers in $\mathbf{g}_{\mathrm{di}}$ as drop centers.
13:              Add centers $v$ outside $\mathbf{g}_{\mathrm{di}}$ to drop centers if $\max_{i:c_{i}\in\mathbf{g}_{\mathrm{di}}}\mathbb{M}_{vi}\geq\delta$.
14:              Remove $\mathbf{g}_{\mathrm{di}}$ and drop centers, and update $G$.
15:         else
16:              Set Continue as False;
17:         end if
18:     end while
19:     Define the first cutoff candidate $\mathrm{cc}_{1}$ as in (12).
20:     Identify additional candidates among centers with indices greater than $\mathrm{cc}_{1}$ and satisfy (13) and (14).
21:     Initialize bp=zero array of size |$\mathrm{bp}=\text{zero array of size }|\mathrm{cc}|$.
22:     for $q=1$ to $|\mathrm{cc}|$ do
23:         Remove local centers in groups with indices larger than $\mathrm{cc}_{q}$ to form $\{\mathbf{g}^{\prime}_{1},\ldots,\mathbf{g}^{\prime}_{G}\}$.
[…]
25:     end for
26:     Define cutoff as $\mathrm{cc}_{p}$ where $p$ is the first index satisfying $\mathrm{bp}_{p+1}-\mathrm{bp}_{p}>1$. If no such $p$ exists, set $cc_{p}$ as the largest cutoff candidate index.
27:else
28:     Drop local centers from $\{c_{1},\ldots,c_{T}\}$ that do not satisfy (15).
29:     Compute current cumulative proportion $\mathcal{P}$ and total cumulative proportion $\mathcal{P}^{\prime}$.
30:     while $\mathcal{P}<0.9\mathcal{P}^{\prime}$ do
31:         Add the local center maximizes the increase in $\mathcal{P}$.
32:         Update $\mathcal{P}$.
33:     end while
34:end if
35:Save remaining local centers and their local depth values.
After Algorithm 2, we reapply Algorithm 1 to group the filtered centers. If any local centers fail to meet the requirements of the min strategy, Algorithm 3 is applied, which is referred to as the group trimming process. The number of clusters is not required, but if users specify $K$, the algorithm forces the number of groups to equal $K$. For simplicity, the left and right sides of inequality (9) are denoted as $\mathrm{mw}(c)$ and $\mathrm{mb}(c)$ in the algorithm, respectively.
Algorithm 3 Group Trimming Process
1:Groups of local centers $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$, similarity matrix $\mathbb{M}$, optional $K$.
2:Trimmed groups satisfying the min strategy.
3:if $K\neq\emptyset$ then
4:     if $G>K$ then
5:         Drop groups with the highest $\{\arg\min_{t:c_{t}\in\mathbf{g}}t\}$ until $G$ equals to $K$.
6:     else
7:         while $G<K$ do
8:              Define a new stable center and increment $G$ by $1$triangleright Details provided in supplementary material
9:         end while
10:         Update $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$.
11:     end if
12:end if
13:while violations of (9) or (10) exist do
14:     while violations of (9) exist do
15:         for each group ${\mathbf{g}}_{l}$ with violations of (9do
16:              Drop the center $c\in{\mathbf{g}}_{l}$ which maximizes:
$\frac{\sum_{c^{\prime}\in{\mathbf{g}}_{l}\setminus\{c\}}\left(\mathrm{mw}(c^{\prime})-\mathrm{mb}(c^{\prime})\right)}{|{\mathbf{g}}_{l}\setminus\{c\}|}.$
17:         end for
18:         Update groups and recompute violations of (9).
19:     end while
20:     if violations of (10) exist then
21:         Drop all centers violating (10).
22:         Update groups and recompute violations.
23:     end if
24:end while
III-A2 Max strategy
Definition 4 (Isolated Local Center).
A local center $c_{i}$ is considered isolated if, for all $c_{j}\neq c_{i}$, the similarity between the neighborhoods associated with $c_{i}$ and $c_{j}$ is zero, i.e.,
$\mathbb{M}_{ij}=0.$
In other words, an isolated local center is one whose neighborhood does not overlap with any other neighborhoods. As the max strategy assumes each cluster consists of various convex shapes, if there are any isolated local centers, they will be discarded.
Definition 5 (Everywhere Non-Convex Cluster).
An everywhere non-convex cluster is characterized by all of its points having non-convex neighborhoods. Examples include one-dimensional manifolds, such as spirals, and self-crossing curves. A region containing such a cluster is termed a locally non-convex region.
The max strategy fails to identify everywhere non-convex clusters because as the depth functions used herein only suit convex shapes, preventing recognition of local centers satisfying $r_{i}(\mathbf{x}_{i})\leq 2$ in a non-convex neighborhood. To assess the presence of everywhere non-convex clusters and mitigate their impact, we categorize local centers into three sets based on the outcome of prior filtering:
  • Set $\mathbf{c}_{1}$ includes all local centers that remain after initial filtering, characterized by non-isolated local centers with $r_{i}(\mathbf{x}_{i})\leq 2$ and $f\neq 1$.
  • Set $\mathbf{c}_{2}$ contains non-isolated local centers $c_{v}$ for which $\forall c_{i}\in\mathbf{c}_{1}$, $c_{v}$ satisfies $\mathbb{M}_{vi}=0$.
  • Set $\mathbf{c}_{3}$ consists of local centers $c_{v}\notin\mathbf{c}_{1}$ satisfying $\exists c_{i}\in\mathbf{c}_{1}$ such that $\mathbb{M}_{vi}>0$.
The presence of everywhere non-convex clusters is inferred if $\mathbf{c}_{2}$ is not empty. In such cases, Algorithm 4 are taken to segregate locally non-convex regions from other clusters.
Algorithm 4 Segregating the locally non-convex region
1:Sets of local centers $\mathbf{c}_{1}$, $\mathbf{c}_{2}$, and $\mathbf{c}_{3}$, neighborhood similarity matrix $\mathbb{M}$
2:Set of filtered centers $\mathbf{C}$
3:if $\mathbf{c}_{2}\neq\emptyset$ then
4:     for $v$ in the indices of local centers in $\mathbf{c}_{3}$ do
5:         if  $\max_{j:c_{j}\in\mathbf{c}_{2}}\mathbb{M}_{vj}>\max_{i:c_{i}\in\mathbf{c}_{1}}\mathbb{M}_{vi}$  then
6:              $\mathbf{c}_{2}=\mathbf{c}_{2}\cup\{c_{v}\}$
7:              for $i$ in the indices of local centers in $\mathbf{c}_{1}$ do
8:                  if $\mathbb{M}_{vi}>0$ then
9:                       $\mathbf{c}_{1}=\mathbf{c}_{1}\setminus\{c_{i}\}$
10:                  end if
11:              end for
12:         end if
13:     end for
14:     $\mathbf{C}=\mathbf{c}_{1}\cup\mathbf{c}_{2}$
15:     if any isolated local centers exist based on $\mathbf{C}$ then
16:         Remove isolated local centers from $\mathbf{C}$
17:     end if
18:else
19:     $\mathbf{C}=\mathbf{c}_{1}$
20:end if
Algorithm 4 ensures that local centers in $\mathbf{c}_{2}$, which capture the essence of locally non-convex regions, remain disjoint with the local centers in other clusters, i.e., local centers in $\mathbf{c}_{1}$.
III-B Clustering in DLCC
Assuming the grouping results for the filtered centers are at hand, with the unique neighbors of each group serving to initialize temporary clusters, let $\mathbf{C}$ represent the set of filtered centers, such that $\mathbf{C}=\bigcup_{k=1}^{G}\mathbf{g}_{k}$, where $G$ denotes the total number of groups, and $\mathbf{g}_{k}$ contains the filtered centers of the $k$th group. Consequently, the number of clusters $K$ is inherently defined to be equal to $G$, as each group’s unique neighbors directly initiate a corresponding temporary cluster. The subsequent step involves defining a score for each observation with respect to each cluster in order to update these temporary clusters.
III-B1 Min strategy
For the min strategy, the score accounts for both clustering and assessing if each observation’s clustering result is acceptable, and it is computed using the similarity matrix $\mathbb{S}$. If $\mathbb{S}$ is asymmetric, it is updated by replacing $\mathbb{S}_{ji}$ and $\mathbb{S}_{ij}$ with their average value. For clarity, we continue to denote the symmetrized matrix as $\mathbb{S}$, ensuring that $\mathbb{S}_{ij}=\mathbb{S}_{ji}$ in this and the subsequent discussion about the DC metric in Section IV. We define the score function for an observation $\mathbf{x}_{i}$ with respect to a temporary cluster $k$ under the min strategy as
$\operatorname{score}_{i|k}=\frac{\max_{j\in\mathcal{I}_{k}}\mathbb{S}_{ji}-\max_{j\in\bigcup_{z\neq k}\mathrm{I_{z}}}\mathbb{S}_{ji}}{\max\{\max_{j\in\mathcal{I}_{k}}\mathbb{S}_{ji},\max_{j\in\bigcup_{z\neq k}\mathrm{I_{z}}}\mathbb{S}_{ji}\}\,},$ (16)
where $\mathcal{I}_{k}$ is the set of indices for filtered centers within the $k$th group, i.e., $\mathcal{I}_{k}=\{j:\mathbf{x}_{j}\in\mathbf{g}_{k}\}$. The score (16) takes a value between $-1$ and $1$ and a positive value indicates that, compared with the nearest filtered center in other groups, the similarity between $\mathbf{x}_{i}$ and the nearest filtered center in group $k$ is larger.
For each cluster, two score pools are generated: $\mathcal{S}_{k}$, containing both the unique neighbors within the cluster ($\mathbf{U}_{\mathbf{g}_{k}}$) and their corresponding scores, and $\mathcal{\hat{S}}_{k}$, holding observations not in $\mathbf{U}_{\mathbf{g}_{k}}$ yet having positive scores toward this cluster, along with their scores. Upon reviewing these pools, observations with high scores may shift from $\mathcal{\hat{S}}_{k}$ to $\mathcal{S}_{k}$, and vice versa, for those with low scores in $\mathcal{S}_{k}$. Lastly, $\mathcal{S}_{k}$ includes observations confidently assigned to the cluster, while those in $\mathcal{\hat{S}}_{k}$ remain unlabelled. Note that during the reassignment of observations between pools, their corresponding scores are also moved. The detailed steps are outlined as Algorithm 5, where the second element of the min function in line $14$ ensures at least $s/2$ observations per temporary cluster, preventing overly small clusters. Figure 4 illustrates an example using the Yale B face recognition dataset [56]. Algorithm 5 removes labels from unique neighbors within each group that are less distinguishable from other points in their own group. Meanwhile, it assigns labels to points that are not unique neighbors but are relatively closer to the filtered centers of their group.
Figure 4: t-SNE visualization [47] of the Yale B dataset. Colors represent the true labels of the points, with only those in temporary clusters being colored. Points enclosed by diamond-shaped boxes represent filtered centers. (a) Initial points in $\mathcal{S}_{k}$, with points enclosed by square boxes indicating those to be moved to $\mathcal{\hat{S}}_{k}$. (b) The temporary clusters obtained after applying Algorithm 5.
Algorithm 5 Update Temporary Cluster
1:Score pools $\mathcal{S}_{k}$ and $\hat{\mathcal{S}}_{k}$ for $k=1,\ldots,K$, where $\mathcal{S}_{k}^{1}$ and $\hat{\mathcal{S}}_{k}^{1}$ represent observations, and $\mathcal{S}_{k}^{2}$ and $\hat{\mathcal{S}}_{k}^{2}$ represent the corresponding scores, neighborhood size $s$.
2:Updated score pool $\mathcal{S}_{k}$, where $\mathcal{S}_{k}^{1}$ constructs the $k$-th temporary cluster.
3:for $k=1$ to $K$ do
4:     $\gamma_{1}=\text{mean}(\hat{\mathcal{S}}_{k}^{2})$
5:     for $i$ in $\mathcal{S}_{k}^{2}$ do
6:         if $i<\gamma_{1}$ then
7:              $\mathcal{S}_{k}^{2}=\mathcal{S}_{k}^{2}\setminus\{i\}$
8:              $\hat{\mathcal{S}}_{k}^{2}=\hat{\mathcal{S}}_{k}^{2}\cup\{i\}$
9:              Move corresponding points from $\mathcal{S}_{k}^{1}$ to $\hat{\mathcal{S}}_{k}^{1}$
10:         end if
11:     end for
12:     $SS=\text{Sort scores in }\mathcal{S}_{k}^{2}\text{ in descending order}$
13:     $B=\text{length of }SS$
14:     $idx=\arg\max_{j=\text{floor}(B/2)\text{ to }B-1}(SS[j]-SS[j+1])$
15:     $\hat{B}=\text{length of }\hat{\mathcal{S}}_{k}^{2}$
16:     $\gamma_{2}=\min(SS[idx],1-\frac{s/2-B}{\hat{B}}\text{quantile of }\hat{\mathcal{S}}_{k}^{2})$
17:     for $u$ in $\hat{\mathcal{S}}_{k}^{2}$ do
18:         if $u>\gamma_{2}$ then
19:              $\hat{\mathcal{S}}_{k}^{2}=\hat{\mathcal{S}}_{k}^{2}\setminus\{u\}$
20:              $\mathcal{S}_{k}^{2}=\mathcal{S}_{k}^{2}\cup\{u\}$
21:              Move corresponding points from $\hat{\mathcal{S}}_{k}^{1}$ to $\mathcal{S}_{k}^{1}$
22:         end if
23:     end for
24:end for
III-B2 Max strategy
The score for the max strategy is only for overlapping neighbors between filtered centers in different groups, and is defined as
$\operatorname{score}_{i|k}=\frac{|\{c\in\mathbf{g}_{k}\mid\mathbf{x}_{i}\in\mathbf{N}_{c}\}|}{|\mathbf{g}_{k}|}.$ (17)
Observations that are neighbors of filtered centers in different groups are allocated to the cluster with the highest score. Unlabeled observations either do not have the highest unique score or are not neighbours of any filtered center.
III-B3 Classification
Among three built-in classification methods for the remaining unlabelled observations, MDC is the non-parametric method based on data depth. MDC assigns each observation to the cluster within which it achieves its maximum depth value. To ensure consistency in our methodology, we employ the same type of data depth in the MDC as we used in defining the local centers. When Mahalanobis depth is chosen, the covariance matrix in the classification step is estimated using traditional moment estimation methods. This is based on the theoretical consideration that temporary clusters are subsets of clusters and all observations in these temporary clusters are selected with a degree of confidence.
As for the KNN method, it is an intuitive classification method and is suitable for non-convex shapes. Given that we have already constructed the similarity matrix $\mathbb{S}$, the KNN method in our algorithm identifies neighbors through depth-based similarity, rather than relying on Euclidean distance. As such, the method identifies neighbors directly within $\mathbb{S}$, thus eliminating the need for additional computational steps.
The last method, RF, is a renowned machine learning algorithm and has a wide range of applications. In classification, it operates without assumptions on the shapes of the clusters and can demonstrate satisfactory performance even in high-dimensional spaces.
IV Exploring Parameter Selections
The choice of parameters $s$ (neighborhood size) and $\delta$ (neighborhood similarity threshold) affect the resulting clustering differently. The neighborhood size $s$ directly affects the defined neighborhoods for each observation and hence impacts the construction of the set $\mathbf{C}$ of filtered centers. Under the max strategy, for any given $s$, the value of $\delta$ further influences how the filtered centers are grouped. Basically, a smaller value of $\delta$ results in fewer, and larger, groups of filtered centers.
IV-A Hierarchical Structure of grouping under the max strategy
Under the max strategy, given a certain $s$, all possible values of the number of groups can be derived. By starting at $\delta=\delta_{0}$ and progressively reducing $\delta$, all $\delta$’s at the time that the number of groups changes can be recorded.
The initial choice for $\delta$ can be based on the following value:
$\delta_{U}:=\min_{c}\max_{c^{\dagger}}\{\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}}):c\neq c^{\dagger}\in\mathbf{C}\}.$ (18)
For any $\delta\geq\delta_{U}$, there will be at least one group with a single filtered center. It is thus desirable to start with a smaller value, say $\delta_{0}=0.99\delta_{U}$, so that multiple smaller subsets are connected to form clusters.
Suppose $\delta_{0}$ is chosen, partitioning $\mathbf{C}$ into groups $\mathbf{g}_{1},\mathbf{g}_{2},\ldots$ Now, if we reduce the $\delta$ value from $\delta_{0}$ to a value below the maximum similarity among neighborhoods of centers between any two groups $\mathbf{g}_{i}$ and $\mathbf{g}_{j}$, as described by:
$\delta<\max\{\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}}):c\in\mathbf{g}_{i},c^{\dagger}\in\mathbf{g}_{j}\},$ (19)
then $\mathbf{g}_{i}$ and $\mathbf{g}_{j}$ will merge. The value in (19) is referred to as the grouping critical point for $\mathbf{g}_{i}$ and $\mathbf{g}_{j}$. Observing how grouping results evolve as $\delta$ decreases from $\delta_{0}$ reveals the hierarchical clustering structure.
$\begin{pmatrix}1&0.6&0&0&0&0\\ 0.6&1&0&0&0&0\\ 0&0&1&0.5&0.3&0.2\\ 0&0&0.5&1&0.4&0.1\\ 0&0&0.3&0.4&1&0.6\\ 0&0&0.2&0.1&0.6&1\end{pmatrix}$ $\begin{pmatrix}1&0&0&0\\ 0&1&0.5&0.3\\ 0&0.5&1&0.4\\ 0&0.3&0.4&1\end{pmatrix}$ $\begin{pmatrix}1&0&0\\ 0&1&0.4\\ 0&0.4&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&1\end{pmatrix}$ $\boldsymbol{G}=\{6,4,3,2\}$$\boldsymbol{\delta}=\{0.6,0.5,0.4,0\}$
Figure 5: A toy example illustrating the hierarchical structure in grouping filtered centers, alongside the relationship between the threshold $\delta$ and the number of groups $G$. This process generates two lists, $\boldsymbol{G}$ and $\boldsymbol{\delta}$, which store the number of groups and the corresponding threshold values at each stage, respectively. The elements of these lists are displayed in the top right.
We use a toy example to illustrate the process, shown in Figure 5. Each square matrix is referred as a “grouping matrix”. The order of the matrix denotes the number of groups, and its entries identify the grouping critical points among these groups for a given $\delta$ value. In each step, the grouping matrix obtained from the previous one by merging columns and rows that contain the largest non-trivial (not 00 or $1$) value. The merging process retains the larger of the values in the corresponding entries of the columns and rows. Eventually, we are left with completely disjoint groups, represented by the identity matrix as the grouping matrix.
The process of grouping generates two key lists: one for the number of groups and another for the corresponding threshold values. More precisely, in the toy example, the potential numbers of groups are $6$, $4$, $3$, and $2$, with the $\delta$ values ranging between the intervals $[0.6,\delta_{U})$, $[0.5,0.6)$, $[0.4,0.5)$, and $[0,0.4)$ at each stage. The hierarchical structure ensures that if the number of clusters is known or can be estimated using any methods (refer to [48] for estimation techniques), a suitable $\delta$ can be directly determined from the relationship between the lists $\boldsymbol{G}$ and $\boldsymbol{\delta}$.
IV-B Depth-Based Clustering Metric
The next challenge of parameter selection is to identify the optimal pair of parameters $(s,\delta)$, which requires an appropriate internal metric to determine the best clustering among various results. However, the DLCC under the max strategy often addresses clustering problems involving non-convex shapes or closely positioned clusters of significantly different sizes, where well-known internal metrics such as ASW and CHI encounter difficulties. The limitations of these metrics are illustrated using two synthetic datasets, SmileyF and Bainba — see Figures 1 and 6, and Table II. In both figures, the clustering results in (a) are the actual labels for the datasets, for which both ASW and CHI yield low scores. Both criteria awarded high scores to the clustering result (c) in the SmileyF dataset, although human assessment considers (c) to perform the poorest. Similarly, both criteria favour clustering result (b) for the Bainba dataset, which contains only two clusters. This discrepancy highlights the need for a metric that better handles clustering of complex shapes.
Figure 6: Clustering results for the synthetic dataset Bainba. Panel (a) shows the ground truth, (b) merges two distinct clusters into one, and (c) incorrectly clusters a point at the top of the central elliptical cluster.
TABLE II: ASW and CHI values corresponding to the clustering results in Figures 6 and 1.
DATA SET Label (a) (b) (c)
Bainba: ASW $0.341$ $0.491$ $0.339$
Bainba: CHI $642.13$ $842.52$ $641.99$
SmileyF: ASW $-0.146$ $-0.152$ $0.435$
SmileyF: CHI $45.25$ $43.18$ $1470.62$
The DC metric is introduced to evaluate and compare different clustering outcomes when the ground truth is not available, facilitating the identification of the optimal pair of parameters. The main idea is to calculate within- and between-cluster similarities in a small, fixed region instead of computing similarities globally to handle non-convex clusters. This approach, inspired by the DBCA methodology [33], is adapted in our context with a broader scope. While DBCA specifies similarity as $D_{\mathrm{MD}}^{i}(\mathbf{z})$, we generalize this concept to accommodate various forms of similarity matrices, $\mathbb{S}$. Some definitions from DBCA are listed below:
Definition 6 (Core Neighbors).
Given a depth similarity threshold $\eta$, a point $\mathbf{x}_{j}$ is a core neighbor of $\mathbf{x}_{i}$ (with respect to $\eta$) if $\mathbb{S}_{ij}\geq\eta$.
Definition 7 (Depth-Connection).
A point $\mathbf{x}_{u}$ is depth-connected to $\mathbf{x}_{v}$ if there is a chain of points $\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{i}$, where each two adjacent points are core neighbors for each other, $\tilde{\mathbf{x}}_{1}=\mathbf{x}_{u}$, and $\tilde{\mathbf{x}}_{i}=\mathbf{x}_{v}$.
Definition 8 (DBCA-Cluster).
A chain $\mathbf{V}=\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{i}\}$ is called a maximal depth-connected chain if there is no $\mathbf{x}_{j}\notin\mathbf{V}$ satisfying $\sup_{\mathbf{x}_{i}\in\mathbf{V}}\mathbb{S}_{ij}\geq\eta$. Let $\tilde{n}$ be the minimum number of points required to form a cluster in DBCA, then each cluster is formed by observations that are in the same maximal depth-connected chain with length at least $\tilde{n}$.
Definition 9 (Outliers).
All observations that remain unlabelled in DBCA are regarded as outliers.
In Definition 6, larger $\eta$ corresponds to fewer core neighbors, which corresponds to a smaller fixed region. Let $\eta_{X}$ be the largest value of $\eta$ such that all observations in the dataset are depth-connected. Let $k$ be the label of a cluster in the clustering result which is to be examined by the DC metric, and $\eta_{k}$ be the maximum $\eta$ value such that the cluster is decomposed into a single DBCA-cluster and outliers (without considering observations in other clusters). Let $\mathcal{X}_{k}$ be the set of observations in cluster $k$. Let $n_{k}=|\mathcal{X}_{k}|$ and $\mathbf{O}_{k}$ be the set of DBCA-outliers. For each observation $\mathbf{x}_{i}\in\mathcal{X}_{k}$, we define $\mathcal{J}_{k}^{i}=\{j|\mathbf{x}_{j}\in\mathcal{X}_{k},\mathbf{x}_{j}\neq\mathbf{x}_{i},\mathbb{S}_{ij}>\eta_{k}\}$ as the set of indices. Following this, the within cluster similarity $J_{k}$ of cluster $k$ is defined by
$J_{k}=\sum_{i=1}^{n_{k}}w_{i}\mu_{i},$
where
$\mu_{i}$ $=$ $\begin{cases}\dfrac{\sum\limits_{j\in\mathcal{J}_{k}^{i}}\mathbb{S}_{ij}}{|\mathcal{J}_{k}^{i}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{O}_{k},\\ \max\{\mathbb{S}_{ij}\mid\mathbf{x}_{j}\in\mathcal{X}_{k}\setminus\mathbf{O}_{k}\},\text{ if }\mathbf{x}_{i}\in\mathbf{O}_{k},\end{cases}$
$w_{i}$ $=$ $\begin{cases}\dfrac{|\mathcal{J}_{k}^{i}|}{\sum\limits_{\mathbf{x}_{v}\in\mathcal{X}_{k}\setminus\mathbf{O}_{k}}|\mathcal{J}_{k}^{v}|+|\mathbf{O}_{k}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{O}_{k},\\ \dfrac{1}{\sum\limits_{\mathbf{x}_{v}\in\mathcal{X}_{k}\setminus\mathbf{O}_{k}}|\mathcal{J}_{k}^{v}|+|\mathbf{O}_{k}|},\text{ if }\mathbf{x}_{i}\in\mathbf{O}_{k}.\end{cases}$
The weights $w_{i}$ in $J_{k}$ give observations with more core neighbors more weight and give outliers less weight. By considering DBCA-outliers, the DC metric can be made more robust by avoiding the significant adverse impact of an individual outlier.
For the cluster $k$, the set of inner observations is
$\mathbf{I}_{k}:=\left\{\mathbf{x}_{i}\in{\mathcal{X}_{k}}:\mathbb{S}_{ij}<\eta_{X}\text{ for all }\mathbf{x}_{j}\in{\mathbf{X}\setminus\mathcal{X}_{k}}\right\},$
where $\mathbf{X}$ is the total dataset. If $\mathbf{x}_{i}\in\mathbf{I}_{k}$ then $\mathbf{x}_{i}$ is called an inner observation of cluster $k$, i.e., observations in $\mathcal{X}_{k}$ that are not similar to any observation outside of $\mathcal{X}_{k}$. For defining the between cluster similarity $H_{k}$ of cluster $k$, we set $\mathcal{H}_{k}^{i}=\{h|\mathbf{x}_{h}\notin\mathcal{X}_{k},\mathbb{S}_{ih}>\eta_{X}\}$ to include indices of observations from other clusters that are core neighbors of points in $\mathcal{X}_{k}$ when $\eta=\eta_{X}$. $H_{k}$ is then defined as
$H_{k}=\sum_{i=1}^{n_{k}}\tilde{w}_{i}\tilde{\mu}_{i},$
where,
$\tilde{\mu}_{i}$ $=$ $\begin{cases}\dfrac{\sum\limits_{h\in\mathcal{H}_{k}^{i}}\mathbb{S}_{ih}}{|\mathcal{H}_{k}^{i}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{I}_{k},\\ \eta_{X},\text{ if }\mathbf{x}_{i}\in\mathbf{I}_{k},\end{cases}$
$\tilde{w}_{i}$ $=$ $\begin{cases}\dfrac{|\mathcal{H}_{k}^{i}|}{\sum\limits_{\mathbf{x}_{u}\in\mathcal{X}_{k}\setminus\mathbf{I}_{k}}|\mathcal{H}_{k}^{u}|+|\mathbf{I}_{k}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{I}_{k},\\ \dfrac{1}{\sum\limits_{\mathbf{x}_{u}\in\mathcal{X}_{k}\setminus\mathbf{I}_{k}}|\mathcal{H}_{k}^{u}|+|\mathbf{I}_{k}|}\,\text{if }\mathbf{x}_{i}\in\mathbf{I}_{k}.\end{cases}$
In $H_{k}$, the weights $\tilde{w}_{i}$ give observations which are close to other clusters more weight, but less weight for inner observations.
The total DC metric for a clustering result is then given by
$\text{DC}=\sum_{k}\frac{n_{k}}{n}(J_{k}-H_{k}),$ (20)
which gives more weight to the clustering performance in clusters with more observations. Some properties of the DC metric are summarized as follows:
  1. Property 1:
    For each observation $\mathbf{x}_{i}\in\mathcal{X}_{k}$, reducing the threshold from $\eta_{k}$ to $\eta_{k^{*}}$ implies $\mu_{i^{*}}\leq\mu_{i}$.
  2. Property 2:
    For each observation $\mathbf{x}_{i}\in\mathcal{X}_{k}$, reducing the threshold from $\eta_{k}$ to $\eta_{k^{*}}$ implies $\tilde{\mu}_{i^{*}}\leq\tilde{\mu}_{i}$.
  3. Property 3:
    For any cluster $k$, the $\mu_{i}$ given by any outlier is less than $\eta_{k}$.
  4. Property 4:
    For any cluster $k$, $H_{k}\geq\eta_{X}$, higher proportions of inner observations in $k$ leads to smaller $H_{k}$.
Table III displays the DC metrics for the aforementioned clustering results for the Bainba and SmileyF datasets. Notably, the ranking of these scores aligns more closely with human assessments. Thus, using the DC metric allows for the selection of the parameters for the DLCC algorithm under the max strategy.
TABLE III: The DC metrics for clustering results in Figures 6 and 1, where $\tilde{n}=3$. Note this score correctly assesses (a) as the best assignment of labels among the three panels.
(a) (b) (c)
Bainba $0.046$ $0.044$ $0.045$
SmileyF $0.174$ $0.164$ $0.079$
V Experiments and Evaluations
In this section, we evaluate the performance of DLCC under various strategies, comparing it with several other clustering algorithms on both synthetic and real-world datasets.
External Metric. To compare the performance of different clustering algorithms on the same datasets, we use the clustering error (CE) [28, p. 2774] and adjusted Rand index (ARI) [51]. The CE simply computes the proportion of mismatched points. This makes CE intuitive, with a value range of 00 (perfect match) to $1$. The ARI, on the other hand, measures pairwise similarity between clustering results and ground truth. The expected value fo the ARI under random class assignment is 0 and it takes a value $1$ under perfect agreement.
Methods for comparison. We compare our approach with several representative clustering methods. These include the center-based algorithm K-means and its state-of-the-art variant LW-K-means; the recent advanced DPC method R-MDPC; the subspace clustering algorithm SSC; the model-based clustering algorithm GMM (GPCM family);
V-A Description of Datasets
Figure 7 presents four synthetic datasets. The first subfigure is simulated using custom code, while the second and third subfigures are generated with the mlbench [53] package in R [52]. The last subfigure provides a two-dimensional visualization of the ray shapes data, which is derived from a study on spectral clustering [54]. These synthetic datasets represent different clustering scenarios with non-convex shapes, clusters of varying sizes and densities in close proximity, and large overlapping regions. Each of these characteristics presents unique challenges for clustering.
Figure 7: Graphs of synthetic datasets: (a) Starbeam; (b) Blend; (c) Bainba; (d) Ray [54]. In $(d)$, $x$ and $y$ axes are the first two variables of ray data. Colors represent true labels.
As for the real-world datasets, we utilize five benchmark datasets from the UCI Machine Learning Repository [55]: iris, seed, wine, optidigits, pendigits. Additionally, we test high-dimensional datasets ($d\geq 100$) with $n>d$, including Yale B [56] and HAR [57], which represent human-face and human-activity datasets, respectively, addressing clustering problems in different domains. Among these, Yale B, optidigits, and pendigit are sourced from the R package PPCI [58]. Furthermore, we consider two microarray datasets: leukemia [59] and SuCancer [60], which have large dimensions $d$ but small sample sizes $n$.
V-B Result comparisons
Table IV provides the settings, used for running DLCC on the above-mentioned datasets. The max strategy tends to perform better with non-convex or unbalanced sized clusters. Conversely, the min strategy is more effective for overlapping clusters of similar sizes.
TABLE IV: DLCC settings for different datasets, where CM stands for classification method.
Datasets Depth Strategy CM maxdepth ifloop
Starbeam MD Max KNN False --
Starbeam SD Max KNN False --
Blend MD Max KNN False --
Blend SD Max KNN False --
Bainba MD Max MDC True --
Bainba SD Max KNN False --
Ray MD Min MDC False True
Ray SD Min RF False False
Iris MD Min MDC True False
Iris SD Min RF False False
Seed MD Min KNN True True
Seed SD Min KNN False False
Wine MD Min MDC True True
Wine SD Min MDC False False
Yale B SD Min RF False False
Optidigit SD Min KNN False False
Pendigits MD Min KNN False False
Pendigits SD Min KNN False False
HAR SD Min KNN False False
Leukemia SD Max MDC False --
SuCancer SD Max RF False --
TABLE V: Comparative analysis of selected clustering algorithms on several datasets ( “–” means a method is not applicable).
Datasets Metric DLCCMD DLCCSD Kmeans LW-Kmeans SSC GMM R-MDPC
Starbeam ARI $\mathbf{1}$ $\mathbf{1}$ $0.4822$ $0.4822$ $0.3052$ $0.0157$ $0.8123$
$n=780,d=2,K=2$ CE $\mathbf{0}$ $\mathbf{0}$ $15.26\%$ $15.26\%$ $22.31\%$ $40.26\%$ $4.87$%
Blend ARI 0.85950.8595{0.8595}0.8595 $\mathbf{0.8605}$ $0.7266$ $0.7365$ $0.5716$ $0.7518$ $0.8453$
$n=2000,d=2,K=10$ CE ${12.50}\%$ $\mathbf{8.65}\%$ $25.70\%$ $18.90\%$ $31.35\%$ $21.45\%$ ${9.10}\%$
Bainba ARI $\mathbf{1}$ $\mathbf{1}$ $0.4961$ $0.9814$ $0.4276$ $0.7436$ $\mathbf{1}$
$n=600,d=2,K=3$ CE $\mathbf{0}$ $\mathbf{0}$ $32.00\%$ $0.67\%$ $30.17\%$ $10.50\%$ $\mathbf{0}$
Ray ARI $\mathbf{0.8261}$ $0.7521$ $0.6388$ $0.7090$ $0.7081$ $0.8245$ $0.5364$
$n=1000,d=5,K=4$ CE $\mathbf{7.00}\%$ $10.00\%$ $15.00\%$ $11.90\%$ $12.60\%$ $7.10\%$ $37.70\%$
Iris ARI $\mathbf{0.9039}$ $0.8857$ $0.7302$ $0.8857$ $0.6138$ $\mathbf{0.9039}$ $0.8857$
$n=150,d=4,K=3$ CE $\mathbf{3.33}\%$ $4.00\%$ $10.67\%$ $4.00\%$ $17.33\%$ $\mathbf{3.33}\%$ $4.00\%$
Seed ARI $0.7612$ $\mathbf{0.7626}$ $0.7167$ $0.6687$ $0.4462$ $0.6299$ $0.7289$
$n=210,d=7,K=3$ CE $\mathbf{8.57}\%$ $\mathbf{8.57}\%$ $10.48\%$ $12.86\%$ $33.33\%$ $14.76\%$ $10.00\%$
Wine ARI $\mathbf{0.9817}$ $0.9295$ $0.8975$ $0.8516$ $0.8804$ $0.9309$ $0.7269$
$n=178,d=13,K=3$ CE $\mathbf{0.56}\%$ $2.25\%$ $3.37\%$ $5.06\%$ $3.93\%$ $2.25\%$ $9.66\%$
Yale B ARI -- $\mathbf{0.9901}$ $0.5663$ $0.8461$ $0.9484$- $0.7635$
$n=2000,d=600,K=10$ CE -- $\mathbf{0.45}\%$ $31.85\%$ $13.25\%$ $2.40\%$- $12.10\%$
Optidigits ARI -- $0.8144$ $0.5682$ $0.5799$ $0.7540$ $0.4878$ $\mathbf{0.8543}$
$n=5620,d=64,K=10$ CE -- $9.40\%$ $28.81\%$ $28.26\%$ $14.84\%$ $36.89\%$ $\mathbf{8.75}\%$
Pendigits ARI $0.6488$ $0.6661$ $0.5526$ $0.5756$ $\mathbf{0.7069}$ $0.6522$ $0.6657$
$n=10992,d=16,K=10$ CE $20.52\%$ $18.22\%$ $32.25\%$ $27.83\%$ $\mathbf{15.56}\%$ $20.58\%$ $19.17\%$
HAR ARI -- $0.6564$ $0.4610$ $0.4556$ $\mathbf{0.7005}$- $0.4373$
$n=10299,d=561,K=6$ CE -- $\mathbf{18.68}\%$ $40.00\%$ $41.51\%$ $23.11\%$- $51.06\%$
Leukemia ARI -- $\mathbf{0.8897}$ $\mathbf{0.8897}$ $\mathbf{0.8897}$ $0.7371$- $0.7851$
$n=72,d=3571,K=2$ CE -- $\mathbf{2.78}\%$ $\mathbf{2.78}\%$ $\mathbf{2.78}\%$ $6.94\%$- $5.56\%$
SuCancer ARI -- $0.1154$ $-0.003$ $-0.003$ $\mathbf{0.1234}$- $0.1154$
$n=174,d=7909,K=2$ CE -- $32.76\%$ $47.70\%$ $47.70\%$ $\mathbf{32.18}\%$- $32.76\%$
Table V illustrates the performance of DLCC with both MD and SD, compared to other clustering algorithms. For all comparing methods, we pick the result achieving the highest ARI as the best. For K-means and LW-K-means, we report the best results from $10$ runs. For the R-MDPC method, following the suggestions, we fix $cv=1$ and test all neighborhood sizes from $0.5\sqrt{n}$ to $2\sqrt{n}$ [18]. For SSC, the best combinations are chosen with regards to the hyper parameter $\alpha$ (ranging from $10$ to $50$ with a step size of $10$) and the logical parameters “affine” (indicating if affine subspaces are considered) and “outlier” (signifying if the data might be contaminated by outliers). In the case of GMM, the R package mixture [61] is used to identify the optimal model from the GPCM family. As for DLCC, when RF is employed as the classification method, its inherent randomness led us to conduct $100$ separate runs with unique seeds. The corresponding internal metric for each run was calculated, and the median performance score was recorded in the table.
Table V clearly demonstrates that DLCC yields impressively competitive results across all datasets. Specifically, in the synthetic datasets, both versions of DLCC perform perfectly in Starbeam and Bainba, while they meet some challenges in Blend. As detailed in Section III-A, the max strategy assumes that, within any given cluster, each subset used for defining local centers is convex. Although the filtering procedures for local centers can identify the areas of non-convex subsets, they are unable to further distinguish individual clusters within these regions and, in the current context, to identify the two distinct spirals. DLCCMD achieves the best performance on the Ray dataset, closely matching that of GMM, as its similarity matrix is constructed based on the EEV model, a member of GMM. DLCCSD proves highly adaptable for high-dimensional data applications, delivering the best results in the Yale B and Leukemia datasets and ranking among the top performers on other high-dimensional datasets. In contrast, each of the other algorithms displays distinct strengths and limitations. For instance, K-means and LW-Kmeans produce satisfactory results for convex datasets, while they do not perform well on datasets with non-convex shapes. R-MDPC performs well on non-convex datasets and excels on the Optidigits dataset. However, it faces challenges with datasets that have large overlapping regions with similar densities, such as Ray and HAR. SSC excels with high-dimensional datasets but falters when the dimension $d$ is less than or close to the number of clusters $K$. Both GMM and DLCCMD face challenges in high-dimensional data due to their reliance on covariance matrix estimations. Accurately estimating covariance matrices requires significantly more observations per cluster than the number of dimensions, which is often not feasible for high-dimensional data.
V-C Performance in estimating the number of cluster
For the max strategy, the relationship between $\delta$ and the number of clusters is detailed in Section IV. Here, we evaluate the effectiveness of Algorithm 1 in estimating the number of clusters under the min strategy. To ensure a fair comparison, the results presented in Table V assume that the true number of clusters is known. In this section, we focus on the scenario where $K$ is not pre-specified, specifically discussing DLCCSD as it is applicable to all datasets.
To provide insights into the performance of our method, we compare it with a well-established spectral clustering method, Spectrum [62], which can automatically estimate the number of clusters during the clustering process without requiring extensive parameter tuning. The implementation of Spectrum is based on the R package Spectrum [63], and we report the better results between the eigengap method (largest gap between eigenvalues) and the multimodality gap method.
Table VI demonstrates that without pre-defined $K$, DLCCSD consistently outperforms Spectrum in terms of ARI. Although DLCCSD estimates two additional clusters in the Pendigits dataset, its ARI is even higher compared to the scenario where $K$ is fixed at $10$. For the Optidigits and HAR datasets, DLCCSD underestimates the number of clusters by one. Specifically, it merges clusters “3” and “8” into a single cluster in Optidigits, and in HAR, it groups walking, walking upstairs, and walking downstairs into one cluster. This clustering also makes sense based on human assessments.
TABLE VI: Comparison of performance between DLCCSD and Spectrum across datasets. Here, $\hat{K}$ denotes the estimated number of clusters, with differences from true $K$ in parentheses.
Method Ray Iris Seed Wine Yale B Optidigits Pendigits HAR
$\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI
DLCCSD $4$ (00) $0.7521$ $3$ (00) $0.8857$ $3$ (00) $0.7626$ $3$ (00) $0.9295$ $10$ (00) $0.9901$ $9$ ($-1$) $0.7174$ $12$ ($+2$) $0.6885$ $4$ ($-2$) $0.4813$
Spectrum $3$ ($-1$) $0.6468$ $3$ (00) $0.8508$ $3$ (00) $0.5702$ $3$ (00) $0.9170$ $9$ ($-1$) $0.7575$ $9$ ($-1$) $0.6527$ $10$ (00) $0.5086$ $10$ ($+4$) $0.3242$
V-D Complexity Analysis
The computational cost of the DLCC algorithm is dominated by constructing the depth-based similarity matrix and estimating local centers. For both versions, the first term corresponds to constructing the similarity matrix, and the second term accounts for estimating local centers, with $s\ll n$.
  • DLCCMD: With pre-estimated covariance matrices (Section II-B), the complexity is $O(n^{2})+O(ns^{2}+ns)$, where $O(ns^{2}+ns)$ covers covariance matrix estimation and depth computation for neighborhoods. As $s\ll n$, the overall complexity simplifies to $O(n^{2})$.
  • DLCCSD: The complexity is $O(n^{3})+O(ns^{2})$. Constructing the similarity matrix requires $O(n^{3})$, as computing SD values for $n$ points ($O(n^{2})$) across $n$ reflected datasets introduces an additional $n$ factor. Estimating local centers for neighborhoods adds $O(ns^{2})$, resulting in an overall complexity of $O(n^{3})$.
Figure 8: Running time of DLCC on simulated datasets without acceleration techniques, where blue represents DLCCSD and red represents DLCCMD: (a) Fixed the dimension at $30$ while increasing the sample size up to $5060$, y-axis up to $700$s; (b) Fixed the sample size at $1000$ while increasing the dimension up to $1000$, y-axis up to $25$s.
Figure 8(a) shows that the experimental running times align with theoretical complexities: $O(n^{2})$ for DLCCMD and $O(n^{3})$ for DLCCSD. In Figure 8(b), the running time of DLCCMD appears irregular as dimensionality increases because when estimating the covariance matrix, the method defaults to the identity matrix if singularity issues arise, which are more likely when the dimensionality is not significantly smaller than the sample size. In contrast, DLCCSD shows a linear trend due to the SD computation involving the $L^{2}$ norm of the average of unit vectors, with a complexity of $O(n^{3}d)$ when $d$ is large.
While DLCC can handle datasets with approximately $10,000$ samples, its computational cost is dominated by the depth-based similarity matrix, limiting its scalability to larger datasets. However, DLCCSD is computationally efficient with increasing dimensionality. Future work could explore approximation methods to accelerate the algorithm. For instance, defining local centers does not necessarily require computing the entire similarity matrix but rather identifying suitable neighborhoods for the centers. Fast approximate neighborhood search methods (e.g., [64]) could be employed to estimate candidate local centers (with size much smaller than $n$), followed by exact depth similarity computations for points relative to these candidates.
VI Conclusions and Future work
In this paper, we have proposed an innovative clustering framework, DLCC, which first defines a depth-based similarity matrix to provide orderings of neighbors for each observation. DLCC then uses depth to define representative points, or “local centers”, with the number of these local centers controlled by the parameter indicating the neighborhood size, $s$. Our results show that DLCC delivers superior clustering performances under a wide range of scenarios. This includes handling both convex and non-convex shapes, managing clusters of balanced and unbalanced sizes, working with well-separated and overlapping clusters, and analyzing clustering problems in both low and high-dimensional spaces, including datasets with small $n$ and large $d$.
For future research, a key direction is to develop an efficient DLCC method for handling larger datasets. Moreover, while this paper focuses on MD and SD for their computational simplicity, exploring other depth definitions, such as the integrated rank-weighted depth [65], could further enhance the DLCC framework. Furthermore, as depth is also a well-defined concept for functional data, e.g., band depth [66], there are many potential avenues for extending the DLCC framework to functional data analysis.
Acknowledgments
This work is supported by the Canada Research Chairs program (PM), a Dorothy Killiam Fellowship (PM), and respective NSERC Discovery Grants (AL, PM).

### v2

@@
 However, data depth typically fails to capture the multimodal characteristics of data, something of the utmost importance in the context of clustering.
-To overcome this, DLCC makes use of a local version of data depth that is based on subsets of data.
+To overcome this, DLCC considers data depth based on depth-defined subsets of data.
 From this, local centers can be identified as well as clusters of varying shapes.
-Furthermore, we propose a new internal metric based on density-based clustering to evaluate clustering performance on non-convex clusters.
 Overall, DLCC is a flexible clustering approach that seems to overcome some limitations of traditional clustering methods, thereby enhancing data analysis capabilities across a wide range of application scenarios.
-Index Terms:
-Clustering, statistical depth, similarity matrix, Mahalanobis depth, spatial depth, internal metric.
 I Introduction
@@
 In turn, this has increased the need for flexible and computationally efficient clustering methods [4] in many fields, such as pattern recognition [5], computer vision [6, 7] and image processing [8].
-The complexity of data distributions has led to the development of numerous clustering algorithms, each having a unique set of strengths and constraints.
-We next briefly discuss different traditional clustering methods to highlight their distinct characteristics and limitations.
-In this, and in what follows, we consider the objects to be clustered as points in the $d$ dimensional real space $\mathbb{R}^{d}$.
+The complexity of real-world data distributions has given rise to numerous clustering algorithms with different assumptions and objectives.
+Across those methodologies, many clustering approaches rely on identifying representative points (exemplars), such as centers, density peaks, or anchors, and forming clusters by grouping these exemplars and assigning remaining points accordingly.
+In the following, we review representative clustering methods under this unified viewpoint.
 I-A Clustering Methods
 I-A1 Center-based Clustering
-Center-based clustering methods, with k-means clustering [9] being a prime example, are often utilized due to their simplicity and effectiveness.
-However, k-means clustering has several limitations.
-Firstly, it is sensitive to outliers as it uses the means of the points in each cluster as cluster centers.
-Secondly, even in outlier-free contexts, this approach is most adequate when clusters are spherically symmetric, which may not always be representative of the actual data distribution.
-Lastly, the algorithm is sensitive to initialization and can easily get trapped in local minima.
-To address some of these issues, several variants of the k-means algorithm have been developed:
-  • Partitioning around medoids (PAM; [10]), was introduced to combat the sensitivity to outliers. Instead of cluster means, PAM uses medoids (a form of multivariate median) as cluster centres.
-  • Weighted k-means (W-k-means; [11]) and lasso weighted k-means (LW-k-means; [12]) were designed to address the underlying spherical symmetry assumption for clusters.
-These both provide a more flexible k-means approach by incorporating a weight vector for each variable.
-LW-k-means improves W-k-means by adding a lasso penalty to the objective function, enabling feature selection in high-dimensional data.
-  • Power k-means [13] and kernel power k-means [14] were developed to address the sensitivity of k-means to initialization and tendency to get trapped in local minima by redefining the objective function using power means.
-Kernel power k-means extends this by incorporating multiple kernel configurations to handle non-linear separable data via higher-dimensional projections.
-I-A2 Density-Based and Hierachical Clustering
+Center-based clustering methods form clusters around a set of representative points defined as cluster centers, typically obtained by minimizing within-cluster dispersion under a prescribed distance measure.
+The classical k-means algorithm [9] exemplifies this paradigm, where exemplars are defined as arithmetic means of cluster members.
+While computationally efficient and conceptually simple, k-means is sensitive to outliers, biased toward spherical cluster shapes, and highly dependent on initialization, often leading to suboptimal local minima.
+Many variants of k-means have been proposed to address those limitations and refine the definition of exemplars.
+Partitioning around medoids (PAM; [10]) replaces means with medoids, improving robustness by restricting exemplars to sample points.
+Weighted k-means (W-k-means; [11]) and lasso weighted k-means (LW-k-means; [12]) provide a more flexible k-means approach by incorporating a weight vector for each variable, where LW-k-means enables feature selection in high-dimensional data by adding a lasso penalty, partially alleviating the spherical symmetry assumption.
+Beyond hard assignments, fuzzy extensions of k-means reinterpret exemplars as prototypes that attract data points with varying membership strengths.
+Classical fuzzy C-means [13] and its variants define cluster centers as membership-weighted averages, thereby softening hard boundaries while retaining a center-based exemplar structure.
+Equilibrium k-means (EKM; [14]), a recently proposed method, addresses the limitations of fuzzy clustering under class imbalance by modifying the interaction between data points and exemplars through an attraction–repulsion mechanism, resulting in more stable and representative centroid estimation.
+Despite these advances, center-based methods fundamentally rely on distance-driven exemplar definitions and a one-exemplar-per-cluster assumption, which makes them sensitive to metric choice and limits their ability to model complex and nonconvex data geometries.
+I-A2 Density-based Clustering
 Density-based clustering assumes that clusters are regions of high density separated by low-density areas, making it well-suited for detecting arbitrarily shaped clusters.
-For instance, density-based spatial clustering of applications with noise (DBSCAN; [15]) sets a threshold and minimum point count for a region to be considered dense.
-Then, dense regions that can be directly connected with each other are merged into a cluster.
-However, DBSCAN’s reliance on a fixed threshold limits its ability to handle clusters with varying densities.
-Hierarchical clustering, on the other hand, adopts a tree-like approach to iteratively merge or split clusters.
-It offers two main strategies: agglomerative (bottom-up) and divisive (top-down).
-The resulting dendrogram represents cluster relationships, from an all-inclusive cluster at the top to individual objects at the bottom [16].
-While hierarchical clustering provides an intuitive structure, its high computational cost and irreversible decisions make it less suitable for large datasets or complex cluster shapes.
-Density peak clustering (DPC) [17], introduced in $2014$, takes a different approach by combining density and distance to identify cluster centers.
-It locates density peaks based on a decision graph, where density is plotted on the y-axis and the distance to the nearest higher-density point on the x-axis.
-This method allows for manual selection of cluster centers and performs well with varying-density clusters.
-However, it can struggle with robustness, as density peaks may correspond to satellite points rather than true cluster centers.
-An improved version, fast main density peak clustering within relevant regions (R-MDPC) [18], proposed in $2024$, incorporates hierarchical concepts by applying single-linkage clustering to density peaks, defining relevant regions as density-connected areas.
-This enhancement better handles complex cluster shapes by merging single-density peak clusters and reducing satellite peak interference.
-However, for highly overlapping or extremely high-dimensional datasets, R-MDPC may collapse all points into one region, behaving similarly to the original DPC.
-I-A3 Model-Based Clustering
-Model-based clustering typically assumes the points to be clustered stem from an underlying mixture distribution [19], whose components are naturally associated with the clusters to be constructed.
-For example, working from a Gaussian mixture model (GMM), one can identify elliptical clusters with varying orientations, shapes, and sizes.
-Other mixture models handling asymmetric distributions have also been developed, see, e.g., [20], [21], [22], and [23].
-These different families of mixture models lead to clustering methods typically relying on the expectation-maximization (EM) algorithm [24] to find optimal parameter values, which can be computationally intensive.
-Moreover, selection of the most appropriate family of mixture models can be challenging unless knowledge about the data-producing mechanism and/or subject area expertise is available.
-I-A4 Spectral and Subspace Clustering
-Spectral clustering [25], a notable graph-based method, treats clustering as a graph partitioning problem, converting data points into nodes and their similarities into connecting edge weights.
-It is well-suited for high-dimensional datasets, often utilizing a sparse adjacency matrix derived from the k-nearest neighbors methods.
-After representing the data as a similarity graph, the clustering problem can essentially be seen as partitioning the graph in such a way that edges between groups are given low weights and edges within groups high weights [26].
-Spectral clustering has been extended and has led to other clustering techniques.
-It is, for instance, an essential component of modern subspace clustering algorithms.
-These algorithms have gained attention due to their capacity for identifying low-dimensional representations of high-dimensional data [27].
-In particular, sparse subspace clustering (SSC) [28] is a key evolution in this domain.
-Making use of the self-expressiveness property, points are expressed as linear combinations of other points within the same subspace.
-Therefore, SSC merely requires the coefficient matrix of these linear combinations.
-As a result, SSC overcomes two major challenges of spectral clustering: selecting an appropriate neighborood size and handling points near the intersection of two subspaces.
-However, for low-dimensional data, the premise that data points can be effectively represented as linear combinations of the lower-dimensional subspace may falter, potentially leading to a diminished efficacy of SSC.
-This makes SSC a tool that is primarily designed for high-dimensional datasets.
+In this paradigm, exemplars are implicitly or explicitly defined as density maxima or points located within dense regions.
+A classical example is density-based spatial clustering of applications with noise (DBSCAN; [15]), which characterizes clusters as density-connected components, where each component is represented by a “core point”.
+While effective for arbitrary shapes, its reliance on a fixed threshold limits its ability to handle clusters with varying densities.
+Mean shift clustering [16] further adopts a density-driven exemplar definition by iteratively shifting points toward local modes of an estimated density function.
+The resulting modes act as exemplars that summarize local structure without requiring a predefined number of clusters.
+However, the effectiveness of mean shift depends critically on hyperparameters for estimating density (e.g., the kernel bandwidth).
+Density peak clustering (DPC; [17]) explicitly defines exemplars as density peaks that are well separated from higher-density points.
+A key limitation of it is that spurious (satellite) peaks may be promoted to exemplars, which degrades robustness.
+Recent work has therefore incorporated hierarchical ideas to improve exemplar grouping.
+In particular, fast main density peak clustering within relevant regions (R-MDPC; [18]) introduces the notion of relevant regions by first forming single-peak clusters led by density peaks and then merging density-connected single-peak clusters into larger regions, within which main peaks are identified.
+This enhancement better handles complex cluster shapes and reducing satellite peak interference.
+More broadly, such hierarchical schemes provides a generic mechanism for grouping representatives via linkage criteria.
+See also [19, 21, 20] for related clustering methods that adopt hierarchical schemes to merge over-fragmented components.
+Density-based approaches still depend on density estimation and/or distance-based relations among representatives, which can be unstable under strong overlap between clusters or in high-dimensional spaces.
+I-A3 Graph-based Clustering
+Graph-based clustering forms clusters by partitioning a similarity graph, where vertices represent samples and edge weights encode pairwise affinities.
+Spectral clustering (SC; [22, 23]) solves a relaxed graph-cut problem via Laplacian eigen-decomposition and typically applies a discretization step (e.g., k-means) to obtain hard labels.
+To avoid potential information loss from post-processing discretization and to couple graph construction with label learning, constrained Laplacian rank (CLR; [24]) approaches directly produce discrete labels from an optimally structured graph.
+However, both SC and CLR-style formulations can be computationally prohibitive on large-scale data due to full-graph construction.
+Exemplars in graph-based clustering are used not only to summarize data structure but also to reduce the computational burden of graph construction and partitioning.
+This has motivated anchor-based graph clustering, which selects $m\!\ll\!n$ representative anchors and builds a sparse bipartite graph to approximate the original affinity graph [25, 26, 27].
+Although anchors improve scalability, the resulting graph can remain sensitive to the choice of distance metric, graph-construction hyperparameters, and the anchor selection strategy.
 I-B Motivation
-Our objective is to develop a versatile clustering framework capable of addressing diverse challenges, from convex to non-convex cluster shapes, and from low- to high-dimensional data.
-A common theme across clustering methods is their reliance on representative points, or exemplars.
-Center-based methods define exemplars by minimizing within-cluster distances, often relying on distance metrics.
-However, a specific distance measure often performs well only for certain types of data, and adapting to more diverse structures may require mapping data into higher-dimensional spaces via kernel methods.
-The effectiveness of such approaches, however, hinges on selecting an appropriate kernel, which can be challenging and highly sensitive to the data.
-Density-based methods, on the other hand, define exemplars through local density variations.
-These methods identify clusters by connecting dense regions, but the appropriate definition of density often depends on the dimensionality of the data.
-For example, density measures that work well in low-dimensional spaces may lose effectiveness in high-dimensional settings due to the sparsity of data.
-Moreover, in overlapping regions, density connections often struggle to distinguish between clusters.
-Both center-based and density-based approaches share a reliance on exemplars but are inherently limited by their respective definitions of exemplars.
-A versatile clustering framework must overcome these limitations by introducing an adaptive similarity measure that avoids strong parametric assumptions and generalizes across diverse data geometries.
-Equally important is the need for a robust exemplar definition that is consistent across dimensions and cluster shapes, along with effective grouping strategies for these exemplars to form meaningful clusters.
-In response, our method introduces a more flexible exemplar definition by leveraging statistical data depth, a nonparametric tool for multivariate analysis.
+Viewed through an exemplar-based lens, many clustering methods succeed or fail primarily due to how representativeness is defined.
+Most existing approaches tie exemplars to distance- or density-driven criteria (and their associated hyperparameters).
+Such criteria can work well when a single notion of “closeness” or “density” is appropriate, but may yield misleading exemplars when the chosen metric or density scale is mismatched to the data (e.g., Euclidean distance under highly skewed distributions, or DBSCAN-style density reachability in high-dimensional spaces).
+On the one hand, center-based methods typically define a single exemplar per cluster through distance minimization, an assumption that can be overly restrictive for irregular or nonconvex structures.
+On the other hand, density-based methods often induce multiple exemplars (e.g., core points or density peaks), but the resulting exemplar quality depends on reliable density definition.
+These limitations motivate an exemplar definition that remains consistent across dimensionalities and geometries, while allowing the number of exemplars to adapt to local structure.
+In response, we propose the depth-based local center clustering (DLCC) algorithm.
+DLCC addresses this by leveraging statistical data depth, a nonparametric tool for multivariate analysis.
 Tukey [28] pioneered the idea of center-outward ordering in multivariate data, with the introduction of the first depth definition, half-space depth.
 Data depth enables the center-outward ordering of data, making it a potent tool for locating a center for any dataset.
-When considering subsets of the data, data depth with respect to each subset aids in defining multiple exemplars.
+When considering subsets of the data, data depth with respect to (w.r.t.) each subset aids in defining multiple exemplars.
 Naturally, as the subset size grows, the number of exemplars diminishes, and vice versa.
-Building on this concept, we propose the depth-based local center clustering (DLCC) algorithm.
-Equipped with two grouping strategies for depth-defined representative points, DLCC inherits strengths from both center-based and density-based methods, while offering greater flexibility to tackle a wide range of clustering problems.
+To inherit the simplicity of center-based clustering and the shape adaptivity of density-based approaches, DLCC combines depth-defined local exemplars with two grouping strategies for exemplar aggregation.
 Here, the term “depth-based” encompasses two aspects: (1) data depth is used to construct a similarity matrix instead of relying on distance, adaptively capturing the geometric characteristics of each point’s location; (2) data depth measures the centrality of points within each point’s neighborhood, helping to identify exemplars.
-Furthermore, ground truth is typically not available in real-world clustering applications, necessitating the use of internal metrics to evaluate performance.
-Established metrics like the average silhouette width (ASW; [49]), which measures the average distance within clusters relative to the nearest cluster, and the Calinski-Harabasz index (CHI; [50]), which relies on distances from points to their cluster centroids, inherently favor spherical clusters.
-This preference limits their effectiveness in capturing non-convex shapes, as illustrated in Figure 1, where they will incorrectly favor clustering result (c) over (a).
-To address these shortcomings, we propose a new internal metric, the depth-based clustering (DC) metric, designed to more accurately handle complex cluster shapes and assist in parameter selection for our algorithm.
-Further details will be discussed in Section IV-B.
-Figure 1: Clustering results for the synthetic dataset SmileyF. Panel (a) shows the ground truth, (b) misclusters several points in the eyes, and (c) fails to capture the face pattern.
-I-C Main Contributions and Organization of the Paper
-This paper makes the following main contributions:
-  • Depth-Based Similarity Matrix: We improve the efficiency when constructing a depth-based similarity matrix using Mahalanobis depth (MD) and Spatial depth (SD).
-For the MD-based similarity matrix, model-based clustering methods are employed to predefine covariance matrices that approximate the data’s distribution shapes.
-For the SD-based similarity matrix, we develop an efficient matrix-based computation method.
-  • DLCC Clustering Framework: We develop a novel clustering framework, DLCC, which includes the definition of a similarity measure and local centers, and the introduction of methods for filtering and grouping local centers.
-Experiments shows it achieves competitive performance across different types of clustering challenges among the representative algorithms from each type of clustering method referenced in Section I-A.
-  • Internal Evaluation Metric: We propose a depth-based clustering metric as an internal evaluation criterion.
-This metric combines the density-based clustering concept with the depth-based similarity matrix to evaluate clustering performance, particularly for non-convex cluster structures.
-The remainder of the paper is organized as follows: Section II reviews key concepts of data depth and introduces the depth-based similarity matrix.
-Section III provides a detailed description of the DLCC framework.
-In Section IV, we discuss the hierarchical structure of DLCC and present the proposed internal metric for parameter selection.
-Section V showcases experimental results of DLCC using both synthetic and real-world datasets.
-Lastly, Section VI concludes the paper.
-Figure 2 provides the flowchart of the DLCC method, and Table I summarizes the key notations used in this paper to help readers navigate its contents more easily.
-Additional graphs providing further insights into our methods are included in the supplementary file.
-DatasetDepth-basedsimilarity matrixSection IINeighborhoods &Local centersSection III Filtered centersSection III-AGroups of filtered centersSection III-A &\&& Section IVTemporaryclustersSection III-BUnlabelled pointsSection III-BFinal clustersclassification
-Figure 2: Flowchart for the DLCC algorithm, detailing the sections that include the key contents represented in the nodes.
-TABLE I: Summary of the main notations
-Notation Description
-$n$ and $d$ Number of observations and dimensions, respectively;
-$D_{\mathrm{MD/SD}}(\mathbf{z}|\mathbf{X})$ Mahalanobis or spatial depth of $\mathbf{z}$ with respect to $\mathbf{X}$;
-$\mathbf{X}$ A given dataset;
-$\mathbf{X}_{Ri}$ Dataset reflected around $\mathbf{x}_{i}$, i.e., $\mathbf{X}_{Ri}=\mathbf{X}\cup\{2\mathbf{x}_{i}-\mathbf{x}_{j}\,|\,j\neq i\}$;
-$\mathbf{E}$, $\mathbf{\tilde{E}}$ Matrices of unit vector sums from $\mathbf{X}$ and reflected points to $\mathbf{X}$;
-$\mathbb{S}$ Depth-based similarity matrix, with $\mathbb{S}_{ij}$ its $ij$-th entry;
-$s$ Neighborhood size;
-$\mathbf{N}_{i}$ Depth-based neighborhood of $\mathbf{x}_{i}$;
-$r_{i}(\mathbf{x}_{j})$ Depth rank of $\mathbf{x}_{j}$ within $\mathbf{N}_{i}$;
-$c$, $\mathbf{c}$, $\mathbf{C}$ A local center, a certain set of local centers, and the set of all filtered centers, respectively;
-$\mathbb{M}$ Local center neighborhoods’ similarity matrix, with $\mathbb{M}_{ij}$ its $ij$-th entry;
-$\delta$ Similarity threshold for local center grouping ($\mathbb{M}$-based);
-$\mathbf{g}$ A group of local centers;
-$\mathbf{U}_{\mathbf{g}}$ Unique neighbors of group $\mathbf{g}$;
-$G$ and $K$ Number of groups and clusters, typically $G=K$;
-$k$ Cluster index;
-$\mathcal{X}_{k}$ Observations in cluster $k$;
-$f$ Frequency of a local center;
-$\mathcal{P}$ Cumulative proportion of neighbors of local centers to all points;
-$\mathcal{S}_{k}$, $\mathcal{\hat{S}}_{k}$ Score pools for accepting and leaving points unlabelled in cluster $k$;
-$\eta_{k}$, $\eta_{X}$ Depth similarity thresholds for cluster $k$ and the dataset ($\mathbb{S}$-based);
-$\mathcal{I}_{k}$ Indices of points in $\mathbf{g}_{k}$;
-$\mathcal{J}_{k}^{i}$ Indices of core neighbors of $\mathbf{x}_{i}$ in $\mathcal{X}_{k}$ based on $\eta_{k}$;
-$\mathcal{H}_{k}^{i}$ Indices of core neighbors of $\mathbf{x}_{i}$ outside $\mathcal{X}_{k}$ based on $\eta_{X}$;
-$\mathbf{I}_{k}$ Inner observations of cluster $k$;
-$\mathbf{O}_{k}$ DBCA-outliers of cluster $k$;
-II Depth-based Similarity Matrix
-II-A Preliminaries
+I-C Preliminaries about depth and related work in clustering
 A depth function quantifies how “central” a point is within a multidimensional data distribution, providing a meaningful center-outward ordering of points.
@@
 A deepest point $\mathbf{m}=\text{argmax}_{\mathbf{z}\in\mathbb{R}^{d}}D(\mathbf{x}\mid F)$ is usually referred to as a depth median.
-In applications, where the inherent distribution $F$ is usually unknown, depth is calculated with respect to the empirical distribution $F_{n}$ of $\mathbf{X}$, that is $D(\mathbf{z}\mid F_{n})$ is the depth of $\mathbf{z}$ with respect to the sample $\mathbf{X}$.
+In applications, where the inherent distribution $F$ is usually unknown, depth is calculated w.r.t.
+the empirical distribution $F_{n}$ of $\mathbf{X}$, that is $D(\mathbf{z}\mid F_{n})$ is the depth of $\mathbf{z}$ w.r.t.
+the sample $\mathbf{X}$.
 Hereinafter, we will denote the latter as $D(\mathbf{z}\mid\mathbf{X})$ to emphasize the role of the sample on calculating the depth of the point $\mathbf{z}$.
@@
 where $\mathbf{z}$ denotes a point in $\mathbb{R}^{d}$, $\bar{\mathbf{X}}$ is the mean vector, $\mathbf{x}_{i}$ is an observation in $\mathbf{X}$, while $\hat{\mathbf{\Sigma}}$ is the estimated covariance matrix of dataset $\mathbf{X}$.
-Unlike DBSCAN, which uses fixed Euclidean distance for neighborhoods, DBCA employs (2) to define elliptical, affine-invariant neighborhoods.
-Recent work has also explored data depth to identify centroids for k-means initialization in diverse datasets [34, 35].
+Unlike DBSCAN, which defines neighborhoods via a fixed Euclidean radius, DBCA uses (2) to induce elliptical, affine-invariant neighborhoods.
+Subsequent extensions further developed this depth-based framework by refining robustness of covariance matrix estimation and rules of core points [33, 34].
+Another recent development in depth-based clustering was proposed by Francisci et al.
+[35], who build upon the concept of local depth introduced by Agostinelli and Romanazzi [36].
+Their work generalizes local depth to lens depth and $\beta$-skeleton depth, and exploits depth as a surrogate for density estimation to develop a mean-shift like clustering algorithm.
+Beyond density-based paradigms, data depth has also been explored for identifying representative centroids in center-based clustering, particularly for k-means initialization across diverse datasets [37, 38].
 So far, the notion of depth serves an auxiliary role to the more traditional clustering methods and is not employed directly in examining the multimodality of data.
-Herein, two depth functions are considered for our DLCC algorithm: MD and spatial depth (SD).
-For SD [36], the sample version is
+In contrast, DLCC places data depth at the core of both exemplar identification and similarity construction, thereby using depth directly to examine local and global cluster structure.
+The DLCC framework itself is not restricted to a specific depth function, and in principle can be instantiated with a broad class of depth notions.
+In this work, we adopt spatial depth (SD; [39]) as a representative choice, as it offers a favorable trade-off between theoretical interpretability and computational simplicity.
+Although SD is invariant only under similarity transformations, its straightforward computation makes it well suited for high-dimensional applications.
+The sample version of SD is given by
 $D_{\mathrm{SD}}(\mathbf{z}\mid\mathbf{X})=1-\left\|\sum_{i=1}^{n}\frac{\mathbf{z}-\mathbf{x}_{i}}{n\|\mathbf{z}-\mathbf{x}_{i}\|}\right\|,$ (3)
 where $n$ is the total number of observations.
-Although SD is only invariant under similarity transformations, it is nevertheless a sensible choice in the current context due to its straightforward computation in high-dimensional spaces.
 DLCC requires a similarity matrix to determine the neighbors of each observation.
@@
 In contrast, data depth offers a nonparametric, robust, and adaptive alternative.
-As the depths (1) and (3) provide a center-outward ordering and their value ranges from 00 to $1$, the depth values of other points can be viewed as a measure of similarity to the depth median.
+As the depth (3) provide a center-outward ordering and their value ranges from 0 to $1$, the depth values of other points can be viewed as a measure of similarity to the depth median.
 Thus, similarity matrices can be constructed based on data depth.
@@
 It is also the case that $D(\mathbf{x}_{i}|\mathbf{X}_{Ri})=1$ and that $D(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ can be used to quantify the similarity of any point $\mathbf{x}_{j}$ ($j\neq i$) to the point of interest $\mathbf{x}_{i}$.
-Subsequently, by repeatedly computing the depth values for points in $\mathbf{X}$ with respect to $\mathbf{X}_{R1},\mathbf{X}_{R2},\ldots,\mathbf{X}_{Rn}$, a $n\times n$ non-mutual depth-based similarity matrix can be defined.
-We denote it as $\mathbb{S}$, where $\mathbb{S}_{ij}=D(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ represents the entry in the $i$th row and $j$th column.
+Subsequently, by repeatedly computing the depth values for points in $\mathbf{X}$ w.r.t.
+$\mathbf{X}_{R1},\mathbf{X}_{R2},\ldots,\mathbf{X}_{Rn}$, a $n\times n$ non-mutual depth-based similarity matrix can be defined.
+We denote it as $\mathbb{S}$, where $\mathbb{S}_{i,j}=D(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ represents the entry in the $i$th row and $j$th column.
 As will be seen shortly, this can cause some computational challenges.
-This approach to localize the concept of depth was introduced by Paindaveine and Van Bever [37] when they introduced local depth, which was shown to yield non-spherical depth neighborhoods in some contexts.
+This approach to localizing the notion of depth was firstly introduced by Paindaveine and Van Bever [40], who proposed a form of local depth that differs from the one of Agostinelli and Romanazzi [36].
+In particular, their construction was shown to yield non-spherical depth neighborhoods in some contexts.
 The same authors [41] used this concept and introduced an affine-invariant k-nearest neighbor classification algorithm that adapts to the local geometry of the sample and was shown to be consistent.
-II-B Mahalanobis Depth based Similarity Matrix
-Previous studies [39] have employed the approach of using individual observations $\mathbf{x}_{i}$, $i=1,2,\ldots,n$, in place of the mean vector, while incorporating the globally estimated covariance matrix, for depth value computations as given in (2).
-In other words, the center used for calculating the Mahalanobis depth is shifted to each observation, and the depth values of the remaining observations are subsequently recorded.
-These studies commonly employ the minimum covariance determinant (MCD) method [40] to achieve robust results.
-In light of the previous discussion, the mean of $\mathbf{X}_{Ri}$ can easily be seen to be $\mathbf{x}_{i}$, which leads to
-$D_{\mathrm{MD}}(\boldsymbol{z}\mid\mathbf{X}_{Ri})=D_{\mathrm{MD}}^{i}(\boldsymbol{z}),$
-assuming the use of a consistent covariance matrix. This implies that making use of (2) to define a neighborhood of any point is in line with the approach of Paindaveine and Van Bever.
-However, the method of Paindaveine and Van Bever necessitates estimating $n$ distinct covariance matrices, substantially increasing computational load.
-In contrast, the DBCA approach estimates only a single global covariance matrix, which may not adequately reflect the diversity within different clusters.
-To mediate between these extremes, we propose a covariance matrix selection strategy that balances computational efficiency with the ability to capture the unique characteristics of each cluster more effectively than a global approach.
-In the framework of GMM, it is feasible to estimate multiple covariance matrices, ideally keeping the number of such matrices less than or equal to the number of clusters — substantially fewer than $n$ — yet more than one.
-The covariance matrix for cluster $k$, i.e., component $k$, can be decomposed as
-${\mathbf{\Sigma}}_{k}=\lambda_{k}\mathbf{\Gamma}_{k}\mathbf{\Delta}_{k}\mathbf{\Gamma}_{k}^{\prime},$ (4)
-where $\lambda_{k}=|{\mathbf{\Sigma}}_{k}|^{1/d}$, $\mathbf{\Gamma}_{k}$ denotes the matrix of eigenvectors, and $\mathbf{\Delta}_{k}$ is a diagonal matrix of eigenvalues with the condition $|\mathbf{\Delta}_{k}|=1$ [41], [42].
-These elements, respectively, represent size, orientation, and shape of the cluster; see [43] for details.
-In our study, we focus on the EEV model from the Gaussian parsimonious clustering models (GPCM) family, wherein different clusters estimated by the EEV model share common $\lambda_{k}$ and $\mathbf{\Delta}_{k}$ — i.e., $\lambda_{k}=\lambda$ and $\mathbf{\Delta}_{k}=\mathbf{\Delta}$ — but possess distinct $\mathbf{\Gamma}_{k}$.
-When model performance metrics, such as the Bayesian information criterion (BIC; [44]), are similar for results with varying numbers of components, those with fewer components are preferred.
-We do not consider more flexible models within the GPCM family as these could result in GMMs that overfit the data, and the size of the estimated covariance matrices may be too small for constructing the similarity matrix.
-If we assume clusters are all convex, we use a test based on principal component analysis (PCA; [45]) to decide on the method for constructing the similarity matrix.
-With the close correlation between the Mahalanobis distance and PCA, the squared Mahalanobis distance is equal to the sum of squares of standardized principal component scores [46].
-If a significant proportion of variance can be explained by the leading principal components, then the Mahalanobis distance in the space spanned by these components is a suitable metric, especially in the presence of strong inter-variable correlations.
-In contrast, if the variables show no clear correlations, the standard Euclidean distance is deemed a more appropriate measure.
-The test procedure can be summarized as follows: Let $\mathbf{vp}=\{vp_{1},vp_{2},\ldots,vp_{H},\ldots,vp_{d}\}$, which contains the proportion of variance explained by each PC from high to low, where $H$ indicates the number of PCs that explain over $5\%$ variance, but if $H=d$, then $H$ redefined to be $d-1$.
-If either (5) or (6) is satisfied, the dataset passes the test:
-$vp_{1}$ $>$ $0.6,$ (5)
-$\sum_{h=1}^{H}vp_{h}$ $>$ $0.95$ (6)
-When the dataset passes the test, we use the estimated global covariance matrix.
-If it fails, the EEV model is brought into play, offering the benefit of estimating clusters with identical eigenvalues.
-The test can be repeated using the eigenvalues derived from the EEV model.
-If it continues to fail, indicating weak inter-variable correlations even within clusters, and the Euclidean distance becomes sufficient for constructing the similarity matrix, setting the covariance matrix in (2) as the identity matrix.
-However, if it succeeds, suggesting the presence of within-cluster correlations, we employ the covariance matrices estimated by the EEV model.
-Specifically, consider two points, $\mathbf{x}_{i}$ and $\mathbf{x}_{j}$, where $\mathbf{x}_{i}$ is assigned to cluster $1$ and $\mathbf{x}_{j}$ to cluster $2$ by the EEV model.
-The estimated covariance matrices, $\hat{\mathbf{\Sigma}}_{1}$ for cluster $1$ and $\hat{\mathbf{\Sigma}}_{2}$ for cluster $2$, are used to compute the $i$th and $j$th rows of the similarity matrix, respectively.
-While both covariance matrices are of the same size and shape, their orientations differ to capture the inter-variable correlation within each cluster.
-II-C Spatial Depth based Similarity Matrix
+I-D Main Contributions and Organization of the Paper
+This paper makes the following main contributions:
+  • Efficient Depth-Based Similarity Construction: We improve the computational efficiency of constructing depth-based similarity matrices under spatial depth by developing a matrix-based implementation.
+  • DLCC Clustering Framework: We propose DLCC, a depth-based clustering framework that uses data depth to measure similarities, define robust neighborhoods, and extract exemplars (“local centers”).
+DLCC then forms clusters by grouping these local centers and propagating labels to all observations.
+  • A Complete Procedure and Empirical Evaluation: We provide a practical end-to-end procedure for DLCC, including filtering, grouping local centers and cluster formation mechanisms, together with guidance on strategy/parameter selection.
+We also provide extensive experiments on synthetic and real-world data sets.
+The remainder of the paper is organized as follows: Section II introduces the depth-based similarity matrix.
+Section III provides a detailed description of the DLCC framework.
+In Section IV, we present details about parameter selections.
+Section V showcases experimental results of DLCC using both synthetic and real-world datasets.
+Lastly, Section VI concludes the paper.
+Figure 1 provides the flowchart of the DLCC method, and Table I summarizes the key notations used in this paper to help readers navigate its contents more easily.
+Additional graphs providing further insights into our methods are included in the supplementary file.
+DatasetDepth-basedsimilarity matrixSection IINeighborhoods &Local centersSection III Filtered centersSection III-AGroups of filtered centersSection III-A &\& Section IV-BTemporaryclustersSection III-BUnlabelled pointsSection III-BFinal clustersclassification
+Figure 1: Flowchart for the DLCC algorithm, detailing the sections that include the key contents represented in the nodes.
+TABLE I: Summary of the main notations
+Notation Description
+$n$ and $d$ Number of observations and dimensions, respectively;
+$D_{\mathrm{SD}}(\mathbf{z}|\mathbf{X})$ Spatial depth of $\mathbf{z}$ w.r.t. $\mathbf{X}$;
+$\mathbf{X}$ A given dataset;
+$\mathbf{X}_{Ri}$ Dataset reflected around $\mathbf{x}_{i}$, i.e., $\mathbf{X}_{Ri}=\mathbf{X}\cup\{2\mathbf{x}_{i}-\mathbf{x}_{j}\,|\,j\neq i\}$;
+$\mathbf{E}$, $\mathbf{\tilde{E}}$ Matrices of unit vector sums from $\mathbf{X}$ and reflected points to $\mathbf{X}$;
+$\mathbb{S}$ Depth-based similarity matrix, with $\mathbb{S}_{i,j}$ its $ij$th entry;
+$s$ Neighborhood size;
+$\mathbf{N}_{i}$ Depth-based neighborhood of $\mathbf{x}_{i}$;
+$r_{i}(\mathbf{x}_{j})$ Depth rank of $\mathbf{x}_{j}$ within $\mathbf{N}_{i}$;
+$c$, $c$, $\mathbf{C}$ A local center, a certain set of local centers, and the set of all filtered centers, respectively;
+$\mathbb{M}$ Local center neighborhoods’ similarity matrix, with $\mathbb{M}_{i,j}$ its $ij$th entry;
+$\delta$ Similarity threshold for local center grouping ($\mathbb{M}$-based);
+$\mathbf{g}$ A group of local centers;
+$\mathbf{U}_{\mathbf{g}}$ Unique neighbors of group $\mathbf{g}$;
+$K$ and $k$ Number of clusters and cluster index;
+$\mathcal{X}_{k}$ Observations in cluster $k$;
+$f$ Frequency of a local center;
+$\mathcal{P}$ Cumulative proportion of neighbors of local centers to all points;
+$\mathcal{S}_{k}$, $\mathcal{\hat{S}}_{k}$ Score pools for accepting and leaving points unlabelled in cluster $k$;
+II Spatial Depth based Similarity Matrix
 A spatial-depth-based similarity matrix can be constructed from calculating $D_{\mathrm{SD}}(\mathbf{x}_{j}|\mathbf{X}_{Ri})$ for all $i,j=1,2,\ldots,n$.
 One advantage of using this depth measure is that the shape of the resulting neighborhoods is determined by the local geometry of the sample and does not rely on any parameters.
-This is a potentially very interesting advantage, but it does entail the computationally intensive process of repeatedly calculating depth values for the $n$ sample points with respect to $n$ distinct datasets $\mathbf{X}_{R1},\mathbf{X}_{R2},\ldots,\mathbf{X}_{Rn}$.
+This is a potentially very interesting advantage, but it does entail the computationally intensive process of repeatedly calculating depth values for the $n$ sample points w.r.t.
+$n$ distinct datasets $\mathbf{X}_{R1},\mathbf{X}_{R2},\ldots,\mathbf{X}_{Rn}$.
 Hence, we delve into certain properties of spatial depth in the current context (working with reflected samples) and introduce an efficient approach to calculate the similarity matrix.
 In order to compute the spatial depth of a point $\mathbf{x}$ from (3), we note that it is necessary to find the vectors from all sample points to $\mathbf{x}$ and their corresponding norms.
-In any of the combined dataset $\mathbf{X}_{Ri}$, the $n$ points from the original dataset $\mathbf{X}$ remain unchanged.
+In any of the combined datasets $\mathbf{X}_{Ri}$, the $n$ points from the original dataset $\mathbf{X}$ remain unchanged.
 Consequently, we can retain a $n\times d$ matrix, denoted as $\mathbf{E}$, to prevent redundant computations.
-Here, the $q^{\text{th}}$ row is the sum of the unit vectors from other observations in $\mathbf{X}$ to $\mathbf{x}_{q}$, expressed as:
+Here, the $q$th row is the sum of the unit vectors from other observations in $\mathbf{X}$ to $\mathbf{x}_{q}$, expressed as:
 $\mathbf{E}_{q}=\sum_{\mathbf{x}\in\mathbf{X}\setminus\{\mathbf{x}_{q}\}}\frac{\mathbf{x}_{q}-\mathbf{x}}{\|\mathbf{x}_{q}-\mathbf{x}\|}.$
@@
   • The $n\times d$ matrix $\mathbf{B}_{i}$ consists of points in $\mathbf{X}$, each multiplied by the corresponding column sum of $\mathbf{R}_{i}$.
-With these, $\mathbf{\tilde{E}}_{i}$ is expressed as: $\mathbf{\tilde{E}}_{i}=\mathbf{B}_{i}-\mathbf{R}^{\prime}_{i}\mathbf{\tilde{X}}_{i}.$ Through the above procedure, $\mathbf{\tilde{E}}_{i}$ can be derived efficiently.
-For instance, in a $2000\times 600$ dataset, while $\mathbf{E}$ requires approximately two seconds to compute, $\mathbf{\tilde{E}}_{i}$ only takes around $0.1$ second per loop — run on a PC with Intel(R) Core(TM) i7-11700K processor (3.6 GHz) with sufficiently large physical memory, Matlab R2023a.
+With these, $\mathbf{\tilde{E}}_{i}$ is expressed as: $\mathbf{\tilde{E}}_{i}=\mathbf{B}_{i}-\mathbf{R}^{\prime}_{i}\mathbf{\tilde{X}}_{i}.$ Through the above procedure, $\mathbf{\tilde{E}}_{i}$ can be computed efficiently.
+For example, on a $2000\times 600$ dataset, computing $\mathbf{E}$ takes approximately $2$ seconds, whereas each $\mathbf{\tilde{E}}_{i}$ requires about $0.1$ seconds per loop.
+All experiments in this paper were conducted on a PC equipped with an Intel(R) Core(TM) i7-11700K CPU (3.6 GHz) and sufficient physical memory, using Matlab R2023a.
 Having obtained $\mathbf{E}$ and $\mathbf{\tilde{E}}$, the reflection spatial depth values of any $\mathbf{x}_{q}$ in $\mathbf{X}$ based on $\mathbf{X}_{Ri}$ can be computed via
 $D_{\mathrm{SD}}(\mathbf{x}_{q}|\mathbf{X}_{Ri})=1-\left\|\frac{\mathbf{E}_{q}+\mathbf{\tilde{E}}_{i,q}}{2n-1}\right\|.$ (4)
+From a theoretical perspective, computing depth-based similarities requires $O(n^{3})$ operations, since evaluating spatial depth for $n$ points incurs $O(n^{2})$ cost and is repeated over $n$ reflected datasets.
+While the proposed implementation does not reduce this complexity, it substantially accelerates computation in practice through efficient reuse of the distance relationship and matrix-level computations.
+Figure 2 illustrates the empirical runtime comparison on synthetic datasets with fixed dimension $d=500$ and sample size ranging from $60$ to $3600$, reported on a logarithmic scale.
+For a fair comparison, we implemented the baseline in Matlab following the same algorithmic logic as the local depth in the DepthProc R package [42].
+As shown, the proposed approach consistently achieves significant runtime reductions.
+Figure 2: Running time comparison (log scale) between the proposed method and the traditional construction.
 III Depth-based Local Center Clustering
@@
 Definition 1 (DLCC Neighborhoods and Local Ranks).
-Neighbors and Neighborhoods: Given an integer parameter $s$, the neighbors of $\mathbf{x}_{i}\in\mathbf{X}$ are defined as the set of $s-1$ points $\mathbf{x}_{q}\in\mathbf{X}\setminus\{\mathbf{x}_{i}\}$ with the highest depth values with respect to the dataset $\mathbf{X}_{Ri}$, i.e., the points with the largest $D\left(\mathbf{x}_{q}\mid\mathbf{X}_{Ri}\right)$.
-The neighborhood of $\mathbf{x}_{i}$, denoted as $\mathbf{N}_{i}$, is thus formed by including $\mathbf{x}_{i}$ and its $s-1$ neighbors, constituting a subset of $\mathbf{X}$ with size $s$.
-Local Depth [37] and Local Ranks: The local depth of an observation $\mathbf{x}_{i}$ is its depth value within its neighborhood $\mathbf{N}_{i}$.
-The local rank of an observation $\mathbf{x}_{j}$ within the neighborhood $\mathbf{N}_{i}$ is defined as:
+Neighbors and Neighborhoods: Given an integer parameter $s$, the neighbors of $\mathbf{x}_{i}\in\mathbf{X}$ are defined as the set of s1s-1 points $\mathbf{x}_{q}\in\mathbf{X}\setminus\{\mathbf{x}_{i}\}$ with the highest depth values w.r.t.
+the dataset $\mathbf{X}_{Ri}$, i.e., the points with the largest $D\left(\mathbf{x}_{q}\mid\mathbf{X}_{Ri}\right)$.
+The neighborhood of $\mathbf{x}_{i}$, denoted as $\mathbf{N}_{i}$, is thus formed by including $\mathbf{x}_{i}$ and its s1s-1 neighbors, constituting a subset of $\mathbf{X}$ with size $s$.
+Local Ranks: The local rank of an observation $\mathbf{x}_{j}$ within the neighborhood $\mathbf{N}_{i}$ is defined as:
 $r_{i}(\mathbf{x}_{j})=1+\sum_{\mathbf{x}_{q}\in\mathbf{N}_{i}}\mathbb{I}\left[D\left(\mathbf{x}_{j}\mid\mathbf{N}_{i}\right)<D\left(\mathbf{x}_{q}\mid\mathbf{N}_{i}\right)\right],$ (5)
@@
 If $r_{i}(\mathbf{x}_{j})=1$, then $\mathbf{x}_{j}$ is characterized as the center of the subset $\mathbf{N}_{i}$.
-We define central points of $\mathbf{N}_{i}$ as a local center.
+We define the central point of $\mathbf{N}_{i}$ as a local center.
 These local centers can be understood as representative points for the entire dataset $\mathbf{X}$.
-Figure 3: Local center example ($s=50$) for the first two PCs of the Iris dataset, where a focus point (green triangle) with its neighbors (pink points) and the corresponding local center (blue diamond) are highlighted.
+Figure 3: Local center example ($s=50$) in the two–PC visualization of the Iris dataset (see Section V for dataset description).
+A focus point (green triangle), its neighbors (pink points), and the resulting local center (blue diamond) are highlighted.
 Notes:
@@
 Hence, it is crucial to filter out unnecessary local centers before grouping them.
-The remaining centers after filtering are denoted as filtered centers.
-DLCC introduces two strategies designed to address diverse types of clustering challenges. The conditions for grouping the filtered centers under these two strategies are summarized as follows:
+The local centers that remain after going through this filtering process are denoted as filtered centers.
+The filtering of local centers will be discussed in details in Section III-A.
+DLCC introduces two strategies for grouping filtered local centers, designed to address different types of clustering challenges.
+The conditions for grouping the filtered centers under these two strategies are summarized as follows.
 Definition 2.
@@
 Define the similarity between any two neighborhoods as the proportion of overlapping observations among them.
-Two grouping strategies are as follows:
+For any two neighborhoods $\mathbf{N}_{c}$ and $\mathbf{N}_{c^{\prime}}$, define their similarity as
+$\mathrm{sim}(\mathbf{N}_{c},\mathbf{N}_{c^{\prime}})=\frac{|\mathbf{N}_{c}\cap\mathbf{N}_{c^{\prime}}|}{s},$ (6)
+where $|\cdot|$ denotes set cardinality and $s$ is the neighborhood size.
+A grouping is a partition $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$ of the centers $\{c_{1},\ldots,c_{A}\}$.
+Two grouping strategies are defined as follows:
   • Min strategy: for any $c\in\mathbf{g}_{i}$:
-[…]
-[…]
+    $\min_{c^{*}\in\mathbf{g}_{i}}\mathrm{sim}(\mathbf{N}_{c},\mathbf{N}_{c^{*}})$ $\!\!\!\!>\!\!\!\!$ $\max_{j\neq i}\min_{c^{\dagger}\in\mathbf{g}_{j}}\mathrm{sim}(\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}})$ (7)
+    $\max_{c^{*}\in\mathbf{g}_{i}\setminus\{c\}}\mathrm{sim}(\mathbf{N}_{c},\mathbf{N}_{c^{*}})$ $\!\!\!\!>\!\!\!\!$ $\max_{c^{\dagger}\notin\mathbf{g}_{i}}\mathrm{sim}(\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}}).$ (8)
   • Max strategy: Given a parameter $\delta$, termed as the similarity threshold, $\forall$ $c$ in group $\mathbf{g}$, $\exists c^{*}$ in the same group s.t.
-$\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{*}})\,>\delta$, and for $\forall c^{\dagger}\notin\mathbf{g}$, $\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}})\,\leq\delta$.
+$\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{*}})\,>Q$, and for $\forall c^{\dagger}\notin\mathbf{g}$,
+    $\mathrm{sim}(\,\mathbf{N}_{c},\mathbf{N}_{c^{\dagger}})\,\leq\delta.$
 Note that, “group” refers specifically to a set of filtered centers and should not be confused with “cluster”.
-Some insights into the strategy chosen for grouping local centers are important in applications.
-Under the min strategy, DLCC attempts to find centroids of clusters from filtered centers.
-The variance in the number of filtered centers across different groups allows for better flexibility than traditional center-based clustering methods, especially when dealing with clustering problems that have overlapping regions.
-The min strategy tends to be more suitable when the sizes of all clusters are comparable.
-The max strategy assumes each cluster can be built by connecting multiple finite convex shapes created by filtered centers.
-Filtered centers in the same group are connected with a chain of similarities.
-For example, if two filtered centers $c_{y}$ and $c_{z}$ are in the same group, then there must exist a chain of filtered centers, $c_{p_{1}},c_{p_{2}},\ldots,c_{p_{h}}$, where $c_{p_{1}}=c_{y}$, $c_{p_{h}}=c_{z}$, and the similarity between any two adjacent centers’ neighborhoods in that chain is larger than $\delta$.
-Generally, the parameter $s$ is small under the max strategy.
-When using Mahalanobis depth to construct the similarity matrix with a small $s$, the max strategy is less sensitive to variations in the matrix.
-In such cases, a globally estimated similarity matrix often suffices, similar to DBCA.
-However, unlike DBCA, which builds clusters with chains of fixed-shape ellipses centered at each observation, DLCC with the max strategy considers only chains of filtered centers, with a fixed number of neighbors rather than a fixed region.
-Like density-based methods, the max strategy works well for disjoint clusters and can handle non-convex clusters of varying sizes.
+Some insights into the strategy chosen for grouping local centers are described in Section IV-A.
 Definition 3 (Unique neighbor).
@@
 Indeed, we could potentially consider any classification method at this stage, depending on the specific requirements of the task at hand.
-DLCC incorporates three built-in classification methods: maximum-depth classifier (MDC), K-nearest neighbors (KNN) based on the similarity matrix, and random forests (RF).
-Lastly, self-improvement procedures of DLCC are discussed.
-Although the clustering problem is converted to a classification problem in the final stage, it has some special features, because the observations that are already labelled may still be updated.
-Two logical parameters are introduced here:
-  • maxdepth: This parameter dictates whether DLCC will continue iterations until all observations in the assigned cluster attain the highest depth values.
-  • ifloop: This parameter, only applicable under the min strategy, determines whether only a single filtered center should remain in each cluster.
-Let the $K$ clusters constructed so far be denoted by $\mathcal{X}_{k}$, where $k=1,2,\ldots,K$.
-For any observation $\mathbf{x}_{j}$, $D(\mathbf{x}_{j}\mid\mathcal{X}_{k})$ corresponds to the depth of this observation within cluster $\mathcal{X}_{k}$.
-If the depth value of $\mathbf{x}_{j}$ in the cluster it is currently assigned to does not match $\max_{k}D(\mathbf{x}_{j}\mid\mathcal{X}_{k})$, and if the “maxdepth” parameter is set to TRUE, $\mathbf{x}_{j}$ will be relocated.
-This procedure loops until all observations are assigned to the cluster where they obtain the largest depth value.
-Regarding the min strategy, it can be adapted to resemble a centroid-based clustering method.
-If the “ifloop” parameter is TRUE, the clustering algorithm updates clusters using only the deepest filtered center within each cluster.
-However, other filtered centers are not disregarded.
-They may become the deepest center for a cluster in future iterations, as clustering results are reassessed after each update.
-This iterative process continues, regenerating clusters and potentially re-evaluating the roles of filtered centers, until the clustering outcome stabilizes.
+DLCC incorporates two built-in classification methods: $k$-nearest neighbors (kkNN) based on the similarity matrix $\mathbb{S}$, and random forests (RF).
 III-A Local Center Selection and Grouping
@@
 $\mathcal{P}_{t}=\frac{1}{n}\left|\bigcup_{p=1}^{t}\mathbf{N}_{c_{p}}\right|,$ (9)
-where ||$\cdot$|| ⋅ | represents the cardinality of a set.
-The filtering procedure consists of two steps: (1) selecting a suitable frequency cutoff, and (2) adjusting the grouping results to satisfy the min strategy defined in Definition 2.
-According to Definition 2, for fixed local centers, the grouping result satisfying the max strategy is unique, whereas multiple solutions may satisfy the min strategy.
-Thus, a reasonable initial grouping method that provides high-quality results and estimates the number of groups is essential.
-We summarize the designed grouping method in Algorithm 1.
-This method leverages local depths and the similarities between neighborhoods of local centers.
-Here, local depth values are computed during defining local centers, and the similarity matrix for neighborhoods of local centers is denoted as $\mathbb{M}$, with $\mathrm{sim}(\mathbf{N}_{c_{i}},\mathbf{N}_{c_{j}})$ as its $ij$-th entry.
-Algorithm 1 Local depth and similarity-based grouping
-1:Local centers $c_{1},\ldots,c_{T}$, corresponding top $s/2$ neighbors, local depth values (LD), similarity matrix $\mathbb{M}$, and a similarity threshold range $[0.4,0.75]$.
-2:Grouping results $\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{G}\}$.
-3:For each local center $c_{i}$, set valid neighbors $\mathrm{VN}_{i}$ as local centers in its top $s/2$ neighbors (excluding itself).
-4:Initialize mean similarities and closest neighbors: MS=zero array of size $\mathrm{MS}=\text{zero array of size }T$, $\mathrm{CN}=1{:}T$
-5:for $t=1$ to $T$ do
-6:     Compute mean similarity: $\mathrm{MS}_{t}=\frac{\sum_{c_{j}\in\mathrm{VN}_{t}}\mathbb{M}_{tj}}{|\mathrm{VN}_{t}|}$.
-7:     Find the closest neighbor: $\mathrm{CN}_{t}=\arg\max_{j\neq t}\mathbb{M}_{tj}$.
-8:end for
-9:Set similarity threshold $\delta_{i}$ for each local center $c_{i}$:
-$\delta_{i}=\min(\mathrm{MS}_{i},\mathrm{MS}_{\mathrm{CN}_{i}})$
-and if $\delta_{i}\notin[0.4,0.75]$ adjust it to the nearest bound.
-10:Initialize stable centers $\mathrm{SC}=\emptyset$.
-11:for $t=1$ to $T$ do
-12:     if All $c_{j}\in\mathrm{VN}_{t}$ with $\mathrm{sim}(\mathbf{N}_{c_{t}},\mathbf{N}_{c})\geq\delta_{t}$ have $\text{LD}_{j}<\text{LD}_{t}$ then
-13:         Add $c_{t}$ to $\mathrm{SC}$.
+where $|\cdot|$ represents the cardinality of a set.
+The filtering procedure under the min strategy consists of two stages: (1) selecting a frequency cutoff and forming an grouping, and (2) refining the grouping to satisfy the min strategy in Definition 2.
+For a fixed set of local centers, the max strategy yields a unique grouping, whereas the min strategy may admit multiple feasible groupings.
+Thus, we use a heuristic procedure that performs filtering and grouping simultaneously.
+Stage (1) is implemented by Algorithms 2 with grouping detailed in Algorithm 1, and stage (2) by the group trimming process in Algorithm 3.
+Implementation details are given next and may be skipped on a first reading without loss of the main ideas of the paper.
+Our method leverages a stability criterion for evaluating a local center $c_{i}$, termed average neighbor-induced depth (AND).
+Recall $\mathbb{S}_{j,i}=D_{\mathrm{SD}}(\mathbf{x}_{i}\mid\mathbf{X}_{Rj})$ denote the depth-based similarity from $\mathbf{x}_{j}$ to $\mathbf{x}_{i}$, we define AND as
+$\mathrm{AND}(\mathbf{x}_{i})\;=\;\frac{1}{s}\sum_{j\in\mathbf{N}_{i}}\mathbb{S}_{j,i}.$ (10)
+Intuitively, $\mathrm{AND}(i)$ aggregates the depth of $\mathbf{x}_{i}$ computed w.r.t.
+reflected datasets centered at its neighbors.
+It tends to be large only when $\mathbf{x}_{i}$ is consistently central relative to the collection of reflected data sets induced by its neighborhood.
+One might ask why local depth is not adopted here.
+As noted in the literature [40, 43], local depth may assign relatively high values to points located between two distributions, which can be central in a geometric sense but not necessarily lie in regions of high data concentration.
+In clustering tasks, such behavior may lead to the selection of misleading representative centers.
+We next identify stable centers from the frequency-sorted local-center candidates.
+For each candidate $c_{t}$, we consider its top $\lfloor s/2\rfloor$ depth-based neighbors (i.e., half points in $\mathbf{N}_{c}$), and two candidates are regarded as mutual neighbors if they appear in each other’s top $\lfloor s/2\rfloor$ neighbors.
+The set of mutual neighbors of local center $c$ is denoted as $\mathrm{MN}(c)$
+A candidate $c_{t}$ is declared a stable center if its AND is not smaller than that of any of its mutual neighbors.
+Let $\mathrm{SC}$ be the set of stable centers.
+We take the cutoff index as the last stable center in the frequency order and retain the local centers up to this index.
+To avoid premature truncation, we further check the cumulative proportion $\mathcal{P}_{t}$.
+If $\mathcal{P}_{t}<0.75$, we either (1) simply extend the frequency cutoff until the coverage first exceeds $0.75$ when the gap is large, or (2) when the coverage is close to $0.75$, greedily add a small number of additional candidates that maximize the product of marginal coverage gain and AND, thereby efficiently increasing coverage without introducing many low-quality centers.
+Let $\mathbf{C}$ denote the retained local-center set, and let $\mathbb{M}$ collect the pairwise neighborhood similarities among centers in $\mathbf{C}$, with entries $\mathbb{M}_{i,j}=\mathrm{sim}(\mathbf{N}_{c_{i}},\mathbf{N}_{c_{j}})$ computed by (6).
+For each stable center $c_{t}$, we form a group by aggregating centers whose similarities to $c_{t}$ exceed an adaptive threshold $\delta_{t}$.
+Throughout the algorithm, all such similarity thresholds are constrained to lie in $\delta_{\min},\delta_{\max}]$ and in all experiments, we set $(\delta_{\min},\delta_{\max})=(0.4,0.7)$ to avoid overly loose or overly stringent grouping.
+$\delta_{t}=\max\!\Bigl\{$ $\min_{c_{u}\in\mathrm{MN}(c_{t})\cap\mathbf{C}}\mathbb{M}_{t,u}\;\text{s.t.}\;\mathbb{M}_{t,u}>\max_{c_{v}\in\mathbf{C}\setminus\mathrm{MN}(c_{t})}\mathbb{M}_{t,v},$ (11)
+$\delta_{\min}\Bigr\}.$
+This definition ensures that $\delta_{t}$ retains only mutual neighbors of $c_{t}$ while excluding non-mutual candidates, subject to a conservative lower bound that prevents overly loose thresholds.
+The initial grouping may not cover all local centers in $\mathbf{C}$.
+We therefore perform a second pass on the unassigned candidates, applying the same rules for defining extra stable centers but restricting mutual-neighbor relations to the unassigned set.
+When a candidate center $c_{t}$ has no remaining mutual neighbors, we assess its stability by whether its neighborhood $\mathbf{N}_{c_{t}}$ contributes sufficient new coverage beyond the neighborhoods of all other local centers in $\mathbf{C}$.
+Specifically, we define $c_{t}$ as a stable center only if
+$\left|\mathbf{N}_{c_{t}}\setminus\bigcup_{c_{u}\in\mathbf{C}\setminus\{c_{t}\}}\mathbf{N}_{c_{u}}\right|>\frac{s}{3}.$ (12)
+We then use the same rule for creating groups of newly defined stable centers.
+Preliminary groups formed around stable centers may overlap.
+We resolve such overlaps and decide possible merges using a lightweight two-stage heuristic.
+Consider two groups $\mathbf{g}_{p}$ and $\mathbf{g}_{q}$ with nonempty intersection (the extension to overlaps involving more than two groups is similar).
+We define their cores by removing the overlapped centers:
+$\mathrm{core}(\mathbf{g}_{p})=\mathbf{g}_{p}\setminus(\mathbf{g}_{p}\cap\mathbf{g}_{q}),\qquad\mathrm{core}(\mathbf{g}_{q})=\mathbf{g}_{q}\setminus(\mathbf{g}_{p}\cap\mathbf{g}_{q}).$
+Each overlapped center $c\in\mathbf{g}_{p}\cap\mathbf{g}_{q}$ is hard-assigned to the group with the larger minimum neighborhood similarity between $c$ and the corresponding core; if both are below $\delta_{\min}$, $c$ is discarded for ambiguity.
+After this, we test whether two previous overlapped groups should be merged. We merge $\mathbf{g}_{p}$ and $\mathbf{g}_{q}$ if
+$\max_{c_{u}\in\mathbf{g}_{p},\,c_{v}\in\mathbf{g}_{q}}\mathbb{M}_{u,v}\;\geq\;\min\!\{\min_{c_{u},c_{v}\in\mathbf{g}_{p}}\mathbb{M}_{c_{u},c_{v}},\;\min_{c_{u},c_{v}\in\mathbf{g}_{q}}\mathbb{M}_{c_{u},c_{v}}\},$ (13)
+and
+$\min_{c_{u},c_{v}\in\mathbf{g}_{p}\cup\mathbf{g}_{q}}\mathbb{M}_{u,v}\geq\delta_{\min}.$ (14)
+In implementation, the within-group minima in (13) are clipped to $\delta_{\min},\delta_{\max}]$.
+Lastly, if some local centers in $\mathbf{C}$ remain unassigned, we assign each such center $c_{t}$ to
+$g^{*}_{t}=\arg\max_{g\in\mathcal{G}(c_{t})}\;\min_{c_{u}\in\mathbf{g}_{g}}\mathbb{M}_{t,u},$ (15)
+where
+$\mathcal{G}(c_{t})=\left\{g:\min_{c_{u}\in\mathbf{g}_{g}}\mathbb{M}_{t,u}\geq\min\!\left(\min_{c_{u},c_{v}\in\mathbf{g}_{g}}\mathbb{M}_{u,v},\,\delta_{\max}\right)\right\},$
+otherwise $c$ is dropped from $\mathbf{C}$.
+Algorithm 1 AND-based stable-center selection and grouping
+1:Candidate local centers, neighborhoods $\{\mathbf{N}_{c}\}$ (size $s$), pointwise similarities $\mathbb{S}$, AND $\mathrm{AND}(\cdot)$ in (10), threshold bounds $\delta_{\min},\delta_{\max}]$, and sample size $n$.
+2:Retained centers $\mathbf{C}$ and groups $\mathbf{G}=\{\mathbf{g}_{1},\ldots,\mathbf{g}_{\hat{K}}\}$.
+3:Mutual-neighbor graph: for each candidate $c$, keep its top $\lfloor s/2\rfloor$ candidate neighbors under $\mathbb{S}$; define $\mathrm{MN}(c)$ as mutual neighbors.
+4:Stable centers: $\mathrm{SC}\leftarrow\{c_{t}:\mathrm{AND}(c_{t})\geq\mathrm{AND}(c_{u}),\ \forall c_{u}\in\mathrm{MN}(c_{t})\}$.
+5:Frequency cutoff: $t_{0}\leftarrow\max\{t:c_{t}\in\mathrm{SC}\}$; initialize $\mathbf{C}\leftarrow\{c_{1},\ldots,c_{t_{0}}\}$.
+6:Compute $\mathcal{P}(\mathbf{C})=\frac{1}{n}\big|\cup_{c\in\mathbf{C}}\mathbf{N}_{c}\big|$ and if $\mathcal{P}(\mathbf{C})<0.75$ update $\mathbf{C}$ using the two-case rule in the text until $\mathcal{P}(\mathbf{C})\geq 0.75$.
+7:Build neighborhood similarity matrix $\mathbb{M}$ over $\mathbf{C}$.
+[…]
+9:for each $c_{t}\in\mathrm{SC}$ do
+10:  Compute $\delta_{t}$ by (11) and form $\mathbf{g}_{t}\leftarrow\{c_{u}\in\mathbf{C}:\mathbb{M}_{t,u}\geq\delta_{t}\}$.
+11:  $\mathcal{G}\leftarrow\mathcal{G}\cup\{\mathbf{g}_{t}\}$.
+12:end for
+13:Second pass: $\mathcal{L}\leftarrow\mathbf{C}\setminus\bigcup_{\mathbf{g}\in\mathbf{G}}\mathbf{g}$.
+14:if $\mathcal{L}\neq\emptyset$ then
+15:  Obtain $\mathrm{SC}_{2}$ by repeating Line 2 on $\mathcal{L}$ (mutual neighbors are defined within $\mathcal{L}$ only); if a candidate $c\in\mathcal{L}$ has no mutual neighbor in $\mathcal{L}$, add it to $\mathrm{SC}_{2}$ only if (12) holds.
+16:  For each $c\in\mathrm{SC}_{2}$, repeat Lines 7–10 to form groups and append them to $\mathbf{G}$.
+17:end if
+18:Check overlap: apply the two-stage heuristic in the text (core-based hard assignment, then merge test (13) and (14)).
+19:If any centers in $\mathbf{C}$ remain unassigned, assign them by (15).
+[…]
+In the above procedure, we target a candidate set whose cumulative neighborhood coverage is at least $0.75$.
+When $\{c_{1},\ldots,c_{T}\}$ yields $\mathcal{P}_{T}<0.75$, the retained local centers may be overly concentrated in overlapping regions, resulting in insufficient coverage.
+To avoid discarding informative but low-frequency candidates, we expand the candidate set by iteratively adding dropped centers from the initial filtering step (i.e., those with $f=1$ or $r_{i}(\mathbf{x}_{i})>2$) until the target coverage is reached.
+Details are summarized in Algorithm 2.
+Algorithm 2 Local-center filtering for the min strategy
+1:Frequency-sorted candidates $\{c_{1},\ldots,c_{T}\}$, full local-center set $\mathcal{C}_{\mathrm{all}}$, neighborhoods $\{\mathbf{N}_{c}\}$ (size $s$), AND scores $\mathrm{AND}(\cdot)$, and sample size $n$.
+2:Grouping results $\mathbf{G}$ and retained centers $\mathbf{C}$.
+3:Compute $\mathcal{P}_{\mathrm{all}}=\mathcal{P}(\mathcal{C}_{\mathrm{all}})$ and set $\tau\leftarrow\min\{0.75,\;0.9\,\mathcal{P}_{\mathrm{all}}\}$.
+4:if $\mathcal{P}(\mathbf{C})<\tau$ then
+5:  Initialize $\mathbf{C}\leftarrow\{c_{1},\ldots,c_{T}\}$. and $\mathcal{N}\leftarrow\bigcup_{c\in\mathbf{C}}\mathbf{N}_{c}$.
+6:  while $\mathcal{P}(\mathbf{C})<\tau$ do
+7:    For each $u\in\mathcal{C}_{\mathrm{all}}\setminus\mathbf{C}$, compute $\Delta(u)=|\mathbf{N}_{u}\setminus\mathcal{N}|$.
+8:    Select $u^{*}=\arg\max_{u}\ \Delta(u)\cdot\mathrm{AND}(u)$.
+9:    Update $\mathbf{C}\leftarrow\mathbf{C}\cup\{u^{*}\}$ and $\mathcal{N}\leftarrow\mathcal{N}\cup\mathbf{N}_{u^{*}}$.
+10:  end while
+11:  Apply Algorithm 1 with the candidate set fixed to $\mathbf{C}$ (i.e., skip frequency cutoff stage).
+12:else
+13:  Apply Algorithm 1 on $\{c_{1},\ldots,c_{T}\}$.
 14:end if
-15:end for
-16:Set $G=|\mathrm{SC}|$.
-17:Assign each remaining local center to the group of its most similar stable center using $\mathbb{M}$.
-18:Return groups $\{\mathbf{g}_{1},\mathbf{g}_{2},\ldots,\mathbf{g}_{G}\}$.
-Algorithm 1 uses local depth to assess the centrality of each center and employs neighborhood similarity to evaluate the relationship between centers.
-The similarity threshold is determined automatically, eliminating the need for manual parameter selection.
-Specifically, the threshold for each center $c$ is set as the minimum of the mean similarity between $c$ and its closest neighbor.
-This approach provides a center-specific threshold while ensuring that centers with similar neighborhoods share comparable thresholds.
-The empirical threshold range $[0.4,0.75]$ prevents the values from being excessively small or large.
-Moreover, the algorithm naturally estimates the number of groups, which corresponds to the number of estimated clusters.
-A limitation of Algorithm 1 is if multiple clusters overlap with each other, the local center defined in the overlapped region may have high local depth values.
-These local centers can be removed by analyzing the number of unique neighbors (Definition 3) within each group, as groups containing local centers from overlapped regions tend to have very few unique neighbors.
-Specifically, we consider a group of centers should be dropped if the number of unique neighbors of this group is less than $s/10$.
-Additionally, local centers highly similar to those in the identified group should also be dropped to avoid ambiguity.
-With the grouping of local centers, we determine candidates for the cutoff in the order of frequency. The first candidate index is defined as
-$\max\left(\arg\min_{t}\{\mathcal{P}_{t}\geq 0.75\},\max_{\mathbf{g}}\{\arg\min_{t:c_{t}\in\mathbf{g}}t\}\right),$ (12)
-where the first term ensures the final cumulative proportion $\mathcal{P}_{t}$ exceeds $0.75$ to prevent excessive elimination of local centers, and the second term guarantees at least one local center from each group is retained.
-Additional candidates include local centers after the begin index that provide complementary information beyond those in the same group with higher frequency.
-Specifically, a candidate index $i$ must satisfy two conditions: First for all $c_{j}$ where $j\leq i$ and $c_{j}$ belongs to the same group $\mathbf{g}$ as $c_{i}$,
-$\sum_{j}\mathbb{M}_{ij}=\min_{l\leq i,c_{l}\in\mathbf{g}}\sum_{j}\mathbb{M}_{lj},$ (13)
-This ensures $c_{i}$ is the least similar to other local centers with higher frequencies in its group. Second, for all $c_{v}$ where $v<i$ and $c_{v}\notin\mathbf{g}$,
-$\max_{v}\mathbb{M}_{iv}\leq\min_{j<i,c_{j}\in\mathbf{g}}\max_{v}\mathbb{M}_{jv},$ (14)
-ensuring $c_{i}$ is less similar to points in other groups compared to any existing local center in its own group.
-In essence, these candidates are distant from local centers in other groups while also comparatively distinct within their own group, making them potential sources of complementary information.
-Lastly, we evaluate each candidate in ascending order and count the number of “bad points” between candidates, defined as local centers that fail to satisfy Equations (9) or (10) after removing local centers in groups with indices larger than the candidate.
-If the number of bad points between two consecutive candidates $c_{i}$ and $c_{i^{\prime}}$ exceeds $1$, we stop at $c_{i}$.
-In the above setting, the initial candidate should have a cumulative proportion of at least $0.75$.
-However, if $\mathcal{P}_{T}<0.75$, it may indicate that overlapping clusters cause local centers to concentrate in shared regions rather than capturing true subcluster centers.
-The robustness of the depth median (e.g.
-SD’s breakdown point is $0.5$) biases local centers toward overlapping regions when most points in a cluster have neighborhoods with similarities exceeding $0.5$ to points in other clusters.
-To address this, if $\mathcal{P}_{T}<0.75$, we retain only centers dissimilar to those in other groups after grouping $c_{1},\ldots,c_{T}$ using Algorithm 1.
-For $c\in\mathbf{g}$, we check:
-$\frac{\sum_{j:c_{j}\in\mathbf{g}\setminus c_{i}}\mathbb{M}_{ij}}{|\mathbf{g}|-1}-\max_{v:c_{v}\notin\mathbf{g}}\mathbb{M}_{iv}>0.$ (15)
-Then, we iteratively add back local centers from the dropped ones, including those with $f=1$ or $r_{i}(\mathbf{x}_{i})>2$.
-In each step, the local center that maximizes the increase in cumulative proportion is added back.
-The process stops when the cumulative proportion exceeds $90\%$ of the cumulative proportion of all local centers.
-Algorithm 2 Local center filtering under the min strategy
-1:All local centers, their ranks, frequencies, local depth values, neighborhoods, size $s$ and similarity matrix $\mathbb{M}$.
-2:Filtered local centers and their local depth values.
-3:Perform initial filtering to obtain $\{c_{1},\ldots,c_{T}\}$.
-4:if $\mathcal{P}_{T}\geq 0.75$ then
-5:     Apply Algorithm 1 to obtain $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$.
-6:     Set Continue as True.
-7:     while Continue do
-8:         Obtain unique neighbors $\{\mathbf{U}_{\mathbf{g}_{1}},\ldots,\mathbf{U}_{\mathbf{g}_{G}}\}$.
-9:         Identify $\mathrm{di}=\arg\min_{i:|\mathbf{U}_{\mathbf{g}_{i}}|<s/10}|\mathbf{U}_{\mathbf{g}_{i}}|$.
-10:         if $\mathrm{di}\neq\emptyset$ then
-11:              Set $\delta=\max_{i,j\in\mathbf{g}_{\mathrm{di}},i\neq j}\mathbb{M}_{ij}$.
-12:              Mark centers in $\mathbf{g}_{\mathrm{di}}$ as drop centers.
-13:              Add centers $v$ outside $\mathbf{g}_{\mathrm{di}}$ to drop centers if $\max_{i:c_{i}\in\mathbf{g}_{\mathrm{di}}}\mathbb{M}_{vi}\geq\delta$.
-14:              Remove $\mathbf{g}_{\mathrm{di}}$ and drop centers, and update $G$.
-15:         else
-16:              Set Continue as False;
-17:         end if
-18:     end while
-19:     Define the first cutoff candidate $\mathrm{cc}_{1}$ as in (12).
-20:     Identify additional candidates among centers with indices greater than $\mathrm{cc}_{1}$ and satisfy (13) and (14).
-21:     Initialize bp=zero array of size |$\mathrm{bp}=\text{zero array of size }|\mathrm{cc}|$.
-22:     for $q=1$ to $|\mathrm{cc}|$ do
-23:         Remove local centers in groups with indices larger than $\mathrm{cc}_{q}$ to form $\{\mathbf{g}^{\prime}_{1},\ldots,\mathbf{g}^{\prime}_{G}\}$.
-[…]
-25:     end for
-26:     Define cutoff as $\mathrm{cc}_{p}$ where $p$ is the first index satisfying $\mathrm{bp}_{p+1}-\mathrm{bp}_{p}>1$. If no such $p$ exists, set $cc_{p}$ as the largest cutoff candidate index.
-27:else
-28:     Drop local centers from $\{c_{1},\ldots,c_{T}\}$ that do not satisfy (15).
-29:     Compute current cumulative proportion $\mathcal{P}$ and total cumulative proportion $\mathcal{P}^{\prime}$.
-30:     while $\mathcal{P}<0.9\mathcal{P}^{\prime}$ do
-31:         Add the local center maximizes the increase in $\mathcal{P}$.
-32:         Update $\mathcal{P}$.
-33:     end while
-34:end if
-35:Save remaining local centers and their local depth values.
-After Algorithm 2, we reapply Algorithm 1 to group the filtered centers.
-If any local centers fail to meet the requirements of the min strategy, Algorithm 3 is applied, which is referred to as the group trimming process.
-The number of clusters is not required, but if users specify $K$, the algorithm forces the number of groups to equal $K$.
-For simplicity, the left and right sides of inequality (9) are denoted as $\mathrm{mw}(c)$ and $\mathrm{mb}(c)$ in the algorithm, respectively.
+After Algorithm 2, we obtain groups of filtered centers, denoted by $\mathbf{G}=\{\mathbf{g}_{1},\ldots,\mathbf{g}_{\hat{K}}\}$.
+If any centers in $\mathbf{G}$ violate the min-strategy conditions (7) or (8), we apply the group trimming process in Algorithm 3.
+The number of clusters is not required; when a user specifies $K$, Algorithm 3 first adjusts the number of groups to $K$ by either dropping less reliable groups (when $\hat{K}>K$) or splitting a weakly cohesive group (when $\hat{K}<K$).
+For notational simplicity, we denote the left- and right-hand sides of (7) by $\mathrm{mw}(c)$ and $\mathrm{mb}(c)$, respectively.
 Algorithm 3 Group Trimming Process
-1:Groups of local centers $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$, similarity matrix $\mathbb{M}$, optional $K$.
-2:Trimmed groups satisfying the min strategy.
-3:if $K\neq\emptyset$ then
-4:     if $G>K$ then
-5:         Drop groups with the highest $\{\arg\min_{t:c_{t}\in\mathbf{g}}t\}$ until $G$ equals to $K$.
-6:     else
-7:         while $G<K$ do
-8:              Define a new stable center and increment $G$ by $1$triangleright Details provided in supplementary material
-9:         end while
-10:         Update $\{\mathbf{g}_{1},\ldots,\mathbf{g}_{G}\}$.
-11:     end if
-12:end if
-13:while violations of (9) or (10) exist do
-14:     while violations of (9) exist do
-15:         for each group ${\mathbf{g}}_{l}$ with violations of (9do
-16:              Drop the center $c\in{\mathbf{g}}_{l}$ which maximizes:
-$\frac{\sum_{c^{\prime}\in{\mathbf{g}}_{l}\setminus\{c\}}\left(\mathrm{mw}(c^{\prime})-\mathrm{mb}(c^{\prime})\right)}{|{\mathbf{g}}_{l}\setminus\{c\}|}.$
-17:         end for
-18:         Update groups and recompute violations of (9).
-19:     end while
-20:     if violations of (10) exist then
-21:         Drop all centers violating (10).
-22:         Update groups and recompute violations.
-23:     end if
-24:end while
+1:Groups of filtered centers $\mathbf{G}=\{\mathbf{g}_{1},\ldots,\mathbf{g}_{\hat{K}}\}$, neighborhood-similarity matrix $\mathbb{M}$, optional $K$, and optional stability scores $\mathrm{AND}(\cdot)$ for centers.
+2:Trimmed groups $\mathbf{G}$ satisfying the min strategy.
+3:if $K$ is provided then
+4:  if $\hat{K}>K$ then
+5:    For each group $\mathbf{g}$, compute $t(\mathbf{g})=\min\{t:c_{t}\in\mathbf{g}\}$.
+6:    Drop the (GK)(G-K) groups from $\mathbf{G}$ with the largest $t(\mathbf{g})$.
+7:  else if $\hat{K}<K$ then
+8:    while $\hat{K}<K$ do
+9:     Select the weakest group $\mathbf{g}_{g}$ where
+$g=\arg\min_{\ell}\{\min_{c_{u},c_{v}\in\mathbf{g}_{\ell}}\mathbb{M}_{u,v}\}.$
+10:     Choose the first prototype center $p\in\mathbf{g}_{g}$ as the most stable one (i.e., it has the highest $\mathrm{AND}$ value within $\mathbf{g}_{g}$).
+11:     Choose the second prototype $q\in\mathbf{g}_{g}\setminus\{p\}$ that maximizes
+$\left\{\min_{c_{u},c_{v}\in\mathbf{g}^{(1)}(q^{\prime})}\mathbb{M}_{u,v},\ \min_{c_{u},c_{v}\in\mathbf{g}^{(2)}(q^{\prime})}\mathbb{M}_{u,v}\right\},$
+where $\mathbf{g}^{(1)}(q^{\prime})$ and $\mathbf{g}^{(2)}(q^{\prime})$ are obtained by assigning each $c\in\mathbf{g}_{g}$ to $\{p,q^{\prime}\}$ with larger neighborhood similarity.
+12:     Replace $\mathbf{g}_{g}$ by the two split groups induced by $\{p,q\}$.
+13:     Update $\mathbf{G}$, set $\hat{K}\leftarrow\hat{K}+1$.
+14:    end while
+15:  end if
+16:end if
+17:while violations of (7) or (8) exist do
+18:  while violations of (7) exist do
+19:    for each group $\mathbf{g}_{l}\in\mathbf{G}$ with violations of (7) do
+20:     Drop the center $c\in\mathbf{g}_{l}$ that maximizes
+$\frac{1}{|\mathbf{g}_{l}\setminus\{c\}|}\sum_{c^{\prime}\in\mathbf{g}_{l}\setminus\{c\}}\bigl(\mathrm{mw}(c^{\prime})-\mathrm{mb}(c^{\prime})\bigr).$
+21:    end for
+22:    Remove any empty groups and update $\mathbf{G}$.
+23:  end while
+24:  if violations of (8) exist then
+25:    Drop all centers violating (8) and update $\mathbf{G}$.
+26:  end if
+27:end while
 III-A2 Max Strategy
 Definition 4 (Isolated Local Center).
-A local center $c_{i}$ is considered isolated if, for all $c_{j}\neq c_{i}$, the similarity between the neighborhoods associated with $c_{i}$ and $c_{j}$ is zero, i.e.,
-$\mathbb{M}_{ij}=0.$
+A local center $c_{i}$ is considered isolated if, for every other local center $c_{j}\neq c_{i}$, the neighborhood similarity satisfies
+$\mathbb{M}_{i,j}=0.$
+Here, $\mathbb{M}$ denotes the neighborhood similarity matrix over local-center candidates, where $\mathbb{M}_{i,j}$ records the similarity between the neighborhoods associated with $c_{i}$ and $c_{j}$.
 In other words, an isolated local center is one whose neighborhood does not overlap with any other neighborhoods.
-As the max strategy assumes each cluster consists of various convex shapes, if there are any isolated local centers, they will be discarded.
+As the max strategy assumes each cluster consists of various convex shapes, isolated local centers are discarded to avoid defining extremely small clusters that may in fact correspond to outliers.
 Definition 5 (Everywhere Non-Convex Cluster).
 An everywhere non-convex cluster is characterized by all of its points having non-convex neighborhoods.
-Examples include one-dimensional manifolds, such as spirals, and self-crossing curves.
+Examples of everywhere non-convex cluster include one-dimensional manifolds, such as spirals, and self-crossing curves.
 A region containing such a cluster is termed a locally non-convex region.
@@
 To assess the presence of everywhere non-convex clusters and mitigate their impact, we categorize local centers into three sets based on the outcome of prior filtering:
-  • Set $\mathbf{c}_{1}$ includes all local centers that remain after initial filtering, characterized by non-isolated local centers with $r_{i}(\mathbf{x}_{i})\leq 2$ and $f\neq 1$.
-  • Set $\mathbf{c}_{2}$ contains non-isolated local centers $c_{v}$ for which $\forall c_{i}\in\mathbf{c}_{1}$, $c_{v}$ satisfies $\mathbb{M}_{vi}=0$.
-  • Set $\mathbf{c}_{3}$ consists of local centers $c_{v}\notin\mathbf{c}_{1}$ satisfying $\exists c_{i}\in\mathbf{c}_{1}$ such that $\mathbb{M}_{vi}>0$.
-The presence of everywhere non-convex clusters is inferred if $\mathbf{c}_{2}$ is not empty. In such cases, Algorithm 4 are taken to segregate locally non-convex regions from other clusters.
+  • Set $\mathcal{C}_{1}$ includes all local centers that remain after initial filtering, characterized by non-isolated local centers with $r_{i}(\mathbf{x}_{i})\leq 2$ and $f\neq 1$.
+  • Set $\mathcal{C}_{2}$ contains non-isolated local centers $c_{v}$ for which $\forall c_{i}\in\mathcal{C}_{1}$, $c_{v}$ satisfies $\mathbb{M}_{v,i}=0$.
+  • Set $\mathcal{C}_{3}$ consists of local centers $c_{v}\notin\mathcal{C}_{1}$ satisfying $\exists c_{i}\in\mathcal{C}_{1}$ such that $\mathbb{M}_{v,i}>0$.
+The presence of everywhere non-convex clusters is inferred if $\mathcal{C}_{2}$ is not empty. In such cases, Algorithm 4 is used to segregate locally non-convex regions from other clusters.
 Algorithm 4 Segregating the locally non-convex region
-1:Sets of local centers $\mathbf{c}_{1}$, $\mathbf{c}_{2}$, and $\mathbf{c}_{3}$, neighborhood similarity matrix $\mathbb{M}$
+1:Sets of local centers $\mathcal{C}_{1}$, $\mathcal{C}_{2}$, and $\mathcal{C}_{3}$, neighborhood similarity matrix $\mathbb{M}$
 2:Set of filtered centers $\mathbf{C}$
-3:if $\mathbf{c}_{2}\neq\emptyset$ then
-4:     for $v$ in the indices of local centers in $\mathbf{c}_{3}$ do
-5:         if  $\max_{j:c_{j}\in\mathbf{c}_{2}}\mathbb{M}_{vj}>\max_{i:c_{i}\in\mathbf{c}_{1}}\mathbb{M}_{vi}$  then
-6:              $\mathbf{c}_{2}=\mathbf{c}_{2}\cup\{c_{v}\}$
-7:              for $i$ in the indices of local centers in $\mathbf{c}_{1}$ do
-8:                  if $\mathbb{M}_{vi}>0$ then
-9:                       $\mathbf{c}_{1}=\mathbf{c}_{1}\setminus\{c_{i}\}$
+3:if $\mathcal{C}_{2}\neq\emptyset$ then
+4:  for $v$ in the indices of local centers in $\mathcal{C}_{3}$ do
+5:    if $\max_{j:c_{j}\in\mathcal{C}_{2}}\mathbb{M}_{v,j}>\max_{i:c_{i}\in\mathcal{C}_{1}}\mathbb{M}_{vi}$ then
+6:     $\mathcal{C}_{2}=\mathcal{C}_{2}\cup\{c_{v}\}$
+7:     for $i$ in the indices of local centers in $\mathcal{C}_{1}$ do
+8:      if $\mathbb{M}_{v,i}>0$ then
+9:         $\mathcal{C}_{1}=\mathcal{C}_{1}\setminus\{c_{i}\}$
 10:      end if
@@
 13:  end for
-14:     $\mathbf{C}=\mathbf{c}_{1}\cup\mathbf{c}_{2}$
-15:     if any isolated local centers exist based on $\mathbf{C}$ then
+14:  $\mathbf{C}=\mathcal{C}_{1}\cup\mathcal{C}_{2}$
+15:  if any isolated local centers exist within the updated set $\mathbf{C}$ then
 16:    Remove isolated local centers from $\mathbf{C}$
@@
 18:else
-19:     $\mathbf{C}=\mathbf{c}_{1}$
+19:  $\mathbf{C}=\mathcal{C}_{1}$
 20:end if
-Algorithm 4 ensures that local centers in $\mathbf{c}_{2}$, which capture the essence of locally non-convex regions, remain disjoint with the local centers in other clusters, i.e., local centers in $\mathbf{c}_{1}$.
+This Algorithm ensures that local centers in $\mathcal{C}_{2}$, which capture the essence of locally non-convex regions, remain disjoint with the local centers in other clusters, i.e., local centers in $\mathcal{C}_{1}$.
 III-B Clustering in DLCC
-Assuming the grouping results for the filtered centers are at hand, with the unique neighbors of each group serving to initialize temporary clusters, let $\mathbf{C}$ represent the set of filtered centers, such that $\mathbf{C}=\bigcup_{k=1}^{G}\mathbf{g}_{k}$, where $G$ denotes the total number of groups, and $\mathbf{g}_{k}$ contains the filtered centers of the $k$th group.
-Consequently, the number of clusters $K$ is inherently defined to be equal to $G$, as each group’s unique neighbors directly initiate a corresponding temporary cluster.
-The subsequent step involves defining a score for each observation with respect to each cluster in order to update these temporary clusters.
-III-B1 Min strategy
-For the min strategy, the score accounts for both clustering and assessing if each observation’s clustering result is acceptable, and it is computed using the similarity matrix $\mathbb{S}$.
-If $\mathbb{S}$ is asymmetric, it is updated by replacing $\mathbb{S}_{ji}$ and $\mathbb{S}_{ij}$ with their average value.
-For clarity, we continue to denote the symmetrized matrix as $\mathbb{S}$, ensuring that $\mathbb{S}_{ij}=\mathbb{S}_{ji}$ in this and the subsequent discussion about the DC metric in Section IV.
-We define the score function for an observation $\mathbf{x}_{i}$ with respect to a temporary cluster $k$ under the min strategy as
-$\operatorname{score}_{i|k}=\frac{\max_{j\in\mathcal{I}_{k}}\mathbb{S}_{ji}-\max_{j\in\bigcup_{z\neq k}\mathrm{I_{z}}}\mathbb{S}_{ji}}{\max\{\max_{j\in\mathcal{I}_{k}}\mathbb{S}_{ji},\max_{j\in\bigcup_{z\neq k}\mathrm{I_{z}}}\mathbb{S}_{ji}\}\,},$ (16)
+The grouping results $\mathbf{G}=\{\mathbf{g}_{1},\ldots,\mathbf{g}_{K}\}$ for the filtered centers are at hand and $\mathbf{C}$ represent the set of filtered centers.
+Consequently, the number of clusters is inherently defined to be equal to $K$, as each group’s unique neighbors directly initiate a corresponding temporary cluster.
+The subsequent step involves defining a score for each observation w.r.t.
+each cluster in order to update these temporary clusters.
+The score accounts for both clustering and assessing if each observation’s clustering result is acceptable, and it is computed using the similarity matrix $\mathbb{S}$.
+Since $\mathbb{S}$ is asymmetric, it is updated by replacing $\mathbb{S}_{j,i}$ and $\mathbb{S}_{i,j}$ with their average value.
+For clarity, we continue to denote the symmetrized matrix as $\mathbb{S}$, ensuring that $\mathbb{S}_{i,j}=\mathbb{S}_{j,i}$.
+We define the score function for an observation $\mathbf{x}_{i}$ w.r.t.
+a temporary cluster $k$ under as
+$\operatorname{score}_{i|k}=\frac{\max_{j\in\mathcal{I}_{k}}\mathbb{S}_{j,i}-\max_{j\in\bigcup_{z\neq k}\mathcal{I_{z}}}\mathbb{S}_{j,i}}{\max\{\max_{j\in\mathcal{I}_{k}}\mathbb{S}_{j,i},\max_{j\in\bigcup_{z\neq k}\mathcal{I_{z}}}\mathbb{S}_{j,i}\}\,}$, (16)
 where $\mathcal{I}_{k}$ is the set of indices for filtered centers within the $k$th group, i.e., $\mathcal{I}_{k}=\{j:\mathbf{x}_{j}\in\mathbf{g}_{k}\}$.
-The score (16) takes a value between $-1$ and $1$ and a positive value indicates that, compared with the nearest filtered center in other groups, the similarity between $\mathbf{x}_{i}$ and the nearest filtered center in group $k$ is larger.
-For each cluster, two score pools are generated: $\mathcal{S}_{k}$, containing both the unique neighbors within the cluster ($\mathbf{U}_{\mathbf{g}_{k}}$) and their corresponding scores, and $\mathcal{\hat{S}}_{k}$, holding observations not in $\mathbf{U}_{\mathbf{g}_{k}}$ yet having positive scores toward this cluster, along with their scores.
-Upon reviewing these pools, observations with high scores may shift from $\mathcal{\hat{S}}_{k}$ to $\mathcal{S}_{k}$, and vice versa, for those with low scores in $\mathcal{S}_{k}$.
-Lastly, $\mathcal{S}_{k}$ includes observations confidently assigned to the cluster, while those in $\mathcal{\hat{S}}_{k}$ remain unlabelled.
-Note that during the reassignment of observations between pools, their corresponding scores are also moved.
-The detailed steps are outlined as Algorithm 5, where the second element of the min function in line $14$ ensures at least $s/2$ observations per temporary cluster, preventing overly small clusters.
-Figure 4 illustrates an example using the Yale B face recognition dataset [56].
-Algorithm 5 removes labels from unique neighbors within each group that are less distinguishable from other points in their own group.
-Meanwhile, it assigns labels to points that are not unique neighbors but are relatively closer to the filtered centers of their group.
-Figure 4: t-SNE visualization [47] of the Yale B dataset.
-Colors represent the true labels of the points, with only those in temporary clusters being colored.
-Points enclosed by diamond-shaped boxes represent filtered centers.
-(a) Initial points in $\mathcal{S}_{k}$, with points enclosed by square boxes indicating those to be moved to $\mathcal{\hat{S}}_{k}$.
-(b) The temporary clusters obtained after applying Algorithm 5.
-Algorithm 5 Update Temporary Cluster
-1:Score pools $\mathcal{S}_{k}$ and $\hat{\mathcal{S}}_{k}$ for $k=1,\ldots,K$, where $\mathcal{S}_{k}^{1}$ and $\hat{\mathcal{S}}_{k}^{1}$ represent observations, and $\mathcal{S}_{k}^{2}$ and $\hat{\mathcal{S}}_{k}^{2}$ represent the corresponding scores, neighborhood size $s$.
-2:Updated score pool $\mathcal{S}_{k}$, where $\mathcal{S}_{k}^{1}$ constructs the $k$-th temporary cluster.
-3:for $k=1$ to $K$ do
-4:     $\gamma_{1}=\text{mean}(\hat{\mathcal{S}}_{k}^{2})$
-5:     for $i$ in $\mathcal{S}_{k}^{2}$ do
-6:         if $i<\gamma_{1}$ then
-7:              $\mathcal{S}_{k}^{2}=\mathcal{S}_{k}^{2}\setminus\{i\}$
-8:              $\hat{\mathcal{S}}_{k}^{2}=\hat{\mathcal{S}}_{k}^{2}\cup\{i\}$
-9:              Move corresponding points from $\mathcal{S}_{k}^{1}$ to $\hat{\mathcal{S}}_{k}^{1}$
-10:         end if
-11:     end for
-12:     $SS=\text{Sort scores in }\mathcal{S}_{k}^{2}\text{ in descending order}$
-13:     $B=\text{length of }SS$
-14:     $idx=\arg\max_{j=\text{floor}(B/2)\text{ to }B-1}(SS[j]-SS[j+1])$
-15:     $\hat{B}=\text{length of }\hat{\mathcal{S}}_{k}^{2}$
-16:     $\gamma_{2}=\min(SS[idx],1-\frac{s/2-B}{\hat{B}}\text{quantile of }\hat{\mathcal{S}}_{k}^{2})$
-17:     for $u$ in $\hat{\mathcal{S}}_{k}^{2}$ do
-18:         if $u>\gamma_{2}$ then
-19:              $\hat{\mathcal{S}}_{k}^{2}=\hat{\mathcal{S}}_{k}^{2}\setminus\{u\}$
-20:              $\mathcal{S}_{k}^{2}=\mathcal{S}_{k}^{2}\cup\{u\}$
-21:              Move corresponding points from $\hat{\mathcal{S}}_{k}^{1}$ to $\mathcal{S}_{k}^{1}$
-22:         end if
-23:     end for
-24:end for
-III-B2 Max strategy
-The score for the max strategy is only for overlapping neighbors between filtered centers in different groups, and is defined as
-$\operatorname{score}_{i|k}=\frac{|\{c\in\mathbf{g}_{k}\mid\mathbf{x}_{i}\in\mathbf{N}_{c}\}|}{|\mathbf{g}_{k}|}.$ (17)
-Observations that are neighbors of filtered centers in different groups are allocated to the cluster with the highest score.
-Unlabeled observations either do not have the highest unique score or are not neighbours of any filtered center.
-III-B3 Classification
-Among three built-in classification methods for the remaining unlabelled observations, MDC is the non-parametric method based on data depth.
-MDC assigns each observation to the cluster within which it achieves its maximum depth value.
-To ensure consistency in our methodology, we employ the same type of data depth in the MDC as we used in defining the local centers.
-When Mahalanobis depth is chosen, the covariance matrix in the classification step is estimated using traditional moment estimation methods.
-This is based on the theoretical consideration that temporary clusters are subsets of clusters and all observations in these temporary clusters are selected with a degree of confidence.
-As for the KNN method, it is an intuitive classification method and is suitable for non-convex shapes.
-Given that we have already constructed the similarity matrix $\mathbb{S}$, the KNN method in our algorithm identifies neighbors through depth-based similarity, rather than relying on Euclidean distance.
-As such, the method identifies neighbors directly within $\mathbb{S}$, thus eliminating the need for additional computational steps.
-The last method, RF, is a renowned machine learning algorithm and has a wide range of applications.
-In classification, it operates without assumptions on the shapes of the clusters and can demonstrate satisfactory performance even in high-dimensional spaces.
-IV Exploring Parameter Selections
-The choice of parameters $s$ (neighborhood size) and $\delta$ (neighborhood similarity threshold) affect the resulting clustering differently.
+The score (16) takes a value between 1-1 and $1$ and a positive value indicates that, compared with the nearest filtered center in other groups, the similarity between $\mathbf{x}_{i}$ and the nearest filtered center in group $k$ is larger.
+For each cluster, two pairs of sets are generated: $\mathbf{T}_{k}$, containing the unique neighbors within the cluster ($\mathbf{U}_{\mathbf{g}_{k}}$), and $\mathcal{S}_{k}$, the corresponding scores for points in $\mathbf{T}_{k}$ w.r.t.
+the $k$th cluster; as well as $\hat{\mathbf{T}}_{k}$, holding unassigned observations with positive scores toward the $k$th cluster, and $\hat{\mathcal{S}}_{k}$, their associated scores.
+During the update process, observations in $\hat{\mathbf{T}}_{k}$ with sufficiently high scores in $\hat{\mathcal{S}}_{k}$ may be moved to $\mathbf{T}_{k}$ and $\mathcal{S}_{k}$, while those in $\mathbf{T}_{k}$ with low scores in $\mathcal{S}_{k}$ may be moved to $\hat{\mathbf{T}}_{k}$ and $\hat{\mathcal{S}}_{k}$.
+After refinement, $\mathbf{T}_{k}$ contains the observations confidently assigned to the $k$th cluster, while $\hat{\mathbf{T}}_{k}$ collects those remaining unassigned.
+We refer to $\{\mathbf{T}_{k}\}_{k=1}^{K}$ as the set of temporary clusters.
+The detailed steps are outlined in Algorithm 5, with only a minor difference between the min and max strategies.
+Under the max strategy, we use a single mean-based cutoff for each cluster, applied uniformly to points in $\mathbf{T}_{k}$ and $\hat{\mathbf{T}}_{k}$ and take $\gamma=0.5$, yielding a looser threshold for admitting points into temporary clusters.
+Under the min strategy, the additional quantile-based cap in Line 9 enforces at least $s/2$ observations per temporary cluster, preventing overly small clusters.
+Figure 4 presents an example based on the Yale B face recognition dataset [44].
+Algorithm 5 updates the temporary clusters by discarding unique neighbors whose scores relative to their currently assigned clusters are insufficient, while simultaneously assigning labels to non–unique neighbors that are sufficiently close to the filtered centers of their respective groups.
+As shown in the figure, this procedure removes visually misclassified points while correctly labeling most points near local centers.
+Figure 4: t-SNE visualization [45] of the Yale B dataset in two-dimensional space (see Section V for dataset description).
+Colors indicate the temporary labels of the points, i.e., membership in $\{\mathbf{T}_{k}\}_{k=1}^{K}$.
+From left to right, the panels illustrate the progression from the initial state (unique neighbors) to the temporary clusters after Algorithm 5.
+Points enclosed by diamond-shaped boxes denote filtered centers.
+Algorithm 5 Updating Temporary Clusters
+1:For $k=1,\ldots,K$, temporary clusters $\mathbf{T}_{k}$, scores $\mathcal{S}_{k}$ for points in $\mathbf{T}_{k}$; Remaining observations for each cluster $\hat{\mathbf{T}}_{k}$ and their scores $\hat{\mathcal{S}}_{k}$; neighborhood size $s$; $\texttt{strategy}\in\{\texttt{min},\texttt{max}\}$.
+2:Updated $\mathbf{T}_{k}$ and $\mathcal{S}_{k}$ for $k=1,\ldots,K$.
+3:Set $\gamma\leftarrow 1$ if $\texttt{strategy}=\texttt{min}$, otherwise $\gamma\leftarrow 0.5$.
+4:for $k=1$ to $K$ do
+5:  $\alpha_{1}\leftarrow\gamma\cdot\mathrm{mean}(\hat{\mathcal{S}}_{k})$ $\triangleright$ if $\hat{\mathcal{S}}_{k}=\emptyset$, use $\mathrm{mean}(\hat{\mathcal{S}}_{\mathrm{all}})$, where $\hat{\mathcal{S}}_{\mathrm{all}}=\bigcup_{\ell:\hat{\mathcal{S}}_{\ell}\neq\emptyset}\hat{\mathcal{S}}_{\ell}$
+6:  for each $i\in\mathbf{T}_{k}$ with score $\alpha_{1}$ do
+7:    $\mathbf{T}_{k}\leftarrow\mathbf{T}_{k}\setminus\{i\}$,   $\hat{\mathbf{T}}_{k}\leftarrow\hat{\mathbf{T}}_{k}\cup\{i\}$; update $\mathcal{S}_{k}$ and $\hat{\mathcal{S}}_{k}$.
+8:  end for
+9:  if strategy = min then
+10:    $SS\leftarrow$ sorted scores in $\mathcal{S}_{k}$ (descending)
+11:    $idx\leftarrow\arg\max_{j=\mathrm{floor}(|\mathcal{S}_{k}|/2)\text{ to }|\mathcal{S}_{k}|-1}(SS[j]-SS[j+1])$
+12:    $\alpha_{2}\leftarrow\min(SS[idx],1-\frac{s/2-|\mathcal{S}_{k}|}{|\hat{\mathcal{S}}_{k}|}$th quantile of $\hat{\mathcal{S}}_{k})$
+[…]
+14:    $\alpha_{2}\leftarrow\alpha_{1}$.
+15:  end if
+16:  for each $u\in\hat{\mathbf{T}}_{k}$ with score $\hat{\alpha}_{2}$ do
+17:    $\hat{\mathbf{T}}_{k}\leftarrow\hat{\mathbf{T}}_{k}\setminus\{u\}$,   $\mathbf{T}_{k}\leftarrow\mathbf{T}_{k}\cup\{u\}$.
+18:  end for
+19:end for
+III-B1 Classification
+After obtaining temporary clusters, we further assign the remaining unlabeled observations via a classification step.
+The motivation is that, for exemplar-based clustering methods, assigning ambiguous points solely according to their similarity to an exemplar can be unreliable, as such points are often weakly associated with any single exemplar.
+Instead, incorporating information from confidently assigned samples provides a more robust basis for assigning these ambiguous points.
+This idea has also been explored in other clustering researches.
+For example, Shi et al.
+[46] noted that, beyond center-based allocation, assignments of determined samples can be used to infer labels of undetermined ones.
+Two built-in options for classifying the remaining observations in DLCC are depth-based kkNN and RF.
+In both cases, the temporary-cluster labels serve as the training labels, and we intentionally avoid additional hyperparameter tuning to keep the procedure lightweight and reproducible.
+The kkNN option is a simple nonparametric classifier that can accommodate non-convex decision boundaries.
+Since DLCC already provides the depth-based similarity matrix $\mathbb{S}$, neighbors are obtained by sorting $\mathbb{S}_{i,\cdot}$, avoiding additional distance evaluations.
+Moreover, the choice of $k$ can be made data-adaptive via the notion of natural neighbors [47].
+Concretely, we take the smallest $k$ such that at least $95\%$ of observations have at least one mutual neighbor within their top-$k$ neighbor lists, where the $95\%$ threshold serves as a safeguard against potential outliers.
+As for RF, it is a widely used nonparametric classifier that makes no explicit shape assumptions and often remains effective in high-dimensional spaces [48].
+Again, we do not perform hyperparameter tuning and fix the number of trees at $100$.
+III-C Complexity Analyses
+The overall computational cost of DLCC is dominated by constructing the depth-based similarity matrix $\mathbb{S}$ and the associated local-center estimation.
+As discussed in Section II, computing $\mathbb{S}$ in its exact form requires $O(n^{3})$ operations.
+Estimating local centers over $n$ neighborhoods of size $s$ incurs an additional cost of $O(ns^{2})$.
+Subsequent filtering and grouping steps operate only on the $T$ local-center candidates, where $T\ll n$, and thus contribute negligibly to the overall complexity.
+The temporary clustering step and the depth-based kkNN classification are performed on top of the precomputed similarity matrix $\mathbb{S}$ and incur at most linear cost in $n$.
+For the RF option, training a forest with $100$ trees has a typical cost roughly $O(100\,n\log n)$.
+Consequently, the overall complexity is dominated by the $O(n^{3})$ cost of constructing $\mathbb{S}$.
+When the data dimension $d$ is large, this complexity is expressed as $O(n^{3}d)$, since the computation of spatial depth involves the evaluation of $L^{2}$ norms of averaged unit vectors.
+IV Exploring Parameter Selection
+IV-A Strategy Chosen
+Under the min strategy, DLCC attempts to find centroids of clusters from filtered centers.
+The variation in the number of filtered centers across different clusters allows for better flexibility than center-based clustering methods.
+The min strategy is more suitable when the sizes of all clusters are comparable.
+Moreover, because it groups centers more conservatively, it is often more reliable for overlapping clusters.
+The max strategy views each cluster as a union of multiple finite convex regions, where each region is approximated by the convex hull of neighborhood points associated with a filtered center.
+Filtered centers are grouped via connectivity.
+Specifically, two filtered centers $c_{y}$ and $c_{z}$ are assigned to the same group if there exists a chain $c_{p_{1}},c_{p_{2}},\ldots,c_{p_{h}}$ with $c_{p_{1}}=c_{y}$, $c_{p_{h}}=c_{z}$, such that the similarity between neighborhoods of any adjacent pair in the chain exceeds $\delta$.
+Consequently, the max strategy is typically more effective for non-convex structures or highly unbalanced cluster sizes.
+When no prior information about the dataset is available, we recommend selecting the strategy via visualization.
+For high-dimensional data, a 2D embedding (e.g., t-SNE [45]) is often informative: if clusters appear well separated and roughly balanced, the min strategy is recommended; otherwise, the max strategy is preferred.
+IV-B Max Strategy
+The choice of the parameters $s$ (neighborhood size) and $\delta$ (neighborhood similarity threshold) affect the resulting clustering differently.
 The neighborhood size $s$ directly affects the defined neighborhoods for each observation and hence impacts the construction of the set $\mathbf{C}$ of filtered centers.
@@
 Basically, a smaller value of $\delta$ results in fewer, and larger, groups of filtered centers.
-IV-A Hierarchical Structure of grouping under the max strategy
-Under the max strategy, given a certain $s$, all possible values of the number of groups can be derived.
+IV-B1 Hierarchical Structure of Grouping
+Given a certain $s$, all possible values of the number of groups can be derived.
 By starting at $\delta=\delta_{0}$ and progressively reducing $\delta$, all $\delta$’s at the time that the number of groups changes can be recorded.
@@
 The order of the matrix denotes the number of groups, and its entries identify the grouping critical points among these groups for a given $\delta$ value.
-In each step, the grouping matrix obtained from the previous one by merging columns and rows that contain the largest non-trivial (not 00 or $1$) value.
+In each step, the grouping matrix is obtained from the previous one by merging columns and rows that contain the largest non-trivial (not 0 or $1$) value.
 The merging process retains the larger of the values in the corresponding entries of the columns and rows.
@@
 The process of grouping generates two key lists: one for the number of groups and another for the corresponding threshold values.
-More precisely, in the toy example, the potential numbers of groups are $6$, $4$, $3$, and $2$, with the $\delta$ values ranging between the intervals $[0.6,\delta_{U})$, $[0.5,0.6)$, $[0.4,0.5)$, and $[0,0.4)$ at each stage.
-The hierarchical structure ensures that if the number of clusters is known or can be estimated using any methods (refer to [48] for estimation techniques), a suitable $\delta$ can be directly determined from the relationship between the lists $\boldsymbol{G}$ and $\boldsymbol{\delta}$.
-IV-B Depth-Based Clustering Metric
-The next challenge of parameter selection is to identify the optimal pair of parameters $(s,\delta)$, which requires an appropriate internal metric to determine the best clustering among various results.
-However, the DLCC under the max strategy often addresses clustering problems involving non-convex shapes or closely positioned clusters of significantly different sizes, where well-known internal metrics such as ASW and CHI encounter difficulties.
-The limitations of these metrics are illustrated using two synthetic datasets, SmileyF and Bainba — see Figures 1 and 6, and Table II.
-In both figures, the clustering results in (a) are the actual labels for the datasets, for which both ASW and CHI yield low scores.
-Both criteria awarded high scores to the clustering result (c) in the SmileyF dataset, although human assessment considers (c) to perform the poorest.
-Similarly, both criteria favour clustering result (b) for the Bainba dataset, which contains only two clusters.
-This discrepancy highlights the need for a metric that better handles clustering of complex shapes.
-Figure 6: Clustering results for the synthetic dataset Bainba.
-Panel (a) shows the ground truth, (b) merges two distinct clusters into one, and (c) incorrectly clusters a point at the top of the central elliptical cluster.
-TABLE II: ASW and CHI values corresponding to the clustering results in Figures 6 and 1.
-DATA SET Label (a) (b) (c)
-Bainba: ASW $0.341$ $0.491$ $0.339$
-Bainba: CHI $642.13$ $842.52$ $641.99$
-SmileyF: ASW $-0.146$ $-0.152$ $0.435$
-SmileyF: CHI $45.25$ $43.18$ $1470.62$
-The DC metric is introduced to evaluate and compare different clustering outcomes when the ground truth is not available, facilitating the identification of the optimal pair of parameters.
-The main idea is to calculate within- and between-cluster similarities in a small, fixed region instead of computing similarities globally to handle non-convex clusters.
-This approach, inspired by the DBCA methodology [33], is adapted in our context with a broader scope.
-While DBCA specifies similarity as $D_{\mathrm{MD}}^{i}(\mathbf{z})$, we generalize this concept to accommodate various forms of similarity matrices, $\mathbb{S}$.
-Some definitions from DBCA are listed below:
-Definition 6 (Core Neighbors).
-Given a depth similarity threshold $\eta$, a point $\mathbf{x}_{j}$ is a core neighbor of $\mathbf{x}_{i}$ (with respect to $\eta$) if $\mathbb{S}_{ij}\geq\eta$.
-Definition 7 (Depth-Connection).
-A point $\mathbf{x}_{u}$ is depth-connected to $\mathbf{x}_{v}$ if there is a chain of points $\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{i}$, where each two adjacent points are core neighbors for each other, $\tilde{\mathbf{x}}_{1}=\mathbf{x}_{u}$, and $\tilde{\mathbf{x}}_{i}=\mathbf{x}_{v}$.
-Definition 8 (DBCA-Cluster).
-A chain $\mathbf{V}=\{\tilde{\mathbf{x}}_{1},\ldots,\tilde{\mathbf{x}}_{i}\}$ is called a maximal depth-connected chain if there is no $\mathbf{x}_{j}\notin\mathbf{V}$ satisfying $\sup_{\mathbf{x}_{i}\in\mathbf{V}}\mathbb{S}_{ij}\geq\eta$.
-Let $\tilde{n}$ be the minimum number of points required to form a cluster in DBCA, then each cluster is formed by observations that are in the same maximal depth-connected chain with length at least $\tilde{n}$.
-Definition 9 (Outliers).
-All observations that remain unlabelled in DBCA are regarded as outliers.
-In Definition 6, larger $\eta$ corresponds to fewer core neighbors, which corresponds to a smaller fixed region.
-Let $\eta_{X}$ be the largest value of $\eta$ such that all observations in the dataset are depth-connected.
-Let $k$ be the label of a cluster in the clustering result which is to be examined by the DC metric, and $\eta_{k}$ be the maximum $\eta$ value such that the cluster is decomposed into a single DBCA-cluster and outliers (without considering observations in other clusters).
-Let $\mathcal{X}_{k}$ be the set of observations in cluster $k$.
-Let $n_{k}=|\mathcal{X}_{k}|$ and $\mathbf{O}_{k}$ be the set of DBCA-outliers.
-For each observation $\mathbf{x}_{i}\in\mathcal{X}_{k}$, we define $\mathcal{J}_{k}^{i}=\{j|\mathbf{x}_{j}\in\mathcal{X}_{k},\mathbf{x}_{j}\neq\mathbf{x}_{i},\mathbb{S}_{ij}>\eta_{k}\}$ as the set of indices.
-Following this, the within cluster similarity $J_{k}$ of cluster $k$ is defined by
-$J_{k}=\sum_{i=1}^{n_{k}}w_{i}\mu_{i},$
-where
-$\mu_{i}$ $=$ $\begin{cases}\dfrac{\sum\limits_{j\in\mathcal{J}_{k}^{i}}\mathbb{S}_{ij}}{|\mathcal{J}_{k}^{i}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{O}_{k},\\ \max\{\mathbb{S}_{ij}\mid\mathbf{x}_{j}\in\mathcal{X}_{k}\setminus\mathbf{O}_{k}\},\text{ if }\mathbf{x}_{i}\in\mathbf{O}_{k},\end{cases}$
-$w_{i}$ $=$ $\begin{cases}\dfrac{|\mathcal{J}_{k}^{i}|}{\sum\limits_{\mathbf{x}_{v}\in\mathcal{X}_{k}\setminus\mathbf{O}_{k}}|\mathcal{J}_{k}^{v}|+|\mathbf{O}_{k}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{O}_{k},\\ \dfrac{1}{\sum\limits_{\mathbf{x}_{v}\in\mathcal{X}_{k}\setminus\mathbf{O}_{k}}|\mathcal{J}_{k}^{v}|+|\mathbf{O}_{k}|},\text{ if }\mathbf{x}_{i}\in\mathbf{O}_{k}.\end{cases}$
-The weights $w_{i}$ in $J_{k}$ give observations with more core neighbors more weight and give outliers less weight.
-By considering DBCA-outliers, the DC metric can be made more robust by avoiding the significant adverse impact of an individual outlier.
-For the cluster $k$, the set of inner observations is
-$\mathbf{I}_{k}:=\left\{\mathbf{x}_{i}\in{\mathcal{X}_{k}}:\mathbb{S}_{ij}<\eta_{X}\text{ for all }\mathbf{x}_{j}\in{\mathbf{X}\setminus\mathcal{X}_{k}}\right\},$
-where $\mathbf{X}$ is the total dataset.
-If $\mathbf{x}_{i}\in\mathbf{I}_{k}$ then $\mathbf{x}_{i}$ is called an inner observation of cluster $k$, i.e., observations in $\mathcal{X}_{k}$ that are not similar to any observation outside of $\mathcal{X}_{k}$.
-For defining the between cluster similarity $H_{k}$ of cluster $k$, we set $\mathcal{H}_{k}^{i}=\{h|\mathbf{x}_{h}\notin\mathcal{X}_{k},\mathbb{S}_{ih}>\eta_{X}\}$ to include indices of observations from other clusters that are core neighbors of points in $\mathcal{X}_{k}$ when $\eta=\eta_{X}$.
-$H_{k}$ is then defined as
-$H_{k}=\sum_{i=1}^{n_{k}}\tilde{w}_{i}\tilde{\mu}_{i},$
-where,
-$\tilde{\mu}_{i}$ $=$ $\begin{cases}\dfrac{\sum\limits_{h\in\mathcal{H}_{k}^{i}}\mathbb{S}_{ih}}{|\mathcal{H}_{k}^{i}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{I}_{k},\\ \eta_{X},\text{ if }\mathbf{x}_{i}\in\mathbf{I}_{k},\end{cases}$
-$\tilde{w}_{i}$ $=$ $\begin{cases}\dfrac{|\mathcal{H}_{k}^{i}|}{\sum\limits_{\mathbf{x}_{u}\in\mathcal{X}_{k}\setminus\mathbf{I}_{k}}|\mathcal{H}_{k}^{u}|+|\mathbf{I}_{k}|},\text{ if }\mathbf{x}_{i}\notin\mathbf{I}_{k},\\ \dfrac{1}{\sum\limits_{\mathbf{x}_{u}\in\mathcal{X}_{k}\setminus\mathbf{I}_{k}}|\mathcal{H}_{k}^{u}|+|\mathbf{I}_{k}|}\,\text{if }\mathbf{x}_{i}\in\mathbf{I}_{k}.\end{cases}$
-In $H_{k}$, the weights $\tilde{w}_{i}$ give observations which are close to other clusters more weight, but less weight for inner observations.
-The total DC metric for a clustering result is then given by
-$\text{DC}=\sum_{k}\frac{n_{k}}{n}(J_{k}-H_{k}),$ (20)
-which gives more weight to the clustering performance in clusters with more observations. Some properties of the DC metric are summarized as follows:
-  1. Property 1:
-    For each observation $\mathbf{x}_{i}\in\mathcal{X}_{k}$, reducing the threshold from $\eta_{k}$ to $\eta_{k^{*}}$ implies $\mu_{i^{*}}\leq\mu_{i}$.
-  2. Property 2:
-    For each observation $\mathbf{x}_{i}\in\mathcal{X}_{k}$, reducing the threshold from $\eta_{k}$ to $\eta_{k^{*}}$ implies $\tilde{\mu}_{i^{*}}\leq\tilde{\mu}_{i}$.
-  3. Property 3:
-    For any cluster $k$, the $\mu_{i}$ given by any outlier is less than $\eta_{k}$.
-  4. Property 4:
-    For any cluster $k$, $H_{k}\geq\eta_{X}$, higher proportions of inner observations in $k$ leads to smaller $H_{k}$.
-Table III displays the DC metrics for the aforementioned clustering results for the Bainba and SmileyF datasets.
-Notably, the ranking of these scores aligns more closely with human assessments.
-Thus, using the DC metric allows for the selection of the parameters for the DLCC algorithm under the max strategy.
-TABLE III: The DC metrics for clustering results in Figures 6 and 1, where $\tilde{n}=3$. Note this score correctly assesses (a) as the best assignment of labels among the three panels.
-(a) (b) (c)
-Bainba $0.046$ $0.044$ $0.045$
-SmileyF $0.174$ $0.164$ $0.079$
+More precisely, in the toy example, the potential numbers of groups are $6$, 44, 33, and $2$, with the $\delta$ values ranging between the intervals $0.6,\delta_{U})$, $[0.5,0.6)$, $[0.4,0.5)$, and $[0,0.4)$ at each stage.
+The hierarchical structure ensures that if the number of clusters is known or can be estimated using any methods, a suitable $\delta$ can be directly determined from the relationship between the lists $\boldsymbol{G}$ and $\boldsymbol{\delta}$.
+IV-B2 Parameter selection
+For the max strategy, clusters are formed by connecting multiple local convex regions.
+Thus, the neighborhood size $s$ should be relatively small.
+In our experiments, we recommend $s\in[20,60]$.
+For each $s$, we generate a corresponding candidate list of thresholds $\boldsymbol{\delta}$ and evaluate the resulting solutions accordingly.
+Note that the construction of the similarity matrix is independent of the parameters.
+Therefore, it can be computed once in advance before the parameter selection stage.
+Without access to ground truth, clustering is inherently exploratory.
+As noted by Hennig [49], there is in general no unique “true” clustering, and validation often rely on a combination of internal validation indices (IVI) and subjective assessment, where visualization plays an important supporting role.
+Our parameter selection rule follow this philosophy.
+In terms of IVIs, classical criteria such as the silhouette width [50] and the Calinski-Harabasz index [51] are primarily designed for compact and approximately convex clusters, and are not suitable for clusters with arbitrary shapes.
+To better reflect arbitrary geometries, we adopt the cluster validity index based on density-involved distance (CVDD) [52].
+[^1] https://github.com/hulianyu/CVDD CVDD replaces the usual notion of separation by a density-aware connectivity distance defined on a kkNN graph.
+Two points are regarded as close only if they can be connected by a path that stays in relatively high-density regions.
+Consequently, CVDD penalizes clustering results that effectively require “crossing” low-density gaps to maintain within-cluster connectivity, while remaining insensitive to non-convex shapes as long as they are density-connected.
+The resulting score balances density-aware between-cluster separation against within-cluster connectivity; further details are referred to Hu and Zhong [52].
+However, IVIs inevitably encode particular assumptions about “separation” and “cohesion”, and tuning based on a single IVI alone is therefore often insufficient in practice.
+Indeed, Lewis et al.
+[53] reported that human assessments can be more consistent than existing IVIs when judging clusterings.
+We therefore combine CVDD with visualization for model selection.
+Specifically, we run DLCC over the grid of $(s,\boldsymbol{\delta})$, which yields solutions with varying estimated numbers of clusters $\hat{K}$.
+For each distinct $\hat{K}$, we retain the solution that attains the largest CVDD value as its representative.
+We then visualize the representatives of the top three $\hat{K}$ values ranked by CVDD and choose the final model by visual assessments.
+In experiments shown in Figure 7, although CVDD identifies suitable parameter pairs for most datasets, it may not always favor the most appropriate $\hat{K}$.
+For instance, on the Bainba dataset (LABEL:subfig:bainba), CVDD actually prefers merging the two upper groups due to their proximity and similar local densities; visually, however, they form two distinct clusters and should remain separated.
+IV-C Min strategy
+Figure 6: Empirical association between the proposed depth-connectivity score (blue; left axis) and the ARI (orange; right axis) as the neighborhood size $s$ varies under the min strategy with a fixed $K$.
+Under the min strategy, particularly for real-world high-dimensional data (e.g., $d>50$), selecting parameters solely using a single IVI is often challenging.
+Nevertheless, IVIs (e.g., CVDD) remain useful for narrowing down reasonable ranges of $s$ and estimate $\hat{K}$.
+With a known or an estimated $K$, we recommend choose $s\in[0.6\,n/K,\,n/K]$.
+To further evaluate candidate clusterings, we then exploit the symmetrized depth-based similarity matrix $\mathbb{S}$.
+Before proceeding, we introduce the reachable similarity, a max-min (bottleneck-path) connectivity measure in graph theory [54, 55].
+Definition 6 (Reachable similarity).
+Let $S$ be a symmetric similarity matrix with entries in $[0,1]$. The reachable similarity between nodes $i$ and $j$ is
+$\mathrm{rs}(i,j)=\max_{\boldsymbol{p}:i\to j}\ \min_{(u,v)\in\boldsymbol{p}}S_{u,v},$
+i.e., the strongest bottleneck similarity over all paths connecting $i$ and $j$.
+Given a candidate partition $\{\pi_{g}\}_{g=1}^{K}$, we quantify connectivity using two pointwise quantities: the intra-cluster reachable similarity (IR) and the between-cluster reachable similarity (BR).
+For each point with label $\widehat{y}_{i}$, its IR is defined as the average reachable similarity from $\mathbf{x}_{i}$ to other points within the same cluster, where $\mathrm{rs}(\cdot,\cdot)$ is computed on the subgraph induced by nodes in $\pi_{\widehat{y}_{i}}$.
+To define BR, we first identify the nearest competing cluster for $\mathbf{x}_{i}$ by comparing $\mathbf{x}_{i}$ to filtered centers from other clusters
+$\tilde{g}_{i}=\arg\max_{g\neq\widehat{y}_{i}}\ \max_{\mathbf{x}_{u}\in\mathbf{g}_{g}}\ \mathbb{S}_{i,u},$
+where $\mathbf{g}_{g}$ is the $g$th group of filtered centers.
+We then set BRi to be the average reachable similarity from $\mathbf{x}_{i}$ to points in the competing cluster $\pi_{\tilde{g}_{i}}$, with $\mathrm{rs}(\cdot,\cdot)$ computed on the subgraph induced by nodes in $\pi_{\tilde{g}_{i}}$.
+Finally, we summarize a clustering result by the mean gap
+$\texttt{Connectivity}=\frac{1}{n}\sum_{i=1}^{n}\big(\texttt{IR}_{i}-\texttt{BR}_{i}\big),$ (19)
+where larger values indicate stronger within-cluster connectivity relative to the most competitive cross-cluster connectivity.
+We observe that when $K$ is fixed and ground truth labels are available, this depth-connectivity score exhibits a strong positive association with external agreement measures such as the adjusted Rand index (ARI; [56]) across a range of datasets; representative examples are shown in Fig.
+6.
+This observation motivates using (19) as a practical tool for ranking candidate neighborhood sizes $s$ under the min strategy when ground truth labels are unavailable.
 V Experiments and Evaluations
-In this section, we evaluate the performance of DLCC under various strategies, comparing it with several other clustering algorithms on both synthetic and real-world datasets.
-External Metric.
-To compare the performance of different clustering algorithms on the same datasets, we use the clustering error (CE) [28, p.
-2774] and adjusted Rand index (ARI) [51].
+In this section, we compare DLCC under both strategies with several competing clustering methods on synthetic and real-world data sets.
+In DLCC, the final-stage classifier is selected via CVDD; when RF is chosen, we report the median ARI, NMI, and CE over $100$ independent runs to account for its stochasticity.
+Detailed records of the selected classifiers are provided in the supplementary material.
+V-A External Metric
+To compare the performance of different clustering algorithms on the same datasets, we use the clustering error (CE; [57, p.
+2774]), ARI and normalized mutual information (NMI; [58]).
 The CE simply computes the proportion of mismatched points.
-This makes CE intuitive, with a value range of 00 (perfect match) to $1$.
+This makes CE intuitive, with a value range of 0 (perfect match) to $1$.
 The ARI, on the other hand, measures pairwise similarity between clustering results and ground truth.
 The expected value fo the ARI under random class assignment is 0 and it takes a value $1$ under perfect agreement.
-Methods for comparison.
-We compare our approach with several representative clustering methods.
-These include the center-based algorithm K-means and its state-of-the-art variant LW-K-means; the recent advanced DPC method R-MDPC; the subspace clustering algorithm SSC; the model-based clustering algorithm GMM (GPCM family);
-V-A Description of Datasets
-Figure 7 presents four synthetic datasets.
-The first subfigure is simulated using custom code, while the second and third subfigures are generated with the mlbench [53] package in R [52].
-The last subfigure provides a two-dimensional visualization of the ray shapes data, which is derived from a study on spectral clustering [54].
-These synthetic datasets represent different clustering scenarios with non-convex shapes, clusters of varying sizes and densities in close proximity, and large overlapping regions.
-Each of these characteristics presents unique challenges for clustering.
-Figure 7: Graphs of synthetic datasets: (a) Starbeam; (b) Blend; (c) Bainba; (d) Ray [54]. In $(d)$, $x$ and $y$ axes are the first two variables of ray data. Colors represent true labels.
-As for the real-world datasets, we utilize five benchmark datasets from the UCI Machine Learning Repository [55]: iris, seed, wine, optidigits, pendigits.
-Additionally, we test high-dimensional datasets ($d\geq 100$) with $n>d$, including Yale B [56] and HAR [57], which represent human-face and human-activity datasets, respectively, addressing clustering problems in different domains.
-Among these, Yale B, optidigits, and pendigit are sourced from the R package PPCI [58].
-Furthermore, we consider two microarray datasets: leukemia [59] and SuCancer [60], which have large dimensions $d$ but small sample sizes $n$.
-V-B Result comparisons
-Table IV provides the settings, used for running DLCC on the above-mentioned datasets.
-The max strategy tends to perform better with non-convex or unbalanced sized clusters.
-Conversely, the min strategy is more effective for overlapping clusters of similar sizes.
-TABLE IV: DLCC settings for different datasets, where CM stands for classification method.
-Datasets Depth Strategy CM maxdepth ifloop
-Starbeam MD Max KNN False --
-Starbeam SD Max KNN False --
-Blend MD Max KNN False --
-Blend SD Max KNN False --
-Bainba MD Max MDC True --
-Bainba SD Max KNN False --
-Ray MD Min MDC False True
-Ray SD Min RF False False
-Iris MD Min MDC True False
-Iris SD Min RF False False
-Seed MD Min KNN True True
-Seed SD Min KNN False False
-Wine MD Min MDC True True
-Wine SD Min MDC False False
-Yale B SD Min RF False False
-Optidigit SD Min KNN False False
-Pendigits MD Min KNN False False
-Pendigits SD Min KNN False False
-HAR SD Min KNN False False
-Leukemia SD Max MDC False --
-SuCancer SD Max RF False --
-TABLE V: Comparative analysis of selected clustering algorithms on several datasets ( “–” means a method is not applicable).
-Datasets Metric DLCCMD DLCCSD Kmeans LW-Kmeans SSC GMM R-MDPC
-Starbeam ARI $\mathbf{1}$ $\mathbf{1}$ $0.4822$ $0.4822$ $0.3052$ $0.0157$ $0.8123$
-$n=780,d=2,K=2$ CE $\mathbf{0}$ $\mathbf{0}$ $15.26\%$ $15.26\%$ $22.31\%$ $40.26\%$ $4.87$%
-Blend ARI 0.85950.8595{0.8595}0.8595 $\mathbf{0.8605}$ $0.7266$ $0.7365$ $0.5716$ $0.7518$ $0.8453$
-$n=2000,d=2,K=10$ CE ${12.50}\%$ $\mathbf{8.65}\%$ $25.70\%$ $18.90\%$ $31.35\%$ $21.45\%$ ${9.10}\%$
-Bainba ARI $\mathbf{1}$ $\mathbf{1}$ $0.4961$ $0.9814$ $0.4276$ $0.7436$ $\mathbf{1}$
-$n=600,d=2,K=3$ CE $\mathbf{0}$ $\mathbf{0}$ $32.00\%$ $0.67\%$ $30.17\%$ $10.50\%$ $\mathbf{0}$
-Ray ARI $\mathbf{0.8261}$ $0.7521$ $0.6388$ $0.7090$ $0.7081$ $0.8245$ $0.5364$
-$n=1000,d=5,K=4$ CE $\mathbf{7.00}\%$ $10.00\%$ $15.00\%$ $11.90\%$ $12.60\%$ $7.10\%$ $37.70\%$
-Iris ARI $\mathbf{0.9039}$ $0.8857$ $0.7302$ $0.8857$ $0.6138$ $\mathbf{0.9039}$ $0.8857$
-$n=150,d=4,K=3$ CE $\mathbf{3.33}\%$ $4.00\%$ $10.67\%$ $4.00\%$ $17.33\%$ $\mathbf{3.33}\%$ $4.00\%$
-Seed ARI $0.7612$ $\mathbf{0.7626}$ $0.7167$ $0.6687$ $0.4462$ $0.6299$ $0.7289$
-$n=210,d=7,K=3$ CE $\mathbf{8.57}\%$ $\mathbf{8.57}\%$ $10.48\%$ $12.86\%$ $33.33\%$ $14.76\%$ $10.00\%$
-Wine ARI $\mathbf{0.9817}$ $0.9295$ $0.8975$ $0.8516$ $0.8804$ $0.9309$ $0.7269$
-$n=178,d=13,K=3$ CE $\mathbf{0.56}\%$ $2.25\%$ $3.37\%$ $5.06\%$ $3.93\%$ $2.25\%$ $9.66\%$
-Yale B ARI -- $\mathbf{0.9901}$ $0.5663$ $0.8461$ $0.9484$- $0.7635$
-$n=2000,d=600,K=10$ CE -- $\mathbf{0.45}\%$ $31.85\%$ $13.25\%$ $2.40\%$- $12.10\%$
-Optidigits ARI -- $0.8144$ $0.5682$ $0.5799$ $0.7540$ $0.4878$ $\mathbf{0.8543}$
-$n=5620,d=64,K=10$ CE -- $9.40\%$ $28.81\%$ $28.26\%$ $14.84\%$ $36.89\%$ $\mathbf{8.75}\%$
-Pendigits ARI $0.6488$ $0.6661$ $0.5526$ $0.5756$ $\mathbf{0.7069}$ $0.6522$ $0.6657$
-$n=10992,d=16,K=10$ CE $20.52\%$ $18.22\%$ $32.25\%$ $27.83\%$ $\mathbf{15.56}\%$ $20.58\%$ $19.17\%$
-HAR ARI -- $0.6564$ $0.4610$ $0.4556$ $\mathbf{0.7005}$- $0.4373$
-$n=10299,d=561,K=6$ CE -- $\mathbf{18.68}\%$ $40.00\%$ $41.51\%$ $23.11\%$- $51.06\%$
-Leukemia ARI -- $\mathbf{0.8897}$ $\mathbf{0.8897}$ $\mathbf{0.8897}$ $0.7371$- $0.7851$
-$n=72,d=3571,K=2$ CE -- $\mathbf{2.78}\%$ $\mathbf{2.78}\%$ $\mathbf{2.78}\%$ $6.94\%$- $5.56\%$
-SuCancer ARI -- $0.1154$ $-0.003$ $-0.003$ $\mathbf{0.1234}$- $0.1154$
-$n=174,d=7909,K=2$ CE -- $32.76\%$ $47.70\%$ $47.70\%$ $\mathbf{32.18}\%$- $32.76\%$
-Table V illustrates the performance of DLCC with both MD and SD, compared to other clustering algorithms.
-For all comparing methods, we pick the result achieving the highest ARI as the best.
-For K-means and LW-K-means, we report the best results from $10$ runs.
-For the R-MDPC method, following the suggestions, we fix $cv=1$ and test all neighborhood sizes from $0.5\sqrt{n}$ to $2\sqrt{n}$ [18].
-For SSC, the best combinations are chosen with regards to the hyper parameter $\alpha$ (ranging from $10$ to $50$ with a step size of $10$) and the logical parameters “affine” (indicating if affine subspaces are considered) and “outlier” (signifying if the data might be contaminated by outliers).
-In the case of GMM, the R package mixture [61] is used to identify the optimal model from the GPCM family.
-As for DLCC, when RF is employed as the classification method, its inherent randomness led us to conduct $100$ separate runs with unique seeds.
-The corresponding internal metric for each run was calculated, and the median performance score was recorded in the table.
-Table V clearly demonstrates that DLCC yields impressively competitive results across all datasets.
-Specifically, in the synthetic datasets, both versions of DLCC perform perfectly in Starbeam and Bainba, while they meet some challenges in Blend.
-As detailed in Section III-A, the max strategy assumes that, within any given cluster, each subset used for defining local centers is convex.
-Although the filtering procedures for local centers can identify the areas of non-convex subsets, they are unable to further distinguish individual clusters within these regions and, in the current context, to identify the two distinct spirals.
-DLCCMD achieves the best performance on the Ray dataset, closely matching that of GMM, as its similarity matrix is constructed based on the EEV model, a member of GMM.
-DLCCSD proves highly adaptable for high-dimensional data applications, delivering the best results in the Yale B and Leukemia datasets and ranking among the top performers on other high-dimensional datasets.
-In contrast, each of the other algorithms displays distinct strengths and limitations.
-For instance, K-means and LW-Kmeans produce satisfactory results for convex datasets, while they do not perform well on datasets with non-convex shapes.
-R-MDPC performs well on non-convex datasets and excels on the Optidigits dataset.
-However, it faces challenges with datasets that have large overlapping regions with similar densities, such as Ray and HAR.
-SSC excels with high-dimensional datasets but falters when the dimension $d$ is less than or close to the number of clusters $K$.
-Both GMM and DLCCMD face challenges in high-dimensional data due to their reliance on covariance matrix estimations.
-Accurately estimating covariance matrices requires significantly more observations per cluster than the number of dimensions, which is often not feasible for high-dimensional data.
-V-C Performance in estimating the number of cluster
-For the max strategy, the relationship between $\delta$ and the number of clusters is detailed in Section IV.
-Here, we evaluate the effectiveness of Algorithm 1 in estimating the number of clusters under the min strategy.
-To ensure a fair comparison, the results presented in Table V assume that the true number of clusters is known.
-In this section, we focus on the scenario where $K$ is not pre-specified, specifically discussing DLCCSD as it is applicable to all datasets.
-To provide insights into the performance of our method, we compare it with a well-established spectral clustering method, Spectrum [62], which can automatically estimate the number of clusters during the clustering process without requiring extensive parameter tuning.
-The implementation of Spectrum is based on the R package Spectrum [63], and we report the better results between the eigengap method (largest gap between eigenvalues) and the multimodality gap method.
-Table VI demonstrates that without pre-defined $K$, DLCCSD consistently outperforms Spectrum in terms of ARI.
-Although DLCCSD estimates two additional clusters in the Pendigits dataset, its ARI is even higher compared to the scenario where $K$ is fixed at $10$.
-For the Optidigits and HAR datasets, DLCCSD underestimates the number of clusters by one.
-Specifically, it merges clusters “3” and “8” into a single cluster in Optidigits, and in HAR, it groups walking, walking upstairs, and walking downstairs into one cluster.
-This clustering also makes sense based on human assessments.
-TABLE VI: Comparison of performance between DLCCSD and Spectrum across datasets. Here, $\hat{K}$ denotes the estimated number of clusters, with differences from true $K$ in parentheses.
-Method Ray Iris Seed Wine Yale B Optidigits Pendigits HAR
-$\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI $\hat{K}$ ARI
-DLCCSD $4$ (00) $0.7521$ $3$ (00) $0.8857$ $3$ (00) $0.7626$ $3$ (00) $0.9295$ $10$ (00) $0.9901$ $9$ ($-1$) $0.7174$ $12$ ($+2$) $0.6885$ $4$ ($-2$) $0.4813$
-Spectrum $3$ ($-1$) $0.6468$ $3$ (00) $0.8508$ $3$ (00) $0.5702$ $3$ (00) $0.9170$ $9$ ($-1$) $0.7575$ $9$ ($-1$) $0.6527$ $10$ (00) $0.5086$ $10$ ($+4$) $0.3242$
-V-D Complexity Analysis
-The computational cost of the DLCC algorithm is dominated by constructing the depth-based similarity matrix and estimating local centers.
-For both versions, the first term corresponds to constructing the similarity matrix, and the second term accounts for estimating local centers, with $s\ll n$.
-  • DLCCMD: With pre-estimated covariance matrices (Section II-B), the complexity is $O(n^{2})+O(ns^{2}+ns)$, where $O(ns^{2}+ns)$ covers covariance matrix estimation and depth computation for neighborhoods.
-As $s\ll n$, the overall complexity simplifies to $O(n^{2})$.
-  • DLCCSD: The complexity is $O(n^{3})+O(ns^{2})$.
-Constructing the similarity matrix requires $O(n^{3})$, as computing SD values for $n$ points ($O(n^{2})$) across $n$ reflected datasets introduces an additional $n$ factor.
-Estimating local centers for neighborhoods adds $O(ns^{2})$, resulting in an overall complexity of $O(n^{3})$.
-Figure 8: Running time of DLCC on simulated datasets without acceleration techniques, where blue represents DLCCSD and red represents DLCCMD: (a) Fixed the dimension at $30$ while increasing the sample size up to $5060$, y-axis up to $700$s; (b) Fixed the sample size at $1000$ while increasing the dimension up to $1000$, y-axis up to $25$s.
-Figure 8(a) shows that the experimental running times align with theoretical complexities: $O(n^{2})$ for DLCCMD and $O(n^{3})$ for DLCCSD.
-In Figure 8(b), the running time of DLCCMD appears irregular as dimensionality increases because when estimating the covariance matrix, the method defaults to the identity matrix if singularity issues arise, which are more likely when the dimensionality is not significantly smaller than the sample size.
-In contrast, DLCCSD shows a linear trend due to the SD computation involving the $L^{2}$ norm of the average of unit vectors, with a complexity of $O(n^{3}d)$ when $d$ is large.
-While DLCC can handle datasets with approximately $10,000$ samples, its computational cost is dominated by the depth-based similarity matrix, limiting its scalability to larger datasets.
-However, DLCCSD is computationally efficient with increasing dimensionality.
-Future work could explore approximation methods to accelerate the algorithm.
-For instance, defining local centers does not necessarily require computing the entire similarity matrix but rather identifying suitable neighborhoods for the centers.
-Fast approximate neighborhood search methods (e.g., [64]) could be employed to estimate candidate local centers (with size much smaller than $n$), followed by exact depth similarity computations for points relative to these candidates.
-VI Conclusions and Future work
+NMI measures the shared information between the predicted and true partitions and is normalized to $[0,1]$, with $1$ indicating perfect agreement and larger values implying better alignment.
+V-B Methods for Comparison
+We compare DLCC with representative clustering methods that rely on exemplars or explicitly group local sub-regions/exemplars which show resemblance of DLCC, including LW-k-means [12], EKM [14], R-MDPC [18], granular-ball clustering (GBCT; [20]), and ACLR [27].
+We also include a recent depth-based alternative to mean-shift [35], denoted as DMS.
+For all those methods except DMS, we use the authors’ released implementations; DMS is implemented by us using lens depth [59] since no existing implementation can be found.
+All benchmark datasets considered in this paper are equipped with ground-truth labels.
+Since a single solution is generally not expected to simultaneously optimize all three external metrics (ARI, NMI, and CE), we adopt ARI as the primary selection metric throughout.
+Whenever a method involves hyperparameter tuning, we select and report the solution with the best ARI, and report the corresponding NMI and CE for the same solution.
+For fair comparisons, we assume the true number of clusters $K$ is known, and set $K$ accordingly for methods that require it.
+For LW-k-means, we report the median performance over $100$ random initializations.
+For EKM, we set the number of replications to $1000$ and report the solution with the smallest objective value (as in the original implementation).
+For R-MDPC, we follow the authors’ recommendation by fixing $cv=1$ and scanning neighborhood sizes in $0.5\sqrt{n},\,2\sqrt{n}]$.
+For ACLR, we scan the hierarchy of binary trees $h\in\{7,\ldots,11\}$ and the anchor-graph neighborhood size in $[10,100]$.
+For DMS, we scan the quantile level used to determine the localization parameter $\tau$ over $[0.01,0.5]$.
+The direction parameter is tuned as follows: we first evaluate two initial values ($20$ and $500$) and choose the better one as the starting point.
+We then perform a one-sided search by increasing from $20$ (or decreasing from $500$) and stop when the ARI drops by more than $0.05$ for two consecutive steps.
+V-C Synthetic Datasets
+We evaluate DLCC on $12$ synthetic datasets designed to cover a broad range of clustering scenarios and to give insights on both the strengths and limitations of the proposed DLCC algorithm.
+Figure 7 visualizes the DLCC partitions.
+The datasets 3 blobs, Aggregation, Atom, Chainlink, Cuboids, Diamond, Flame, and Jain are benchmark examples from [60].
+Bainba and Blend are generated using the mlbench package [61] in R [62].
+Ray is adapted from a spectral-clustering study [63], and Starbeam is constructed by us.
+Following the order in Figure 7, the synthetic datasets represent the following clustering challenges:
+  1. (a)
+    well-separated, convex shapes.
+  2. (b)
+    bridge-connected shapes with varying sizes.
+  3. (c)
+    a non-convex component enclosing a convex core.
+  4. (d)
+    proximity-driven ambiguity under strong size imbalance.
+  5. (e)
+    arbitrary shapes with substantial heterogeneity.
+  6. (f)
+    non-convex interlocked rings.
+  7. (g)
+    well-separated shapes with different orientations.
+  8. (h)
+    close convex shapes.
+  9. (i)
+    close non-convex shapes with similar density.
+  10. (j)
+    two curved non-convex shapes separated by a narrow gap, with varying densities.
+  11. (k)
+    overlapping/entangled shapes.
+  12. (l)
+    non-convex connected shapes with varying densities.
+In these experiments, we adopt the min strategy for 3 blobs, Diamond, and Ray, and the max strategy for the remaining datasets (refer to Section IV-A for reasons).
+Figure 7: DLCC clustering visualizations on 12 synthetic datasets. For the Ray dataset ($d=5$), the plot is produced by projecting the data onto the first two principal components.
+TABLE II: Comparative analysis of selected clustering algorithms on synthetic datasets (For methods without a fixed $K$, when the estimated $\hat{K}$ differs from the true $K$, the CE is not meaningful and we report $\hat{K}$ instead).
+Datasets Metric DLCC LW-k-means EKM R-MDPC GBCT ACLR DMS
+3 blobs $n=300,d=2,K=3$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.9900$ $0.9830$ $0.33\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.9900$ $0.9830$ $0.33\%$ $0.9900$ $0.9830$ $0.33\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$
+Aggregation $n=788,d=2,K=7$ ARI NMI CE $0.9927$ $0.9883$ $0.38\%$ $0.7434$ $0.8409$ $13.07\%$ $0.6325$ $0.7646$ $29.31\%$ $0.8415$ $0.9091$ $14.09\%$ $0.7359$ $0.8378$ $21.57\%$ $\boldsymbol{0.9956}$ $\boldsymbol{0.9923}$ $\boldsymbol{0.25\%}$ $0.9109$ $0.9503$ $\hat{K}=6$
+Atom $n=800,d=3,K=2$ ARI NMI CE $0.0193$ $0.1220$ $43.00\%$ $0.1929$ $0.3044$ $28.00\%$ $0.1066$ $0.2332$ $33.63\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.7587$ $0.8201$ $20.25\%$
+Bainba $n=600,d=2,K=3$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.9814$ $0.9686$ $0.67\%$ $0.4600$ $0.5016$ $38.50\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.7578$ $0.7764$ $9.67\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.9600$ $0.9362$ $\hat{K}=4$
+Blend $n=2000,d=2,K=10$ ARI NMI CE $0.8579$ $0.9588$ $\hat{K}=9$ $0.7310$ $0.8432$ $19.58\%$ $0.5793$ $0.7363$ $32.50\%$ $0.8337$ $0.9448$ $16.35\%$ $0.9656$ $0.9795$ $45\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.8218$ $0.8985$ $\hat{K}=14$
+Chainlink $n=1000,d=3,K=2$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ 0.0010-0.0010 $0.0000$ $49.80\%$ 0.0010-0.0010 $0.0000$ $49.70\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$
+Cuboids $n=1002,d=3,K=4$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}a$ $0.6713$ $0.7390$ $28.14\%$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$
+Diamond $n=3000,d=2,K=9$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.7855$ $0.8600$ $16.33\%$ $0.9992$ $0.9990$ $0.03\%$ $0.8758$ $0.9473$ $0.27\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$
+Flame $n=240,d=2,K=2$ ARI NMI CE $0.9666$ $0.9269$ $0.83\%$ $0.4880$ $0.4422$ $15.00\%$ $0.3988$ $0.3544$ $8.33\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.9501$ $0.8991$ $1.25\%$ $0.9666$ $0.9269$ $0.83\%$ $0.9666$ $0.9269$ $0.83\%$
+Jain $n=373,d=2,K=2$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.6352$ $0.5908$ $9.92\%$ $0.5681$ $0.5088$ $2.06\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.2563$ $0.2463$ $9.03\% 𝟏\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.7146$ $0.6315$ $\hat{K}=3$
+Ray $n=1000,d=5,K=4$ ARI NMI CE $\boldsymbol{0.7577}$ $\boldsymbol{0.7541}$ $\boldsymbol{9.65\%}$ $0.7090$ $0.7439$ $11.90\%$ $0.4189$ $0.4847$ $31.40\%$ $0.5364$ $0.6175$ $37.70\%$ $0.6117$ $0.6595$ $21.10\%$ $0.6794$ $0.7013$ $13.10\%$ $0.2814$ $0.4245$ $\hat{K}=14$
+Starbeam $n=780,d=2,K=2$ ARI NMI CE $\boldsymbol{1}$ $\boldsymbol{1}$ $\boldsymbol{0.00\%}$ $0.4822$ $0.5072$ $15.26\%$ $0.4680$ $0.4982$ $15.77\%$ $0.5957$ $0.4671$ $11.28\%$ 0.0248-0.0248 $0.0379$ $39.23\%$ $0.9948$ $0.9859$ $0.13\%$ $0.2823$ $0.4351$ $\hat{K}=4$
+Table II summarizes the clustering performance on the $12$ synthetic benchmarks.
+Overall, DLCC delivers consistently strong results and remains competitive across a broad range of challenges.
+In particular, DLCC attains perfect recovery on $7$ datasets, next to ACLR (an advanced graph-based method), and achieves favorable ARI/NMI with low CE on most of the remaining examples, illustrating its ability to handle both convex and non-convex structures as well as varying densities and close proximity.
+Nevertheless, DLCC faces difficulties on Atom and on the spiral structures in Blend.
+As discussed in Section III-A, the local neighborhoods used to define local centers are expected to be approximately convex.
+This assumption may fail for manifold-like structures, where the neighborhood geometry around most points is non-convex.
+A representative failure case is Atom, in which the depth-defined local centers fail to adequately capture this geometry.
+A similar issue arises for the spiral components in Blend.
+This limitation of depth-defined local centers explains the observed performance degradation and highlights an inherent weakness of the current DLCC model.
+V-D Real-world Datasets
+As for the real-world benchmarks, we evaluate DLCC on $13$ datasets.
+Specifically, Iris, Seed, Wine, Parkinsons (Pa), Breast Cancer (BC), Segmentation (Seg), Optidigits, and Pendigits are obtained from the UCI Machine Learning Repository [64].
+A highly imbalanced dataset, Anuran Calls (Anuran) is also included, whose cluster sizes ranging from as few as $68$ to as many as $3478$ [65].
+We further include high-dimensional datasets ($d\geq 100$) with $n>d$, namely Yale B [44] and HAR [66], representing face images and human-activity signals, respectively.
+Among these, Yale B, Optidigits, and Pendigits are accessed via the R package PPCI [67].
+In addition, we consider two microarray datasets, Leukemia [68] and SuCancer [69], which are characterized by extremely high dimensionality and small sample sizes.
+Overall, this benchmark suite is representative, spanning tabular, image, and time-series datasets, as well as a wide range of dimensionalities (low-dimensional to microarray-scale) and sample sizes.
+For these real-world datasets, we adopt the max strategy for Pa, Anuran, Leukemia, and SuCancer due to severe class imbalance, and use the min strategy for all remaining datasets.
+Table III shows that DLCC is consistently competitive across the $13$ real-world benchmarks.
+It achieves the best performance on $6$ datasets and ranks within the top three on all datasets.
+DLCC also proves highly adaptable for high-dimensional data applications, attaining the strongest results on Yale B and HAR; on SuCancer it is still among the best, although the absolute performance is modest for all methods.
+The table further indicates that competing methods are often dataset-dependent: some methods can be outstanding on particular datasets (e.g., EKM on Wine, R-MDPC on Optidigits, and ACLR on several benchmarks), yet no single method dominates uniformly.
+This is largely because different algorithms have different notions of what constitutes a “cluster” (e.g., centroid-based compactness, graph connectivity, or density/shape regularity), which may align well with some datasets but not with others.
+This variability underscores the practical value of DLCC as a robust alternative with strong performance across different real-world scenarios.
+TABLE III: Comparative analysis of selected clustering algorithms on real datasets.
+Datasets Metric DLCC LW-k-means EKM R-MDPC GBCT ACLR DMS
+Iris $n=150,d=4,K=3$ ARI NMI CE $\boldsymbol{0.8857}$ $\boldsymbol{0.8642}$ $\boldsymbol{4.00\%}$ $\boldsymbol{0.8857}$ $\boldsymbol{0.8642}$ $\boldsymbol{4.00\%}$ $0.6101$ $0.6526$ $17.33\%$ $\boldsymbol{0.8857}$ $\boldsymbol{0.8642}$ $\boldsymbol{4.00\%}$ $0.5638$ $0.7038$ $26.00\%$ $0.7592$ $0.7960$ $9.33\%$ $0.6956$ $0.7711$ $12.67\%$
+Wine $n=178,d=13,K=3$ ARI NMI CE $0.8975$ $0.8759$ $3.37\%$ $0.8516$ $0.8267$ $5.06\%$ $\boldsymbol{0.9134}$ $\boldsymbol{0.8920}$ $\boldsymbol{2.81\%}$ $0.7269$ $0.7435$ $9.55\%$ $0.7033$ $0.7397$ $10.67\%$ $0.4041$ $0.4127$ $26.97\%$ $0.2748$ $0.3933$ $36.52\%$
+Seed $n=210,d=7,K=3$ ARI NMI CE $0.7868$ $0.7409$ $\boldsymbol{7.62}\%$ $0.6687$ $0.6728$ $12.86\%$ $0.7715$ $0.7315$ $8.10\%$ $0.7289$ $0.6946$ $10.00\%$ $0.1614$ $0.3061$ $49.52\%$ $0.7132$ $0.6841$ $10.48\%$ $\boldsymbol{0.7870}$ $\boldsymbol{0.7423}$ $\boldsymbol{7.62}\%$
+Pa $n=195,d=22,K=2$ ARI NMI CE $\boldsymbol{0.4226}$ $\boldsymbol{0.3265}$ $\boldsymbol{14.36}\%$ 0.0937-0.0937 $0.1189$ $44.10\%$ 0.0710-0.0710 $0.0558$ $31.79\%$ $0.1256$ $0.2776$ $31.79\%$ 0.0515-0.0515 $0.0360$ $29.23\%$ $0.2890$ $0.1965$ $17.44\%$ $0.0499$ $0.1767$ $\hat{K}=3$
+BC $n=569,d=30,K=2$ ARI NMI CE $0.7361$ $0.6352$ $7.03\%$ $0.7177$ $0.6215$ $7.56\%$ $0.6466$ $0.5532$ $9.67\%$ $0.5175$ $0.5009$ $13.71\%$ 0.0302-0.0302 $0.0553$ $42.18\%$ $\boldsymbol{0.8241}$ $\boldsymbol{0.7324}$ $\boldsymbol{4.57\%}$ $0.0024$ $0.0188$ $37.08\%$
+Seg $n=2086,d=18,K=7$ ARI NMI CE $\boldsymbol{0.6278}$ $\boldsymbol{0.6971}$ $\boldsymbol{21.86}\%$ $0.4782$ $0.6141$ $41.04\%$ $0.4862$ $0.6839$ $43.29\%$ $0.5832$ $0.7265$ $34.42\%$ $0.2407$ $0.5731$ $53.40\%$ $0.3829$ $0.5133$ $48.95\%$ $0.4981$ $0.6178$ $\hat{K}=8$
+Yale B $n=2000,d=600,K=10$ ARI NMI CE $\boldsymbol{0.9846}$ $\boldsymbol{0.9851}$ $\boldsymbol{0.70\%}$ $0.8461$ $0.9306$ $13.25\%$ $0.1549$ $0.4090$ $60.20\%$ $0.7427$ $0.8765$ $17.35\%$ $0.1725$ $0.5722$ $61.10\%$ $0.8421$ $0.9097$ $14.65\%$ $0.5882$ $0.7744$ $33.80\%$
+Optidigits $n=5620,d=64,K=10$ ARI NMI CE $0.8233$ $0.8394$ $8.83\%$ $0.4923$ $0.6486$ $35.96\%$ $0.0144$ $0.1290$ $82.99\%$ $\boldsymbol{0.8543}$ $\boldsymbol{0.8978}$ $\boldsymbol{8.75\%}$ $0.0003$ $0.0426$ $88.19\%$ $0.7326$ $0.7900$ $22.14\%$ $0.5671$ $0.6549$ $\hat{K}=11$
+Pendigits $n=10992,d=16,K=10$ ARI NMI CE $0.6649$ $0.7433$ $18.31\%$ $0.5749$ $0.6912$ $27.85\%$ $0.5174$ $0.6484$ $30.04\%$ $0.6739$ $0.7840$ $19.17\%$ $0.0013$ $0.0754$ $86.14\%$ $\boldsymbol{0.7695}$ $\boldsymbol{0.8368}$ $\boldsymbol{12.50\%}$ $0.6696$ $0.7497$ $\hat{K}=17$
+HAR $n=10299,d=561,K=6$ ARI NMI CE $\boldsymbol{0.6693}$ $\boldsymbol{0.7248}$ $\boldsymbol{17.74}\%$ $0.4494$ $0.5873$ $42.81\%$ $0.2932$ $0.4633$ $55.20\%$ $0.4995$ $0.6468$ $35.51\%$ $0.3279$ $0.5455$ $64.21\%$ $0.5994$ $0.6355$ $35.74\%$ $0.4075$ $0.5633$ $\hat{K}=9$
+Anuran $n=7195,d=22,K=10$ ARI NMI CE $0.8592$ $0.7146$ $\hat{K}=11$ $0.5859$ $0.6551$ $34.37\%$ $0.5145$ $0.6229$ $58.21\%$ $0.7895$ $0.6489$ $25.70\%$ $0.0289$ $0.0727$ $50.02\%$ $\boldsymbol{0.8737}$ $\boldsymbol{0.7559}$ $\boldsymbol{24.04\%}$ $0.8296$ $0.6857$ $\hat{K}=9$
+Leukemia $n=72,d=3571,K=2$ ARI NMI CE $0.8897$ $0.8056$ $2.78\%$ $0.8897$ $0.8056$ $2.78\%$ $0.2127$ $0.2178$ $26.39\%$ $0.7851$ $0.6790$ $5.56\%$ $0.0095$ $0.0036$ $34.72\%$ $\boldsymbol{0.9440}$ $\boldsymbol{0.8954}$ $\boldsymbol{1.39\%}$ $0.5905$ $0.5536$ $11.11\%$
+SuCancer $n=174,d=7909,K=2$ ARI NMI CE $\boldsymbol{0.1154}$ $\boldsymbol{0.2317}$ $\boldsymbol{32.76}\%$ 0.0030-0.0030 $0.0041$ $47.70\%$ 0.0045-0.0045 $0.0006$ $48.28\%$ $\boldsymbol{0.1154}$ $\boldsymbol{0.2317}$ $\boldsymbol{32.76}\%$ 0.0022-0.0022 $0.0290$ $49.43\%$ $\boldsymbol{0.1154}$ $0.1808$ $\boldsymbol{32.76\%}$ $0.0169$ $0.0214$ $42.53\%$
+VI Conclusions and Discussions
 In this paper, we have proposed an innovative clustering framework, DLCC, which first defines a depth-based similarity matrix to provide orderings of neighbors for each observation.
@@
 Our results show that DLCC delivers superior clustering performances under a wide range of scenarios.
-This includes handling both convex and non-convex shapes, managing clusters of balanced and unbalanced sizes, working with well-separated and overlapping clusters, and analyzing clustering problems in both low and high-dimensional spaces, including datasets with small $n$ and large $d$.
-For future research, a key direction is to develop an efficient DLCC method for handling larger datasets.
-Moreover, while this paper focuses on MD and SD for their computational simplicity, exploring other depth definitions, such as the integrated rank-weighted depth [65], could further enhance the DLCC framework.
-Furthermore, as depth is also a well-defined concept for functional data, e.g., band depth [66], there are many potential avenues for extending the DLCC framework to functional data analysis.
+This includes handling both convex and non-convex shapes, managing clusters of balanced and unbalanced sizes, working with well-separated or overlapping clusters, and analyzing clustering problems in both low and high-dimensional spaces.
+Despite these advantages, DLCC has several limitations that deserve further investigation.
+First, although DLCC can handle datasets with on the order of $10^{4}$ observations, its computational cost is dominated by constructing the depth-based similarity matrix, which can limit scalability to substantially larger datasets.
+One potential direction is to incorporate fast approximate neighborhood search to estimate neighborhoods of points for defining local centers (e.g., [70]), at the cost of weakening some desirable properties of depth-induced neighborhoods (e.g., invariance under similarity transformations for SD).
+Another direction is to adopt a sampling or landmark strategy i.e., estimate local centers from a representative subset.
+Then compute depth-based similarities only between each observation and these local centers, thereby avoiding the full $n\times n$ similarity matrix construction.
+Second, DLCC currently requires user intervention in selecting parametes, the neighborhood size $s$ (and, under the max strategy, the threshold $\delta$), typically based on evaluations of clustering outputs.
+This reduces ease of use in practice.
+A promising future direction is an adaptive DLCC that automatically determines observation-specific neighborhood sizes, performs grouping in a data-driven manner, and reduces or eliminates manual strategy selection.
+Finally, while this paper focuses on spatial depth due to its computational simplicity in high-dimensional spaces, the DLCC framework itself is not tied to a single depth notion.
+It would be of interest to explore alternative depth definitions, such as integrated rank-weighted depth [71], and to study how different depth choices affect the performance of DLCC across diverse data sets.
 Acknowledgments
 This work is supported by the Canada Research Chairs program (PM), a Dorothy Killiam Fellowship (PM), and respective NSERC Discovery Grants (AL, PM).
+Supplementary materials
+Classification method selection
+The last stage of our proposed DLCC algorithm is classification based on pseudo-labels (temporary clusters).
+For the synthetic data sets under the max strategy, we simply apply k-nearest neighborhood (kNN) on the similarity matrix $\mathbb{S}$, since kNN is well suited to non-convex and arbitrarily shaped clusters by relying only on local neighborhood information.
+For the min-strategy synthetic data sets and all real-world data sets, we consider both kNN and random forests (RF), and select the classification method using the CVDD criterion.
+Specifically, we compute the CVDD score under each candidate classifier and choose the one with the larger value.
+In the event of a tie, we select kNN, as it is deterministic (no randomness) and computationally more efficient than RF.
+Table IV reports the CVDD values for kNN and RF across all data sets, along with the selected classifier.
+TABLE IV: CVDD-based selection of the classifier in the final DLCC stage.
+For each data set, we report the CVDD scores under kNN and RF, and select the classifier with the larger CVDD.
+Ties are resolved in favor of kNN.
+Dataset CVDD (kNN) CVDD (RF) Selected
+3_blobs $3459.50$ $3243.30$ kNN
+diamond9 $12451.00$ $12531.00$ RF
+simubyPassino $529.22$ $564.42$ RF
+Iris $142.55$ $154.87$ RF
+seed $83.14$ $65.80$ kNN
+seg $1.69$ $1.48$ kNN
+wine $30.75$ $27.66$ kNN
+bc $9.10$ $7.22$ kNN
+leukemia $3.25$ $3.25$ kNN
+SuCancer $1.69$ $1.69$ kNN
+pa $6.86$ $7.96$ RF
+yale $9.48$ $10.44$ RF
+digit $7.22$ $6.82$ kNN
+pendigi $124.07$ $73.86$ kNN
+har $375.40$ $562.92$ RF
+frog $3.15$ $1.68$ kNN
+Graphs
+Here, we provide visualizations for all $25$ data sets described in the paper, including the ground truth labels, grouped local centers, the resulting temporary clusters, and the final DLCC clustering outputs.
+These graphs are intended to give readers additional insight into the mechanics of DLCC, as well as to highlight the differing behaviors of the min and max strategies.
+Specifically, the min strategy tends to act more like center-based approaches, whereas the max strategy more closely resembles density-based methods.
+For data sets with dimension greater than three, we use tSNE to embed the observations into two dimensions prior to visualization.
+We caution that inter-point distances in tSNE plots should not be interpreted quantitatively, as the embedding may distort global geometry and can be visually misleading.
+Figure 8: DLCC visualization on the 3 Blobs dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 9: DLCC visualization on the Breast Cancer dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 10: DLCC visualization on the Diamond dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 11: DLCC visualization on the HAR dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 12: DLCC visualization on the Iris dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 13: DLCC visualization on the Optidigits dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 14: DLCC visualization on the Pendigits dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 15: DLCC visualization on the Ray dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 16: DLCC visualization on the Seed dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 17: DLCC visualization on the Segmentation dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 18: DLCC visualization on the Wine dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 19: DLCC visualization on the Yale dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 20: DLCC visualization on the Aggregation dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 21: DLCC visualization on the Atom dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 22: DLCC visualization on the Bainba dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 23: DLCC visualization on the Blend dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 24: DLCC visualization on the Chainlink dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 25: DLCC visualization on the Cuboids dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 26: DLCC visualization on the Flame dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 27: DLCC visualization on the Anuran calls dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 28: DLCC visualization on the Jain dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 29: DLCC visualization on the Leukemia dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 30: DLCC visualization on the Parkinsons dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 31: DLCC visualization on the Starbeam dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.
+Figure 32: DLCC visualization on the SuCancer dataset. (a) Ground truth labels; (b) Grouped local centers; (c) Temporary clusters; (d) Final DLCC clustering result.